### v3

# Improving Expressivity of Graph Neural Networks using Localization

Abstract
In this paper, we propose localized versions of Weisfeiler-Leman (WL) algorithms in an effort to both increase the expressivity, as well as decrease the computational overhead. We focus on the specific problem of subgraph counting and give localized versions of $k-$WL for any $k$. We analyze the power of Local $k-$WL and prove that it is more expressive than $k-$WL and at most as expressive as $(k+1)-$WL. We give a characterization of patterns whose count as a subgraph and induced subgraph are invariant if two graphs are Local $k-$WL equivalent. We also introduce two variants of $k-$WL: Layer $k-$WL and recursive $k-$WL. These methods are more time and space efficient than applying $k-$WL on the whole graph. We also propose a fragmentation technique that guarantees the exact count of all induced subgraphs of size at most 4 using just $1-$WL. The same idea can be extended further for larger patterns using $k>1$. We also compare the expressive power of Local $k-$WL with other GNN hierarchies and show that given a bound on the time complexity, our methods are more expressive than the ones mentioned in Papp and Wattenhofer (2022a).
11footnotetext: Indian Institute of Technology, Gandhinagar, India**footnotetext: Equal Contribution
1 Introduction
Graphs have been used for representing relational and structural data that appear in a variety of domains, ranging from social network analysis and combinatorial optimization to particle physics and protein folding Dill et al. (2008). In order to learn representations of these data for various downstream learning tasks such as graph classification, graph neural networks (GNNs) have emerged as very effective models. Given the various types of GNN-based models developed in recent years Kipf and Welling (2017); Veličković et al. (2018); Hamilton et al. (2018); Xu et al. (2019), researchers have attempted to characterize the expressive power of these models. Morris et al. (2019) showed the equivalence between message-passing GNNs and 1-Weisfeiler-Leman (WL) algorithm Weisfeiler and Leman (1968), which is a well known combinatorial technique for checking graph isomorphism and similarity. They also showed the equivalence between $k$-GNNs and $k-$WL. Thus, in this paper, by $k-$WL, we will be referring to the equivalent $k$-GNN model. In general, the expressiveness of $k-$WL, for any $k,$ is measured by its ability to identify non-isomorphic graphs and subgraphs. In this paper, we are using Folklore WL and state the results accordingly.
It has been shown that $(k+1)-$WL is more expressive than $k-$WL. The time and space complexity increases exponentially with $k$. Thus, it is infeasible to run $k-$WL on large graphs. Also, the $k-$WL hierarchy is crude as $3-$WL identifies almost all non-isomorphic graphs. Arvind et al. (2017) characterized the graphs that can be identified by $1-$WL. So, we are interested in coming up with a GNN hierarchy that can be easily extended without much computational overhead. More specifically, we want to define a GNN hierarchy whose expressiveness lies between $k-$WL and $(k+1)-$WL.
A count of specific patterns is very useful in determining the similarity between two graphs. However, detecting and counting the number of subgraphs is generally NP-complete as it is a generalization of the clique problem. There have been various works on efficient algorithms for some fixed patterns and restricted host graph classes Bressan (2018); Shervashidze et al. (2009); Bouritsas et al. (2020); Zhao et al. (2022); Shervashidze et al. (2011); Komarath et al. (2023); Ying et al. (2019); Liu et al. (2019). In Arvind et al. (2020) characterizes patterns whose count is invariant for $1-$WL and $2-$WL equivalent graphs. Also, there exists a GNN hierarchy, $S_{k}$ Papp and Wattenhofer (2022a), where each node has an attribute that counts the number of induced subgraphs of size at most $k$, that the node is participating in. It would be interesting to see whether a scalable GNN hierarchy exists, that is comparable to the $k-$WL and $S_{k}$ hierarchy. There also exists a GNN hierarchy, $M_{k},$ Papp and Wattenhofer (2022a); Huang et al. (2023)in which $k$ vertices are marked or deleted and a GNN model is run on the modified graph.
Various subgraph based GNN models have been proposed that tackle these questions Zhao et al. (2022); Zhang and Li (2021); Morris et al. (2018); Alvarez-Gonzalez et al. (2022); Maron et al. (2019); You et al. (2021); Frasca et al. (2022); Morris et al. (2021); Bevilacqua et al. (2021); Papp and Wattenhofer (2022b); Barceló et al. (2021); Huang et al. (2023). These GNNs have been effective in capturing more fine-grained patterns and relationships within the graphs, and are scalable for large graphs. Also, it has been shown that subgraph GNNs are more expressive than the traditional ones. Frasca et al. (2022) gave an upper bound on the expressive power of subgraph $1-$WL. This leads to the question of coming up with upper and lower bounds for the expressiveness of subgraph $k-$WL for arbitrary $k$.
Consider the task of counting the occurrence of $H$ as subgraph or induced subgraph in the host graph $G$. Given the effectiveness of subgraph $k-$WL methods in increasing the expressiveness of GNNs, we want to extend this method to the subgraph itself and check whether we can fragment the subgraph and learn the count of fragments of the subgraphs to get the actual count of $H$ in the graph. We are interested in evaluating its expressiveness in terms of subgraph and induced subgraph counting. Also, if there exists a GNN hierarchical model, we want to compare its expressiveness to pre-existing GNN hierarchies, as done in Papp and Wattenhofer (2022a).
1.1 Our Contributions
In this paper, we attempt to answer these questions. The main contributions of our work are listed below:
  1. 1.
    Characterizing expressiveness of Local $k-$WL: Given a graph $G=(V,E),$ we extract a $r$-hop subgraph rooted at each vertex $v\in V,$ say $G_{v}^{r},$ and run $k-$WL on $G_{v}^{r}.$ While GNNs based on subgraphs have been proposed in recent papers , this is the first work that gives both upper and lower bounds for the expressiveness of Local $k-$WL. We are also the first to characterize patterns that can be counted exactly by Local $k-$WL.
  2. 2.
    Layer $k-$WL: To improve the space and time complexity of Local $k-$WL, we propose the Layer $k-$WL method. For this method, instead of running $k-$WL on $G_{v}^{r}$, we run it on two consecutive layers of vertices. Here, the $i$th layer of vertices refers to the vertices that appear at an $i$-hop distance from $v$(or the $i$th layer breadth-first search(BFS)).
  3. 3.
    Recursive WL : Recursive WL is an alternative to $k-$WL. In this method, we first run $1-$WL to get a partitioning of vertices. Then we run $(k-1)-$WL on the vertices of each partition separately. It can be shown that this method is more expressive than $(k-1)-$WL and less expressive than $k-$WL. Also, since we are running $(k-1)-$WL on a smaller set of vertices, it has better space and time complexity than running $k-$WL.
  4. 4.
    Fragmentation : For the counting task, based on the pattern $H$ to be counted, the subgraph $G_{v}^{r}$ is further decomposed into simpler patterns for which the exact counts of subpatterns are already known. Thus, we need to learn the easier tasks in the subgraphs of $G_{v}^{r}$. So, a smaller $k$ is sufficient to count the number of patterns. Using this method, we show that all the patterns appearing as induced subgraphs of size four can be counted using just $1-$WL. This technique can be useful for counting larger and more complicated patterns. Thus, instead of training a GNN for the larger subgraph, we can train GNN models for the smaller patterns for counting and then combine their counts to get the count of the larger subgraph. Using the fragmentation technique, we use the model learned for predicting $K_{3}$ or a triangle to predict the number of $K_{4}$ in the graph. Similarly, if we have a model that can predict $K_{n}$ in a graph, then we can use it to predict $K_{n+1}$. In other words, we can reduce the counting of $K_{n+1}$ to a triangle counting problem with a minimal increase in the number of parameters.
  5. 5.
    Comparison with other GNN models : Papp and Wattenhofer (2022a) shows an analysis of four GNN models. We do a similar analysis for our models and compare them with the models mentioned in Papp and Wattenhofer (2022a). We show that our models are more expressive than the ones presented in that paper.
Outline of the paper :
In Section 2, we introduce some of the terms used throughout the paper. In Section 4, we introduce the localized variants of the $k-$WL algorithm and analyze their space and time complexities. In Section 5, we give theorems that characterize the expressiveness of the localized $k-$WL variants proposed in our work. In Section 6, we characterize the expressiveness of our methods in terms of subgraph and induced subgraph counting. We also discuss how to count the occurrences of $H$ in $G,$ using localized algorithms. We discuss the fragmentation approach in Section 7, followed by a theoretical comparison of GNN models in Section 8. The model architecture, along with the parameters used for the experiment, is explained in Section 9 We report the results of our experiments in Section 10 and conclude the paper with Section 11.
2 Preliminaries
We consider a simple graph $G(V,E)$. For basic definitions of graph theory, we refer the reader to West et al. (2001). The neighbourhood of a vertex $v$ is the set of all vertices adjacent to it in $G$ (denoted as $N_{G}(v)$). The closed neighbourhood of $v$ is the set of all neighbours, including the vertex $v$ (denoted as $N_{G}[v]$). A graph whose all the vertices are of same degree are called regular graph.
A graph $H$ is called a subgraph of $G$ if $V(H)\subseteq V(G)$ and $E(H)\subseteq E(G)$. The subgraph induced on $S\subseteq V(G)$ is a graph whose vertex set $S$ contains all the edges in $G$ whose endpoints are in $S$ and is denoted by $G[S]$. The induced subgraph on a $r$-hop neighbourhood around vertex $v$ is denoted by $G_{v}^{r}$. Attributed subgraphs are coloured subgraphs, also referred to as Motifs.
The maximum distance from a vertex to all other vertices is called the eccentricity of the vertex. The minimum of the eccentricity over all the vertices is called the radius of the graph. The center of the graph is set of vertices such that eccentricity is minimum. For pattern counting, we pick one of the center vertex and call it key vertex.
Homomorphism from graph $H$ to $G$ is a function from $V(H)$ to $V(G)$ such that if $\{u,v\}\in E(H)$ then $\{f(u),f(v)\}\in E(G)$. Given a pattern $H$, the set of all of its homomorphic images is called spasm of $H$. Two graphs $G$ and $H$ are isomorphic if there exists a bijective function $f:V(G)\xrightarrow{}V(H)$ such that $\{u,v\}\in E(G)$ if and only if $\{f(u),f(v)\}\in E(H)$. The orbit of a vertex $v$ in $G$ is the set of vertices to which $v$ can be mapped, and that mapping can be extended to automorphism (denoted by $Orbit_{G}(v)$).
We mention some of the structural parameters of graphs. Problems on graph with bounded structural parameters can be solved efficiently on bounded graph parameters.
Graph Parameters:
We first define tree decomposition of the graph as:
Given a graph $G$, we decompose it into tree structures, say $T,$ where the set of vertices of $T$ is a subset of set of vertices of $G.$ This decompostion has to satisfy the following constraints:
  1. 1.
    Every vertex of $G$ must lie in some bag associated with a vertex of $T$.
  2. 2.
    For each edge $\{v_{i},v_{j}\}$, there exists a bag containing having $v_{i}$ and $v_{j}$.
  3. 3.
    If a vertex $v_{i}\in V(G)$ belongs to two bags $B_{i}$ and $B_{j}$ associated with two vertices $u_{i}$ and $u_{j}$ of $T,$ then $v_{i}$ must be present in all the bags associated with the vertices belonging to the path connecting $u_{i}$ and $u_{j}$ in $T.$
The width of the tree decomposition is the maximum size of the bag minus one. The treewidth of the graph $G,$ $tw(G)$ is the minimum over all such decompositions. It is NP-hard to compute tree-width of graphs. However, there exists an efficient algorithm that checks for a fixed $k$, whether $tw(G)$ is at most $k$ Korhonen (2022); Korhonen and Lokshtanov (2022). Graph of bounded treewidth implies sparse graphs. However, for some sparse graph, the treewidth is unbounded. For example, grid graph on $n$ vertices has treewidth $\sqrt{n}$. The maximum of the treewidth over all its homomorphic images is called the hereditary treewidth of pattern $H,$ denoted by $htw(H)$.
Planar graphs are graphs that are $k_{5}$ and $k_{3,3}$ minor free. One can also say the graph that can be redrawn such that no edges cross each other are planar graphs. Also, the number of edges can be at most linear in the number of vertices. The Euler genus of a graph is defined in a similar manner. The genus of a graph is the minimum number such that the graph can be drawn on circle without intersecting edges.
Now, we look at graph classes that are dense but have nice structure such as complete graphs. Clique width has been defined for dense graphs. However, there does not exist efficient algorithm to check whether clique width of the given graphs is $k$ for $k\geq 4$. Rankwidth has been defined by Robertson and Seymour to handle dense graph classes. Given a graph $G$, to construct rankwidth decomposition, we define subcubic tree $T$. Given a bijection from the set of leaves of a tree to the set of vertices of $G$, we can construct rankwidth decomposition with respect to that bijection. The parent vertices in the tree contains union of vertices belonging to its two children. Note that deletion of single edge disconnects the tree and vertices of graph get partitioned into two subparts say $X$ and $V(G)\setminus X$. We define submatrix of adjacency matrix $A(X,V(G)\setminus X)$ where $a_{i,j}=1$ if and only if $i\in X$ and $j\in V(G)\setminus X$ are adjacent. Also, it has been shown that bounded clique width means bounded rank-width and vice versa Oum (2017).
3 Weisfeiler Leman Algorithm
Weisfeiler-Leman (WL) is a well known combinatorial algorithm that has many theoretical and practical applications. Color refinement($1-$WL) was first introduced in 1965 in Morgan (1965). The algorithm goes as follows:
  • Initially, color all the vertices as color 1.
  • In the next iteration $i,$ we color the vertices by looking at the number of colors of vertices adjacent to each vertex $v,$ in the $(i-1)$th iteration, as
    $C_{i}(v)=(C_{i-1}(v),\{\{C_{i}(w)\}\}_{w\in N_{G}(v)})$
    We assign a new color to the vertices, according to the unique tuple it belongs to. This partitions the vertex set in every iteration according to their color classes.
  • The algorithm terminates if there is no further partition. We call the color set a stable color set.
  • We can also observe that if two vertices get different colors at any stage $i,$ then they will never get the same color in the later iterations. We can observe that the number of iterations is at most $n$ as a vertex set ,$V(G),$ can be partitioned at most $n$ many times.
  • The color class of any vertex $v\in V(G)$ can appear at most $1+\log n$ times and the running time is $\mathcal{O}(n^{2}\log n)$ Immerman and Sengupta (2019).
In case we need to run only $h$ iterations and stop before getting the stable color, then the running time is $O(nh)$.
The same idea was extended by Weisfeiler and Leman in which instead of coloring vertex, they colored all the two tuples based on edge, non-edge and $(v,v)$. In later iteration, the color gets refined for each two tuples based on their neighbourhood and common neighbourhood. This partition the set of two tuples of vertices. The iteration in which no further partition is being done are called stable coloring. Weisfeiler Leman algorithm which is known as $2-$WL algorithm.
Similar approach was extended later for coloring $k$-tuples and then do refinement of coloring in later iterations.
Definition 1.
Let $\vec{x}=(x_{1},...,x_{k})\in V^{k}$,$y\in V$ , and $1\leq j\leq k$. Then, let $x[j,y]\in V^{k}$ denote the $k$-tuple obtained from $x$ by replacing $x_{j}$ by $y$. The $k$-tuples $\vec{x}[j,y]$ and $\vec{x}$ are said to be $j$-neighbors for any $y\in V$ We also say $\vec{x}[j,y]$ is the j-neighbor of $\vec{x}$ corresponding to $y$.
  • Color all the $k$-tuple vertices according to their isomorphic type. Formally, $(v_{1},v_{2},....,v_{k})$ and $(w_{1},w_{2},....,w_{k})$ get the same color if $v_{i}=v_{j}$ then $w_{i}=w_{j}$ and also, if $(v_{i},v_{j})\in E(G),$ then $(w_{i},w_{j})\in E(G)$.
  • In every iteration, the algorithm updates the color of the tuple after seeing the color of its adjacent $k$ tuple vertices.
    $C_{i+1}^{k}(\vec{v}):=(C_{i}^{k}(\vec{v},M_{i}(\vec{v})$
    where $M_{i}(\vec{v})$ is the multiset
    $\{\{(C_{i}^{k}(v_{1},v_{2},...,v_{k1},w),...,C_{i}^{k}(v_{1},v_{2},..,w,..,v_{k}),...,C_{i}^{k}(w,v_{2},...,v_{k}))\mid w\in V\}\}$
  • The algorithm terminates if there is no further partition. We call the color set a stable color set.
  • We also observe that if two tuples get different colors at any stage $i,$ then they will never get the same color in the later iterations. We can observe that the number of iterations is at most $n^{k}$ as $V^{k}$ can be partitioned at most $n^{k}$ many times.
  • The color class of any vertex $\vec{v}\in V^{k}$ can appear at most $\mathcal{O}(k\log n)$ times and running time is $\mathcal{O}(k^{2}n^{k+1}\log n)$ Immerman and Sengupta (2019).
Two graphs $G$ and $H$ are said to be $k-$WL equivalent ($G\simeq_{k}H$), if their color histogram of the stable colors matches. We say that $G$ is $k-$WL identifiable if there doesn’t exist any non-isomorphic graphs that are $k-$WL equivalent to $G$.
Color refinement ($1-$WL) can recognise almost all graphs Babai et al. (1980), while $2-$WL can recognise almost all regular graphs Bollobás (1982). The power of $WL$ increases with an increase in the value of $k$. The power of $k-$WL to distinguish two given graphs is same as with counting logic $C^{k+1}$ with $(k+1)$-variable. Also, the power of $k-$WL to distinguish two non-isomorphic graphs is equivalent to spoiler’s winning condition in $(k+1)$-bijective pebble game. Recently, Dell et al. (2018) has shown that the expressive power of $k-$WL is captured by homomorphism count. It has been shown that $G_{1}\simeq_{k}G_{2}$ if and only if $Hom(T,G_{1})=Hom(T,G_{2}),$ for all graphs $T$ of treewidth at most $k$.
The graphs that are identified by $1-$WL are Amenable graphs. There is a complete characterization of the amenable graphs in Arvind et al. (2017); Kiefer et al. (2015). In the original algorithm, we initially color all the vertices with color $1$. However, if we are given a colored graph as input, we start with the given colors as the initial colors. Also, we can color the edges, and run $1-$WL Kiefer et al. (2015).
Even if ${k-WL}$ may not distinguish two non-isomorphic graphs, two ${k-WL}$ equivalent graphs have many invariant properties. It is well known that two ${1-WL}$ equivalent graphs have the same maximum eigenvalue. Two graphs that are ${2-WL}$ equivalent are co-spectral and have the same diameter. Recently, V. Arvind et al. have shown the invariance in terms of subgraph existence and counts Arvind et al. (2020). They show the complete characterization of subgraphs whose count and existence are invariant for ${1-WL}$ equivalent graph pairs. They also listed down matching , cycles and path count invariance for ${2-WL}$. Also, there is a relation between homomorphism count and subgraph count Curticapean et al. (2017). The count of subgraphs is a function of the number of homomorphism from set of all homomorphic image of patterns. Hereditary treewidth of graph is defined as maximum of treewidth over all homomorphic images. So, if two graphs are ${k-WL}$ equivalent, then the count of all subgraphs whose $htw(G)$ is almost $k$ are same . However, running ${k-WL}$ takes $O(k^{2}\cdot n^{k+1}logn)$ time and $O(n^{k})$ space Immerman and Sengupta (2019). So, it is not practically feasible to run ${k-WL},$ for large $k,$ for graphs .
The expressive power of $k-$WL is equivalent to first order logic on $(k+1)$ variables with a counting quantifier. Let $G=(V,E),$ where $V$ is a set of vertices and $E$ is a set of edges. In logic, we define $V$ as the universe and $E$ as a binary relation. In Cai et al. (1992), they have proved that the power to distinguish two non-isomorphic graphs using $k-$WL is equivalent to $C^{k+1},$ where $C^{k+1}$ represents first order logic on $(k+1)$ variables with counting quantifiers (stated in Theorem 1). To prove this, they define a bijective $k$-pebble game, whose power is equivalent to $C^{k}$.
Bijective k-Pebble Game
The bijective k-Pebble game ($BP_{k}(G,H)$) has been discussed in Kiefer (2020); Cai et al. (1992); Grohe and Neuen (2021). Let graphs $G$ and $H$ have the same number of vertices and $k\in\mathbb{N}$. Let $v_{i},v\in V(G)$ and $w_{i},w\in V(H)$.
Definition 2.
The position of the game in bijective pebble game is the tuples of the vertices where the pebbles are placed.
The bijective $k$-pebble game is defined as follows:
  1. 1.
    Spoiler and Duplicator are two players.
  2. 2.
    Initially, no pebbles are placed on the graphs. So, the position of the game is ((),()) (the pairs of empty tuples.)
  3. 3.
    The game proceeds in the following rounds as follows:
    1. (a)
      Let the position of the game after the $i^{th}$ round be $((v_{1},...,v_{l}),(w_{1},w_{2},...,w_{l}))$. Now, the Spoiler has two options: either to play a pebble or remove a pebble. If the Spoiler wants to remove a pebble, then the number of pebbles on the graph must be at least one and if Spoiler decides to play a pebble then number of pebbles on that particular graph must be less than $k$.
    2. (b)
      If the Spoiler wants to remove a pebble from $v_{i},$ then the current position of the game will be $((v_{1},v_{2},...v_{i-1},v_{i+1},..,v_{l}),(w_{1},w_{2},...w_{i-1},w_{i+1},..,w_{l}))$. Note that, in this round, the Duplicator has no role to play.
    3. (c)
      If the Spoiler wants to play a pair of pebbles, then the Duplicator has to propose a bijection $f:V(G)\xrightarrow{}V(H)$ that preserves the previous pebbled vertices. Later, the Spoiler chooses $v\in V(G)$ and sets $w=f(v)$. The new position of the game is $((v_{1},...v_{l},v),(w_{1},w_{2},...,w_{l},w))$.
The Spoiler wins the game if for the current position $((v_{1},...v_{l},v),(w_{1},w_{2},...,w_{l},w)),$ the induced graphs are not isomorphic. If the game never ends, then the Duplicator wins. The equivalence between the bijective $k$-pebble game and $k-$WL was shown in the following theorem.
Theorem 1.
Cai et al. (1992) Let $G$ and $H$ be two graphs. Then $G\simeq_{k}H$ if and only if the Duplicator wins the pebble game $BP_{k+1}(G,H).$
A stronger result, namely, the equivalence between the number of rounds in the bijective $(k+1)$-pebble game and the iteration number of $k-$WL was stated in the following theorem.
Theorem 2.
Kiefer (2020) Let $G$ and $H$ be graphs of same size. The vertices may or may not be colored. Let $\vec{u}:=(u_{1},...,u_{k})\in(V(G))^{k}$ and $\vec{v}:=(v_{1},...,v_{k})\in(V(H))^{k}$ be any two arbitrary elements. Then, for all $i\in\mathbb{N}$, the following are equivalent :
  1. 1.
    The color of $\vec{u}$ is same as the color of $\vec{v}$ after running $i$ iterations of $k-$WL.
  2. 2.
    For every counting logic formulae with $(k+1)$ variables of quantifier depth at most $i$, $G$ holds the formula if and only if $H$ does so.
  3. 3.
    Spoiler does not win the game $BP_{k+1}(G,H)$ with the initial configuration $(\vec{u},\vec{v})$ after at most $i$ moves.
4 Local k-WL based Algorithms for GNNs
In this section, we present the local $k-$WL based algorithms for GNNs. We also give runtime and space requirements for such GNNs.
4.1 Local k-WL
Given a graph $G$, we extract the subgraph induced on a $r$-hop neighbourhood around every vertex. We refer to it as $G_{v}^{r},$ for the subgraph rooted at vertex $v$ in $G$. Then, we colour the vertices in $G_{v}^{r}$ according to their distances from $v$. Now, we run $k-$WL on the coloured subgraph $G_{v}^{r}$. The stable colour obtained after running $k-$WL is taken as the attributes of vertex $v$. Then, we run a GNN on the graph $G$ with the attributes on each vertex $v$. This is described in Algorithm 1.
Algorithm 1 Local k-WL
1:Input: $G,r,k$
2:for each vertex $v$ in $V(G)$ do
3:     Find the subgraph induced on the $r$-hop neighborhood rooted at vertex $v$ ($G_{v}^{r}$).
4:     Color the vertices whose distance from $v$ is $i,$ by color $i$
5:     Run $k-$WL on the colored graph until the colors stabilize.
6:end for
7:Each vertex has as an attribute the stable coloring of vertex $v$ obtained from $G_{v}^{r}$.
8:Run GNN on the graph $G$ with each vertex having attributes as computed above.
Runtime and Space requirement Analysis : The time required to run $k-$WL on $n$ vertices is $O(n^{k+1}\log(n))$. Here, we run $k-$WL on a $r$-hop neighborhood instead of the entire graph. So, $n$ is replaced by $n_{1},$ where $n_{1}$ is the size of the neighborhood. If a graph has bounded degree $d$, and we run $k-$WL for a $2$-hop neighborhood, then $n_{1}$ is $O(d^{2})$. Also, we have to run Local $k-$WL for each vertex. Hence, the total time required is $O(n\cdot d^{2k+2}\log(d))$. Also, running a traditional GNN takes time $O((n+m)\log n),$ where $m$ is the number of edges. So, if we assume that $d$ is bounded, then the time required is linear in the size of the graph. Furthermore, the space required to run $k-$WL on $n$ vertices graph is $O(n^{k})$. Hence, for Local $k-$WL, it follows that the space requirement is $O(n_{1}^{k})$.
4.2 Layer k-WL
In order to make Local $k-$WL more time and space efficient, while maintaining the same expressive power, we propose a modification to Local $k-$WL. Instead of running $k-$WL on the entire $r$-hop neighbourhood, we run $k-$WL on consecutive layers of $G_{v}^{r}$ (i.e., run $k-$WL on the set of vertices with colour $i$ and colour $(i+1)$). Initially, we run $k-$WL on the set of vertices that are at a distance of $1$ and $2$ from $v.$ Then, we run $k-$WL on the set of vertices with colors $2$ and $3,$ and so on. While running $k-$WL, initially, it partitions the $k$-tuples based on the isomorphism type. However, in this setting, we incorporate the stabilized colouring obtained in the previous round. For $l<k,$ we define the color of $l$ tuples as $col(u_{1},u_{2},...,u_{l}):=col(u_{1},u_{2},...,u_{l},\underbrace{u_{1},u_{1}..,,u_{1}}_{(k-l)\text{times}})$. Consider the mixed tuple (we call a tuple to be mixed if some of the vertices have been processed in the previous iteration and the remaining have not yet been processed) $(u_{1},v_{1},\ldots,u_{k})$ where $col(u_{j})=i$ and $col(v_{j})=i+1$ (i.e $u_{i}^{\prime}$s are the set of processed vertices and $v_{i}^{\prime}$s are yet to be processed). So, even if $(u_{1},v_{1},\ldots,u_{k})$ and $(u_{1}^{{}^{\prime}},v_{1}^{{}^{\prime}},\ldots,u_{k}^{{}^{\prime}})$ may be isomorphic, if $col(u_{1},u_{2},\ldots u_{l})\neq col(u_{1}^{{}^{\prime}},u_{2}^{{}^{\prime}},\ldots u_{l}^{{}^{\prime}})$ then $col(u_{1},v_{1},\ldots,u_{k})\neq col(u_{1}^{{}^{\prime}},v_{1}^{{}^{\prime}},\ldots,u_{k}^{{}^{\prime}})$. The algorithm is described in Algorithm 2. A GNN model incorporating Local+Layer $k-$WL is equivalent to running Layer $k-$WL in line 5 in Algorithm 1.
Algorithm 2 Layer k-WL($v$)
1:Given $G_{v}^{r},k$.
2:Run $k-$WL on the induced subgraph of levels $1$ and $2$.
3:for each layer $i$ of BFS($v$), $i\geq 2$ do
4:     Initial colour of $k-tuple$ incorporates the stabilized colour obtained from the previous iteration.
5:     Run $k-$WL on the subgraph induced on the vertices in layer $i$ and $(i+1)$
6:end for
Runtime and Space requirement Analysis.
The running time and space requirement for Layer $k-$WL depends on the maximum number of vertices in any two consecutive layers, say $n_{2}$. The time required to run $k-$WL is $O(r\cdot(n_{2})^{k+1}\log(n_{2}))$. However, running only Local $k-$WL will require $O((r\cdot n_{2})^{k+1}\log(r\cdot n_{2}))$ time. The space requirement is $O(n_{2}^{k}).$ Hence, running Layer $k-$WL is more efficient than running Local $k-$WL, especially when $r$ is large.
4.3 Recursive WL
Here, we present another variant of WL. The central idea is to decompose the graphs initially by running $1-$WL. Then, further, decompose the graphs by running $2-$WL and so on. One can note that the final vertex partition that $1-$WL outputs after color refinement are regular if we restrict to a single color class. In other words, let $G[X]$ be the induced graph on the vertices of same color. Then, $G[X]$ is regular. Also, $G[X,Y]$ where $X$ and $Y$ are sets of vertices of two different color classes. $G[X,Y]$ is also a bi-regular graph. We run $2-$WL on the regular graph. Using Bollobás (1982), we can guarantee that it would distinguish almost all regular graphs. Similarly, running $2-$WL on $G[X,Y]$ is bi-regular and thus can be distinguished by $2-$WL. We again run $1-$WL on $G,$ using the colors obtained after running $2-$WL. This further refines the colors of the vertices in $G$. One can easily check that it is more expressive than $1-$WL and less expressive than $2-$WL. We give the graph Figure 1 that can not be distinguished by Recursive $(1,2)-$WL and $1-$WL but can be distinguished by $2-$WL. This gives an intermediate hierarchy in the $k-$WL hierarchy. Also, the space and time required for running $2-$WL on the entire graph is more than that of Recursive $(1,2)-$WL. The running time and space required depend on the partition size obtained after running $1-$WL.
Note that the color of vertex $v$ is $col(v,v)$ after running $2-$WL.
Algorithm 3 Recursive(1,2) WL
1:Given $G$
2:Run $1-$WL and get the partition of vertices into colour classes.
3:Let $S=\{C_{1},C_{2},\ldots C_{l}\}$ be the color classes obtained after running $1-$WL.
4:for each color class $C_{i}$ in $S$ do
5:     Run $2-$WL on the induced subgraph in $C_{i}$ and get color partition.
6:     Let $C_{i}$ get partitioned into $C_{i,1},C_{i,2},\ldots,C_{i,l}$
7:end for
8:Run $1-$WL on the colored graph $G$ whose colors are given by steps 5 and 6.
9:for each new color class $C^{\prime}_{i}$ and $C^{\prime}_{j}$ do
10:     Run $2-$WL on the induced subgraph on the vertices in color partitions $C^{\prime}_{i}$ and $C^{\prime}_{j}$ and get new color partition.
11:end for
12:Repeat 5-11 till the colour stabilizes.
This idea can be generalized for any suitable $k.$ We can run a smaller dimensional $k_{1}-$WL and then use the partition of $k_{1}$ tuple vertices. Later, we can use this partition to get a finer partition of $k_{2}$ tuples. Assuming $k_{1}<k_{2}$, one can see that we have to run $k_{2}-$WL on smaller graphs. This reduces the time and space required for running $k_{2}-$WL on the entire graph. One can easily see that it is less expressive than $k_{2}-$WL, however, more expressive than $k_{1}-$WL.
More specifically, initially run $1-$WL and further run $(k-1)-$WL on the colored classes. One can check that it is more expressive than $(k-1)-$WL and less expressive than $k-$WL.
Figure 1: Graph identifiable by 2-WL but not by Recursive $1-$WL
5 Theoretical Guarantee of Expressive Power
In this section, we theoretically prove the expressive power of GNN models that we proposed in Section 4 in terms of graph and subgraph isomorphism. In the discussion below, we say that a GNN model $A$ is at most as expressive as a GNN model $B$ if any pair of non-isomorphic graphs $G$ and $H$ that can be distinguished by $A$ can also be distinguished by $B$. Also, we say a GNN model $A$ is at least as expressive as a GNN model $B$ if $A$ can identify all the non-isomorphic graph pairs that can be identified by GNN model $B$. The proof of the theorem and lemmas presented in the section mainly use pebble bijective game. Also, as mentioned earlier, there is equivalence between the expressivity of $k-$WL and $(k+1)$-pebble bijective game.
5.1 Local k-WL
It has been shown in recent works that running Local $1-$WL has more expressive power as compared to running $1-$WL. The upper bound on the expressive power of Local $1-$WL has been shown in Frasca et al. (2022). However, the expressive power of Local $k-$WL, for arbitrary $k,$ has not been studied. In the Theorem 3, we give a description of the expressive power of Local $k-$WL and show that it has more expressive power than $k-$WL. We also show that it is less expressive than $(k+1)-$WL. The proof techniques that we used are different from Frasca et al. (2022).
Theorem 3.
Running Local $k-$WL is more expressive than running $k-$WL on the entire graph. Also, running Local $k-$WL is at most as expressive as running $(k+1)-$WL on the entire graph.
Proof.
Let $G_{1}$ and $G_{2}$ be two graphs distinguishable by $k-$WL. So, the Spoiler has a winning strategy in the game ($G_{1}$,$G_{2}$). Suppose $G_{1}$ and $G_{2}$ do not get distinguished after running $k-$WL locally. That means the Duplicator has a winning strategy for all vertices individualized. Let $v$ in $G_{1}$ and $u$ in $G_{2}$ be the vertices that are individualized.
We play the $(k+1)$-bijective pebble game on the entire graphs $(G_{1},G_{2})$ and the local subgraphs $(G_{1}^{v},G_{2}^{u})$ simultaneously. Let $S_{1}$ and $D_{1}$ be the spoiler and duplicator in game $(G_{1},G_{2})$ respectively, and $S_{2}$ and $D_{2}$ be the spoiler and duplicator in game $(G_{1}^{v},G_{2}^{u})$. We use the strategy of $D_{2}$ to determine the move for $D_{1}$ and the strategy of $S_{1}$ to determine the move for $S_{2}$.
Initially, $D_{2}$ gives a bijection $f$ from the vertex set of $G_{1}^{v}$ to $G_{2}^{u}$. We propose the same bijection $f$ by $D_{1},$ extending it by mapping $v$ to $u$. Now, the spoiler $S_{1}$ places a pebble at some vertex $(v_{i},f(v_{i})).$ The spoiler $S_{2}$ also places a pebble at vertex $(v_{i},f(v_{i}))$. We can show using induction on the number of rounds that if $S_{1}$ wins the game, then $S_{2}$ also wins the game. Our induction hypothesis is that $S_{1}$ has not won till the $j^{th}$ round and the positions of both the games are same. Let the current position of both the games after the $j^{th}$ round be $((v_{1},v_{2},\ldots,v_{l}),(f(v_{1}),f(v_{2}),\ldots,f(v_{l}))$. Now, $S_{1}$ either decides to play a pair of pebbles or remove.
Case(1): If $S_{1}$ decides to remove a pebble. In this case, the Duplicator $D_{1}$ has done nothing to do. $S_{2}$ will copy the same strategy as $S_{1}$. Here, $S_{1}$ cannot win in this round. Also, note that the positions of both games are the same.
Case(2): If $S_{1}$ decides to play a pebble. In this case, $S_{2}$ also decides to play a pebble. Duplicator $D_{2}$ proposes a bijective function $f.$ The same bijective function is proposed by $D_{1}$. Now, $S_{1}$ places a pebble at $(v,f(v))$. $S_{2}$ also chooses the same vertices. So, the position of both the game is the same. Therefore, if $S_{1}$ wins the game, then $S_{2}$ also wins the game. Thus, running $k-$WL locally is at least as expressive as running $k-$WL on the entire graph.
We can show that it is more expressive by looking at the simple example that running $1-$WL on a local substructure can count the number of triangles, whereas running $1-$WL on an entire graph does not recognize graphs having different triangle counts. Also, one can observe that running $k-$WL locally is running $k-$WL on the colored graph where vertices at distinct distances get distinct colors. Its power is the same as individualizing one vertex and running $k-$WL. Thus, running $k-$WL locally is more expressive than running $k-$WL on the entire graph.
Let $G_{1}$ and $G_{2}$ be two graphs that can be distinguished by running $k-$WL locally. Recall that, the key vertices refer to $u$ and $v$ in $G_{1}$ and $G_{2}$ such that they are the root vertices corresponding to $G_{1}$ and $G_{2}$, respectively. This means the Spoiler has a winning strategy in the $(k+1)$ bijective pebble game, where the key vertices are matched to each other. Now, we use the strategy of the Spoiler in the local substructure to get a winning strategy for the Spoiler in the entire graph. At first, when the Duplicator gives a bijective function, the Spoiler places a pebble on the paired vertices. For the remaining moves, we copy the strategy of the Spoiler in the local structure, and the Duplicator’s strategy of the entire graph is copied to the Duplicator’s strategy of the local structures. Thus, if the Spoiler has a winning strategy in the local substructure, then the Spoiler wins the $(k+2)-$ bijective pebble game on entire graphs. ∎
5.2 Layer k-WL
We presented an algorithm (Algorithm 2) for applying $k-$WL to consecutive layers in a $r$-hop subgraph for a vertex $v\in V.$ This improves the time and space efficiency of the Local $k-$WL method as we have discussed above. We now describe the expressive power of Layer $k-$WL. In the following lemmas, we show that the expressive power of Layer $k-$WL is the same as that of Local $k-$WL.
Lemma 1.
Running $k-$WL on the entire $r$-hop neighbourhood is at least as expressive as running Layer $k-$WL.
Proof.
Let $G$ and $H$ be the subgraphs induced on the $r$-hop neighborhood. Let $(S,D)$ be the Spoiler-Duplicator pair for the game $(G,H)$. Similarly, let $(S_{i},D_{i})$ be the Spoiler-Duplicator pair for the game $(G_{i},H_{i}),$ where $G_{i}$ and $H_{i}$ are the subgraphs induced on the vertices at the $i$th and $(i+1)$th layers of $G$ and $H,$ respectively. We claim that if any of the $S_{i}^{\prime}$s has a winning strategy in the game $(G_{i},H_{i})$, then $S$ has a winning strategy in the game $(G,H)$. Here, the strategy of $D$ is copied by $D_{i}$, and the strategy of $S_{i}$ is copied by $S$. We prove this using induction on the number of rounds of the game. Our induction hypothesis is that the positions of both the games are same, and if $S_{i}$ wins after $t$ rounds, then $S$ also wins after $t$ rounds.
Base case: $D$ proposes a bijective function $f:V(G)\xrightarrow{}V(H)$. Note that the bijection must be color-preserving; otherwise, $S$ wins in the first round. Thus, we can assume that $f$ is color-preserving. So, $D_{i}$ proposes the restricted function $f_{i}$ as a bijective function from $V(G_{i})$ to $V(H_{i})$. Now, $S_{i}$ plays a pair of pebbles in $(G_{i},H_{i})$, and $S$ also plays the same pair of pebbles in the game $(G,H)$. It is easy to see that both games’ positions are the same. Also, if $S_{i}$ wins, then the number of vertices of a particular color is different. Hence, $S$ also has a winning strategy.
By the induction hypothesis, assume that after the $t^{th}$ round $S_{i}$ did not win and the position of the game is the same for both games.
Consider the $(t+1)^{th}$ round in both games. $S_{i}$ either chooses to play or remove a pebble. If $S_{i}$ chooses to remove a pebble, so will $S$. Again, the position of both the games is same. Now, if $S_{i}$ decides to play a pair of pebbles, then $S$ also decides to play a pair of pebbles. So, $D$ proposes a bijective function and $D_{i}$ proposes a restricted bijective function. Now, if $S_{i}$ plays a pair of pebbles at $(v,f_{i}(v))$, then $S$ also decides to play a pair of pebbles at $(v,f(v))$. Thus, the position of the game is same in both of the games. This ensures that if $S_{i}$ has won, then $S$ also wins.
Lemma 2.
Running Layer $k-$WL is as expressive as running $k-$WL on the entire induced subgraph.
Proof.
Let $G$ and $H$ be the subgraphs induced on a $r$-hop neighborhood. Let $(S,D)$ be the Spoiler-Duplicator pair for the game $(G,H)$. Similarly, let $(S_{i},D_{i})$ be the Spoiler-Duplicator pair for the game $(G_{i},H_{i}),$ where $G_{i}$ and $H_{i}$ are the subgraphs induced on the vertices at the $i$th and $(i+1)$th layers of $G$ and $H,$ respectively. We claim that if $S$ has a winning strategy in the game $(G,H),$ then there exists $S_{i}$ such that it has a winning strategy in the game $(G_{i},H_{i}).$ Here, the strategy of $D$ is copied by $D_{i}$ and the strategy of $S_{i}$ is copied by $S.$ We prove the lemma using induction on the number of rounds of the game. Our induction hypothesis is that the position of the game $(G,H)$ is same for $(G_{i},H_{i}),$ for all $i,$ if we restrict it to the subgraph induced by the vertices of color $i$ and $(i+1).$ Also, if $S$ wins after round $t,$ then there exists $S_{i}$ that wins after $t$ rounds.
Base case: $D$ proposes a bijective function $f:V(G)\longrightarrow V(H)$. Note that the bijection must be color-preserving; otherwise $S$ wins in the first round. Thus, we can assume that $f$ is color-preserving. So, $D_{i}$ proposes the restricted function $f_{i}$ as a bijective function from $V(G_{i})$ to $V(H_{i}),\forall i\in[r]$. Now, $S$ will play a pair of pebbles in the game $(G,H).$ Suppose $S$ plays the pebbles at $(v,f(v))$ and $color(v)=i,$ then $S_{i}$ and $S_{i-1}$ play pebbles at $(v,f_{i}(v))$ in their first round. It is easy to see that the position of the games $(G,H)$ and $(G_{i},H_{i}),$ for all $i\in[r],$ is same if we restrict it to the subgraph induced by vertices of colors $i$ and $(i+1)$. Also, if $S$ wins, then the number of vertices of particular color are not same. So, there exists some $i,$ such that $S_{i}$ also has a winning strategy.
By induction hypothesis, assume that after the $t^{th}$ round, $S$ did not win and position of the game is same as defined.
Consider the $(t+1)^{th}$ round in both the games. $S$ either chooses to play or remove a pebble. If $S$ chooses to remove a pebble from vertex $(v,f(v)),$ then, if $v$ is colored with color $i,$ then $S_{i}$ and $S_{i-1}$ will remove a pebble from vertex $(v,f_{i}(v))$. Again, the position of both the games is same. Now, if $S$ decides to play a pair of pebbles, then each $S_{i}$ also decides to play a pair of pebbles. So, $D$ propose a bijective function and $D_{i}$ proposes a restricted bijective function. Now, suppose $S$ plays a pair of pebbles at $(v_{1},f(v_{1}))$. If $color(v_{1})=i,$ then $S_{i}$ and $S_{i-1}$ also decides to play pebbles at $(v_{1},f_{i}(v_{1}))$. Thus, the position of the game is same as defined. Now, if $S$ wins, then there exists $u$ and $v$ such that either $(u,v)\in E(G)$ and $(f(u),f(v))\notin E(H)$ or $(u,v)\notin E(G)$ and $(f(u),f(v))\in E(H)$. Similarly, there exists $S_{i}$ for which these things happen as the position induced is same. Therefore, $S_{i}$ wins for some $i.$
Thus, from above two lemmas we can say that the expressive power of Layer $k-$WL is the same as local $k-$WL.
6 Subgraph Counting Algorithms and Characterization of Patterns
Here, we characterize the expressive power of the proposed methods in terms of subgraph as well as induced subgraph counting. In this section, we provide algorithms and characterization of subgraphs that can exactly count the number of patterns appearing as subgraph or induced subgraph. As described above, we can see that the running time is dependent on the size of the local substructure and the value of $k$. The size of the subgraph is dependent on the radius of the patterns. So, we have to take a $r$-hop neighbourhood for each vertex $v$ in the host graph $G$.
In Section 6.1, we show how the value of $k$ can be decided based on the local substructure of the host graph. It is independent of the structure of the pattern. Also, it gives an upper bound on the value of $k$ that can count pattern appearing as subgraph and induced subgraph. In Section 6.2, we first show that doing local count is sufficient for induced subgraph count and later we give the upper bound on $k$ based on the pattern size. Note that the value of $k$ for induced subgraph count is based only on the size of pattern not its structure. In Section 6.3, we again show that locally counting subgraph is sufficient. Also, we explore the value of $k$ based on the structure of pattern. For subgraph counting the structure of pattern can be explored to get a better upperbound for the value of $k$. Later, for the sake of completeness we give algorithm to count triangle, pattern of radius one and $r.$
6.1 Deciding k based on local substructure of host graph
Here, we explore the local substructure of the host graph in which we are counting pattern appearing as graphs and subgraphs. For a given pattern of radius $r$, we explore $r$-hop neighbourhood around every vertex $v$ in the host graph $G$. If two graphs $G_{1}$ and $G_{2}$ are isomorphic then the number of subgraphs and induced subgraphs of both the graphs are the same. We use the same idea to count number of subgraphs.
Cai et. al. Cai et al. (1992), shown that $\Omega(n)$ dimension is needed to guarantee graph isomorphism. However, for restricted graph classes, we can still guarantee isomorphism for small dimensions. It has been shown that ${3-WL}$ is sufficient for planar graph Kiefer et al. (2019), ${k-WL}$ for graphs with treewidth at most $k$ Kiefer and Neuen (2019), $(3k+4)-WL$ for graphs with rankwidth at most $k$ Grohe and Neuen (2021), and $(4k+3)$ for graphs with Euler genus at most $k$ Grohe and Kiefer (2019). We say these graph classes as good graph classes. Note that, for non-isomorphic graphs, the graphs is not $k-$WL equivalent. Thus, running corresponding $k-$WL can count pattern of radius $r$, appearing as subgraph and induced subgraph.
Theorem 4.
Let $G_{v}^{r}$ denote the $r$-hop neighborhood around $v$. Given a pattern of radius $r,$ the values of $k$ that are sufficient to guarantee the count of patterns appearing either as subgraphs or induced subgraphs are:
  1. 1.
    $(3-WL)$ if $G_{v}^{r}$ planar
  2. 2.
    $(k-WL)$ if $tw(G_{v}^{r})\leq k$***tw denotes treewidth
  3. 3.
    $((3k+4)-WL)$ if $rankwidth(G_{v}^{r})\leq k$
  4. 4.
    $((4k+3)-WL)$ if $Euler-genus(G_{v}^{r})\leq k$
Proof.
Consider the subgraph induced by the vertex $v$ and its $r$-hop neighborhood in $G,$ say $G_{v}^{r},$ and the subgraph induced by the vertex $u$ and its $r$-hop neighborhood in $H,$ say $H_{u}^{r}.$ Suppose both structures belong to good graph classes. Now, we run corresponding $k$ based on the local substructure as mentioned in the theorem. If the color histogram of the stable color matches for both the graphs. This implies that both the graphs are isomorphic. Thus, the number of subgraphs and induced subgraphs in both of the substructures are also the same.
Also, we run respective $k-$WL on a colored graph, where vertices at a distance $i$ from $v$ is colored $i$. So, it is at least as expressive as running $k-$WL on an uncolored graph. We can also show that it is strictly more expressive in distinguishing non-isomorphic graphs. Thus, all the $k-$WL mentioned corresponding to good graph classes are sufficient for counting the number of patterns appearing as subgraphs and induced subgraphs. ∎
Corollary 1.
If $G_{v}^{r}$ is amenable, for all $v\in V(G),$ then Local $1-$WL outputs the exact count of the patterns appearing as subgraph and induced subgraph.
Corollary 2.
Running $1-$WL guarantees the exact number of subgraphs and induced subgraphs of all patterns of radius one, when the maximum degree of the host graph is bounded by 5.
Similarly, if the maximum degree of the host graph is bounded by $15$, then running $2-$WL is sufficient to count subgraphs and induced subgraphs of all patterns with a dominating vertex.
6.2 Counting Induced Subgraphs
The following lemma shows that we can easily aggregate the local counts of the pattern $H$ appearing as an induced subgraph to get the count of $H$ over the entire graph.
Lemma 3.
$IndCount(H,G)=\frac{\sum_{v\in V(G)}IndCount_{(u,v)}(H,G_{v}^{r})}{|Orbit_{H}(u)|}$ (1)
Proof.
Suppose a pattern $H$ in $G$ appears as an induced subgraph. So, an injective homomorphism from $V(H)$ to $V(G)$ exists, such that it preserves the edges. We fix one subgraph and find the number of mappings possible. Suppose one of the mappings maps $u_{i}$ to $v_{j}$ where $j\in|V(H)|$. Now, we can see that the number of mappings of $u_{i}$(key vertex) is the same as the size of the orbit of $u_{i}$ in $H.$ This proves the claim that every induced subgraph has been counted, the size of the orbit many times. ∎
Assume that we want to count the number of occurrences of pattern $H$ in $G$ as (an induced) subgraph. Let $u$ be the key vertex of $H$ and $r$ be the radius of $H$.
Lemma 4.
It is sufficient to look at the $r-$hop neighborhood of $v_{i}$ to compute $Count_{(u,v)}(H,G)$ or $IndCount_{(u,v)}(H,G)$.
Proof.
Suppose a subgraph exists in $G$ that is isomorphic to $H$ or isomorphic to some supergraph of $H$ with an equal number of vertices, where $u$ is mapped to $v_{i}$. The homomorphism between the graphs preserves edge relations. Consider the shortest path between $u$ and any arbitrary vertex $u_{i}$ in $H$. The edges are preserved in the homomorphic image of $H$. Therefore, the shortest distance from $f(u)$ to any vertex $f(u_{i})$ in $G$ is less than equal to $r$. So, it is sufficient to look at the $r$-hop neighborhood of $v_{i}$ in $G$. ∎
From the above two lemmas, we can conclude the following theorem:
Theorem 5.
It is sufficient to compute $IndCount_{(u,v)}(H,G_{v}^{r}),$ for $i\in[n],$ where $G_{v}^{r}$ is induced subgraph of $r-$hop neighborhood vector of $v$.
The following theorem gives a direct comparison with the $S_{k}$ model Papp and Wattenhofer (2022a), where each node has an attribute which is the count of induced subgraphs of size at most $k.$
Theorem 6.
Local $k-$WL can count all induced subgraphs of size at most $(k+2)$.
Proof.
Suppose, if possible $G$ and $H$ are Local $k-$WL equivalent and $|P|\leq(k+2),$ where $P$ is the pattern to be counted. We will choose one vertex $v$ as a key vertex. Now, we want to count $P-v$ locally. Assume that the distance from $v$ to any vertex is ${}^{\prime}r^{\prime}$. So, we take the $r$-hop neighborhood of every vertex in $G$ and $H,$ respectively. It is easy to see that the number of induced subgraphs or subgraphs of size $k$ is the same locally if they are $k-$WL equivalent since we do an initial coloring of $k$-tuples based on isomorphism type. Now, suppose $P^{\prime}=P-v$ is of size $(k+1)$.
Let $P_{i}=P^{\prime}-v_{i},$ for $i\in[k+1].$ It is possible that there may exist two pairs of subgraphs that are isomorphic. In that case, we remove such graphs. Let $P_{1},P_{2},\ldots,P_{l}$ be all pairwise non-isomorphic graphs. $k-$WL would initially color the vertices according to isomorphism type. So, the subgraph count of $P_{i}$ is the same in any two $k-$WL equivalent graphs. Let $V(P_{i})=(u_{1},u_{2},\ldots u_{k})$. We see the refined color after one iteration in equation 3. Now, we can observe that by looking at the first coordinate of the color tuples in a multiset, we can determine the adjacency of $u$ with $(u_{2},u_{3},\ldots u_{k})$. Similarly, after seeing the second coordinate of the color tuples in the multiset, we can determine the adjacency of $u$ with $(u_{1},u_{2},\ldots u_{k})$.
Consider, $\forall{u}\in V(G),$ $P_{i}\cup\{u\}=H$ will give the count of induced subgraph $P^{\prime}$. Thus, if $G$ and $H$ are $k-$WL equivalent, then the size of each color class after the first iteration will be the same.
Now, for each $P^{\prime}$ with $v$ will form $P$ if it has exactly $|N_{P}(v)|$ many vertices of color $1$. Also, as mentioned earlier that $k-$WL is equivalent to $C^{k+1}$ logic, and we have to add unary logic stating that the color of neighbor to be $1$. The $k-$WL and $C^{k+1}$ are equivalent, so we have to add the unary relation putting the condition of the required colors. Again, using Lemma 3 we can say that running $k-$WL locally can count all patterns of size $(k+2)$ appearing as an induced subgraph. ∎
We can see the corollary below in which we mention the set of patterns shown in Chen et al. (2020).
Corollary 3.
Local $2-$WL on the subgraphs induced by neighborhood of each vertex can count each pattern appearing as induced subgraphs as well as subgraphs of (a) 3-star (b) triangle (c) tailed triangle (d) chordal cycle (e) attributed triangle.
Based on the above results, we now present the algorithm 4 for counting patterns appearing as an induced subgraph in $G$ using the localized algorithm. The function $IndCount_{u,v}(H,G_{v}^{r})$ takes as input the pattern $H$, the attributed version of $G_{v}^{r}$, and returns the induced subgraph count of $H$ in $G_{v}^{r},$ where $u\in H$ is mapped to $v\in G_{v}^{r}$. Notice that the function $IndCount_{u,v}(H,G_{v}^{r})$ is a predictor that we learn using training data.
Algorithm 4 Counting induced subgraph H in G
1:Given $G,H$.
2:Find $r=radius(H)$ and let $u\in H$ be a corresponding center vertex.
3:for each vertex $v$ in V(G) do
4:     Extract subgraph $G_{v}^{r}$.
5:     Find suitable $k$, which will give an exact count based on the local substructure.
6:     Run Local+Layer $k-$WL on $G_{v}^{r}$.
7:     Calculate $IndCount_{u,v}(H,G_{v}^{r})$.
8:end for
9:return $\frac{\sum_{v\in V(G)}IndCount_{(u,v)}(H,G_{v}^{r})}{|Orbit_{H}(u)|}$
The running time and space requirement for Algorithm 4 is dependent on the value of $k$ and $r$. We can make informed choices for the values of $k$ and $r.$ Notice that the value of $k$ is chosen based on the local substructure. Also, the value of $r$ is the radius of $H.$ Suppose the local substructure is simple (planar, bounded treewidth, bounded rankwidth Theorem 4). In that case, $k-$WL, for small values of $k,$ is sufficient for counting induced subgraph $H.$ Otherwise, we have to run $(|H|-2)-$WL in the worst case.
6.3 Deciding k based on the pattern for counting subgraphs
For any pattern $H$, it turns out that the number of subgraph isomorphisms from $H$ to a host graph $G$ is simply a linear combination of all possible graph homomorphisms from $H^{\prime}$ to $G$ ($Hom(H^{\prime},G)$ is the number of homomorphisms from $H^{\prime}$ to $G$) where $H^{\prime}$ is the set of all homomorphic images of $H$. That is, there exists constants $\alpha_{H^{\prime}}\in\mathbb{Q}$ such that:
$Count(H,G)=\sum_{H^{\prime}}\alpha_{H^{\prime}}Hom(H^{\prime},G)$ (2)
where $H^{\prime}$ ranges over all graphs in $H^{\prime}$. This equation has been used to count subgraphs by many authors (Please refer Alon et al. (1997); Curticapean et al. (2017)).
Theorem 7.
Cai et al. (1992); Dell et al. (2018) For all $k\geq 1$ and for all graphs $G$ and $H$, the following are equivalent:
  1. 1.
    $HOM(F,G)=HOM(F,H)$ for all graph $F$ such that $tw(F)\leq k$.
  2. 2.
    $k-$WL does not distinguish $G$ and $H$ and
  3. 3.
    Graph $G$ and $H$ are $C^{k+1}$ equivalent***Counting logic with (k+1) variables.
Using Equation 2 and Theorem 7, we arrive at the following theorem:
Theorem 8.
Let $G_{1}$ and $G_{2}$ be $k-$WL equivalent and $htw(H)\leq k$. Then subgraph count of $H$ in $G_{1}$ and $G_{2}$ are the same.
Lemma 5.
$Count(H,G)=\frac{\sum_{v\in V(G)}Count_{(u,v)}(H,G_{v}^{r})}{|Orbit_{H}(u)|}$ (3)
Proof.
Suppose $H$ appears as subgraph in $G$. Then, there must exist an injective function matching $u\in V(H)$ to some $v\in V(G)$. Thus, counting locally, we can easily see that every subgraph would be counted. Now to prove Equation 3, it is sufficient to show that for a given subgraph it is counted exactly $|Orbit_{H}(u)|$ many times. Note that two subgraphs are same if and only if their vertex sets and edge sets are same. We fix the vertex set and edge set in $G$, which isomorphic to $H$. Now, consider an automorphism of $H$ which maps $u$ to one of the vertices $u^{\prime}$ in its orbit. Note that we can easily find the updated isomorphic function that maps $u^{\prime}$ to $v$. Now, the number of choices of such $u^{\prime}$ is exactly $|Orbit_{H}(u)|$. Thus, the same subgraph is counted at least $|Orbit_{H}(u)|$ many times. Suppose $x\in V(H)$ is a vertex such that $x\notin Orbit_{H}(u)$. Considering the fixed vertex set and edge set, if we can find an isomorphism, then it is a contradiction to the assumption that $x\notin Orbit_{H}(u)$. Thus, the same subgraph is counted exactly $|Orbit_{H}(u)|$ many times. ∎
Using Theorem 8 and Lemma 5, one can easily see that for counting pattern $H$ as a subgraph, it is sufficient to run Local $k-$WL on the local substructure and count the subgraph locally.
Theorem 9.
Local $k-$WL can exactly count any subgraph $H$ if $htw(H-v)\leq k$.
The upper bound on the choice of $k$ for running $k-$WL can be improved from the default $|H|-2$ bound that we used for the induced subgraph count. The value of $k$ is now upper bound by $htw(H)$. Hence, we pick the minimum $k$ based on the local substructure of $G$ as well as the hereditary treewidth of pattern $H$ for computing the subgraph count of $H$ in $G$. The algorithm for counting the subgraph is similar to the induced subgraph.
Corollary 4.
Local $1-$WL can exactly count the number of patterns $P$ or $P-v$ appearing as a subgraph, when $P$ or $P-v,$ with a dominating vertex $v$, is $K_{1,s}$ and $2K_{2}.$
Proof.
In a star, $K_{1,s},$ all the leaves are mutually independent. By the definition of homomorphism, the edges are preserved in homomorphic images. So, the only possibility of a homomorphic image of the star is getting another star with less number of leaves. Note that the star is a tree, and its treewidth is one. Also, for $2K_{2}$, the homomorphic image is either itself or the star. So, the treewidth of all its homomorphic images is $1$
Corollary 5.
Local $1-$WL can exactly count the number of $C_{4}$ appearing as a subgraph.
Proof.
We can see that $H-v$ is $K_{1,2}$ when choosing any vertex as a key vertex. Also, the orbit size is $4$. So, we can directly use Lemma 5 to compute the count of the $C_{4}$ in the graph locally and then sum it over all the vertices and divide it by $4$. ∎
Corollary 6.
Local $1-$WL can exactly count patterns appearing as subgraphs of (a) 3-star (b) triangle (c) tailed triangle (d) chordal cycle (e)attributed triangle and patterns appearing as induced subgraphs of (a) triangle and (c) attributed triangle.
Proof.
For subgraph counting, we can see that for all of the $5$ patterns, there exists a vertex $v$ such that $htw(P-v)=1.$ One can note that the attributed triangle can also be handled using Corollary 1. Since every pattern has a dominating vertex, running $1-$WL on the subgraph induced on the neighborhood is sufficient. Now, we only have to argue for the patterns appearing as induced subgraphs. Note that the induced subgraph count of the triangle and the attributed triangle is same as the subgraph count of the triangle and attributed triangle. ∎
Note that all the subgraph or induced subgraph counting can be easily extended to attributed subgraph or attributed-induced subgraph counting (graph motif). We will be given a coloured graph as an input, and we will incorporate those colours and apply a similar technique as described above to get the subgraph count.
Corollary 7.
If $C(G)=C(H),$ where $C(.)$ is the color histogram, then $Count(P,G)=Count(P,H)$ where $P$ is the attributed subgraph.
In particular, for counting the number of triangles, we can see that it is enough to count the number of edges in the subgraph induced on the neighbourhood of the vertices. Thus, Local $1-$WL can give the exact count of the number of triangles. For more details, please see 6.4. The running time of $1-$WL depends on the number of iterations, $h$. In general, it takes $O((n+m)\log n)$ time, where $m$ is the number of edges, whereas when we look at it in terms of iteration number it requires $O(nh)$ time.
Lemma 6.
It requires $O(n)$ time to guarantee the count of patterns which can be written using $2$-variable with a counting quantifier where the depth of the quantifier is constant.
For more details, please see the equivalence between the number of iterations and quantifier depth in Theorem 2. A list of patterns that can be counted as subgraph and induced subgraphs using local $1-$WL and local $2-$WL are mentioned in Table 1. Also, the patterns including 3-star, triangle, chordal $4$-cycle (chordal $C_{4}$), and attributed triangle have been studied in Chen et al. (2020) and have been shown that it cannot be counted by $1-$WL.
Table 1: List of all the patterns that can be counted exactly(as a subgraph or induced subgraph), given $G$, using Local $k-$WL, for different $k.$
Restriction on $G_{v}^{r}$ k Patterns, $H$
Induced
subgraph
Subgraph Reference
$G_{v}^{r}$ is amenable 1
All
Corollary 1
Max degree $\leq$ 5 1
Patterns with a dominating vertex
Corollary 2
Max degree $\leq$ 15 2
Pattern with a dominating vertex
Corollary 2
No restriction 2
3-star, triangle, tailed triangle, chordal cycle,
attributed triangle
Corollary 3
No restriction 1
Either $H$ or $H-v$ is $K_{1,s}$ or $2K_{2},$ where $v$ is the
dominating vertex
Corollary 4
No restriction 1 $C_{4}$ Corollary 5
No restriction 1
3-star, tailed triangle, chordal cycle
Corollary 6
No restriction 1
triangle, attributed triangle
Corollary 6
6.4 Algorithms for subgraph counting
Triangle counting in the host graph
We describe an algorithm for counting the number of triangles in a given host graph $G$ in Algorithm 5. Note that counting the number of triangles is the same as counting the number of edges in the subgraph induced by $N_{G}(v).$ It is well known that two $1-$WL equivalent graphs have the same number of edges. This ensures that if we run $1-$WL on the induced subgraphs in the neighborhood of $v$, taking color as a feature, we can guarantee the count of the triangles. On the other hand, we can see that running $1$- WL on graph $G$ will not guarantee the count of triangles. Running $1-$WL on the entire graph takes $O(n+m)\log(n)$ and $O(n)$ space, where $m$ is the number of edges. Thus, running $1-$WL locally in the neighborhood is more space and time efficient. Note that the running time is also dependent on the number of iterations, $h.$ Running $1-$WL for $h-$ iteration requires $O(nh)$ time. The quantifier depth of counting logic with $(k+1)$ variables is equivalent to the number of iterations of $k-$WL ( See Theorem 2). For the case of triangle counting, we just need to count the number of edges, which can be done by running just one iteration of $1-$WL. So, the time required is $O(deg(v))$ for each $v.$ This can be done in parallel for each vertex.
Algorithm 5 Counting the number of triangles
1:Let $G$ be the host graph.
2:$num\_edges=0$
3:for each vertex $v$ in $V(G)$ do
4:     Find the induced subgraph on $N_{G}(v)$
5:     Find the number of edges in the induced subgraph on $N_{G}(v)$
6:     Add it to $num\_edges$
7:end for
8:$c=num\_edges/3$
9:Output: The number of triangles in graph $G$ is $c$.
Counting subgraph of radius one
We begin by explaining the procedure for counting the number of subgraphs having a dominating vertex (radius one). For this purpose, we fix a dominating vertex $u$. If a subgraph exists, then the dominating vertex must be mapped to some vertex. We iteratively map the dominating vertex to each vertex in the host graph and count the number of patterns in the neighborhood of the dominating vertex.
Here we present an algorithm for counting patterns of radius one Algorithm 6. Note that running $k-$WL on the entire graph takes $O(k^{2}\cdot n^{k+1}\log n)$ time and $O(n^{k})$ space, whereas when we run locally, it requires less time and space. Suppose we run only on the neighborhood of each vertex. Then, it requires $\sum_{v\in V(G)}(deg(v))^{k+1}\log(deg(v))$ and space $O(max_{i}(deg(v_{i}))^{k}+n).$ More specifically, suppose the given graph is $r$-regular. Then it requires $O(r^{k+1}\log(r)n)$ time and $O(r^{k}+n)$ space. Therefore, if the graph is sparse, then we can implement Local $k-$WL for a larger value of $k$. We can see that running $k-$WL locally is not dependent on the size of the graph exponentially. However, it is not feasible to run $k-$WL on the entire graph for a larger value of $k.$
Algorithm 6 Counting the number of patterns of radius one
1:Let $H$ be a pattern having a dominating vertex and $G$ be the host graph.
2:for each vertex $v$ in $V(G)$ do
3:     Find the induced subgraph on $N_{G}(v)$
4:     if degree$(v)+1<|V(H)|$ then
5:         skip this iteration
6:     end if
7:     run $k-$WL on the induced subgraph on $N_{G}(v)$
8:     Calculate $Count_{u,v}(H,G_{v}^{r})$
9:end for
10:return $\frac{\sum_{v\in V(G)}Count_{u,v}(H,G_{v}^{r})}{|Orbit_{H}(u)|}$
Counting subgraphs of radius r
Here, in Algorithm 7, we describe how to count the number of subgraphs of radius $r.$ We iterate over all the vertices and take the $r$-hop neighborhood around that vertex, say $v,$ and choose a suitable $k$ according to the structure of the pattern that can guarantee the count of subgraphs in the local substructure.
Algorithm 7 Counting the number of subgraphs
1:Let $P$ be a pattern having a key vertex $u\in V(P)$ and $G$ be the host graph.
2:for each vertex $v$ in $V(G)$ do
3:     Run BFS on $G$ rooted at $v$
4:     $c=1$
5:     for each layer of the BFS tree rooted at $v$ do
6:         if ##\## vertices at layer $i$ from $v$ $<$ ##\## vertices at layer $i$ from $u$ then
7:              $c=0$
8:              skip this iteration
9:         else
10:              Color each vertex with color $i$ for layer $i$
11:              This gives a colored graph at each layer
12:         end if
13:     end for
14:     if $c==0$ then
15:         Skip this iteration
16:         $c=1$
17:     end if
18:     Find an induced subgraph on the set of vertices in the $r$-hop neighborhood
19:     Run $k-$WL on the induced subgraph
20:     Calculate $Count_{u,v}(H,G_{v}^{r})$
21:end for
22:return $\frac{\sum_{v\in V(G)}Count_{u,v}(H,G_{v}^{r})}{|Orbit_{H}(u)|}$
7 Fragmentation
Here, we discuss the Fragmentation technique that is different from the localized methods we have seen so far. From Table 1, we have seen that Local $k-$WL (or ((Local + Layer) $k-$WL)) is sufficient for getting an exact count for the patterns given in the table. Given a pattern $P,$ that is more complicated than the patterns in Table 1, we fragment $P$ into simpler patterns such that their exact count is known. In the subgraph GNN proposed earlier, look into subgraph of the host graph. We have seen that this technique is scalable on large graphs. Also, we have seen that subgraph GNN is more expressive and efficient than traditional GNN. So, we tried to explore the expressibility when the pattern is also fragmented into smaller subpatterns.
The fragmentation method involves fragmenting the pattern, $P$, into smaller subpatterns and counting these subpatterns to get the count of the $P$ in the host graph. As described in Section 6, the value of $k$ depends on the size of the pattern (induced subgraph) and its structure (subgraph count). Thus, even though the $htw(P)$ may be large, if we divide it into subpatterns, then $k$ is required to guarantee the count would be reduced. Thus, it provides more expressiveness for smaller $k$ in order to count the patterns that cannot be counted if we directly apply Local $k-$WL. Thus, given a pattern, we apply the same fragmentation on $G_{v}^{r}.$ Thus, the number of occurrences of $H$ in $G_{v}^{r}$ can be computed by combining the counts of the simpler patterns. Instead of training a GNN for counting $H,$ we can design GNNs for learning the easier tasks (i.e., for counting the simpler models) and combine the outputs of those models. It should be noted that the fragmentation into smaller subgraphs depends on the structure of the pattern $H.$
We demonstrate this technique for counting induced tailed triangles in Figure 2. As seen in the figure, the tailed triangle pattern can be fragmented into two parts : the pendant vertex as the key vertex, and an attributed triangle. The colors assigned to the nodes of the attributed triangle depend on the distance of the nodes from the key node. Thus, the task of counting tailed triangles reduces to counting attributed triangles, as all the vertices at level 1 are connected to the root.
Suppose the task is to count the number of chordal cycles appearing as induced subgraphs. If we pick the vertex of degree three as the key vertex, then it is enough to search the neighbourhood of $v$ in the host graph. Now, in $N_{G}(v),$ if we count the number of $2-stars,$ then it gives the count of chordal cycle appearing as subgraph. If we eliminate the appearance of $K_{4}$, then it would give the exact count of chordal cycles appearing as induced subgraphs. In that case, we count the number of triangles in $N_{G}(v),$ which gives the exact count of $K_{4}$.
Using the fragmentation technique, we show that just $1-$WL is sufficient for getting exact counts of induced subgraphs of certain sizes.
Figure 2: Fragmentation for counting tailed triangles.
Idea of the fragmentation algorithm 8: Given a graph $G$ and a pattern $P$, we first fix a vertex $u\in V(P)$ as the key vertex. Now, assume that the radius of the pattern is $r$. Thus, for counting $P$ locally, it is sufficient to take the $r$-hop neighbourhood for each vertex $v$ of $G,$ say $G_{v}^{r},$ as has been shown in Lemma 3 and Lemma 5. Also, we have proved above that doing count locally is sufficient for both subgraph and induced subgraph. Now, we fragment pattern $P$ into smaller subpatterns, say $P_{1},P_{2},\ldots P_{l}$. Based on the structure of $P$, we consider the subgraphs of $G_{v}^{r}$ where the subpattern $P_{i}$ is required to be counted. For all subpattern $P_{i}$ in $P$, we make a list of subgraphs $G_{v}^{r}(1),G_{v}^{r}(2),\ldots G_{v}^{r}(t)$ of $G_{v}^{r}$ where $P_{i}$ needs to be counted. We aggregate these lists into a dictionary, $\mathcal{L},$ with $P_{i}$ as the keys. It should be noted that the decomposition of $P$ into $P_{i}$s is such that their counts can be calculated. That is, we have learnt models $M_{i},$ corresponding to each $P_{i}$, which counts the number of subpatterns $P_{i}$. The array $c$ stores the count of $P_{i}$’s in each subgraph of $G_{v}^{r}$. Now, for each vertex, we use appropriate functions $\alpha$ and $\beta$ to combine the counts in $c$ to get the count of $P$ in $G_{v}^{r}$. Finally, the function $\gamma$ finds the normalizing factor to get the actual count of the pattern in the $G.$
Algorithm 8 Fragmentation Algorithm
1:Let $G$ be the host graph and $P$ be the list of patterns,
$\mathcal{L}$ be the dictionary of subgraph rules associated with the subpattern $P_{i},$
3:$M=\{M_{1},\ldots,M_{l}\}$ the list of learnt models for counting $P_{i}$’s, where $l=|P|$
4:$a\leftarrow$ Zero array of size $|V(G)|$
5:for each vertex $v$ in $V(G)$ do
6:     Extract $G_{v}^{r}$
7:     $b\leftarrow$ Zero array of size $l$
8:     for each pattern $P_{i}$ in $\mathcal{L}$ do
9:         $c\leftarrow$ Zero array of size $s$, where $s=|\mathcal{L}(P_{i})|$
10:         for each rule $k$ in $\mathcal{L}(P_{i})$ do
11:              Extract $G_{v}^{r}(k)$
12:              $c[k]$ = $M_{i}(G_{v}^{r}(k))$
13:         end for
14:         $b[i]$ = $\alpha(c)$
15:     end for
16:     $a[v]$ = $\beta(b)$
17:end for
18:$Count(P,G)$ = $\gamma(a)$
19:return $Count(P,G)$
Theorem 10.
Using the fragmentation method, we can count induced tailed triangle, chordal $C_{4}$ and $3-star,$ by running Local $1-$WL.
Proof.
For tailed triangle, we fix the pendant vertex as the key vertex (refer to Figure 2). Now, we have to look for the count of triangles such that exactly one of the node of triangle is adjacent to the key vertex. We find induced subgraph of $2$-hop neighborhood for each key vertex. Now, we color the vertices at distance $i$ by color $i$. Then, the problem of counting the number of tailed triangles reduces to counting the number of colored triangles in the colored graph such that one node of triangle is colored $1$ and remaining two nodes are colored $2$. We can find the count of colored triangles using $1-$WL on the induced subgraph by Corollary 1. This number of count of colored triangle is same as $IndCount_{(u,v}(tailed-triangle,G_{v}^{r})$. Now, using Lemma 3, we can say that fragmentation technique can count tailed triangles appearing as induced subgraphs using $1-$WL.
Consider the pattern, chordal $C_{4}.$ We have proved that $1-$WL can count the number of subgraphs of chordal $C_{4}.$ So, to count the number of induced subgraphs of chordal $C_{4},$ we only have to eliminate the count of $K_{4}.$ When we fix one vertex of degree $3$ as key vertex, we can easily compute the count of $K_{1,2}$ in the neighborhood. Now, we have to eliminate all three tuples appearing as triangles in the neighborhood of the key vertex. We can easily count the number of triangles in the neighborhood of each vertex. This gives the exact count of chordal $C_{4}$ appearing as subgraph in the local structure. Using Lemma 3, we can find $IndCount(ChordalC_{4},G)$.
Consider the pattern, $3-star$. Here, we choose any of the pendant vertex as key vertex. Now, we have to compute $K_{1,2},$ where one center vertex of the star is connected to the key vertex. We can easily count the number of colored $K_{1,2}$ in $2$-hop neighborhood of the key vertex. However, a triangle can be also be included in this count. So, we have to eliminate the $3$ tuples forming a triangle. Again, using the approach discussed above, we can count the number of colored triangles and this will output the exact count of colored induced $K_{1,2}$. Again, using lemma 3, we can find $IndCount(3-star,G).$
Theorem 11.
Using the fragmentation technique we can count all patterns appearing as induced subgraphs of size $4$ by just running Local $1-$WL.
Proof.
In Table 2, we describe how the fragmentation technique can be leveraged to count all the induced subgraphs of size $4$. This shows that for $k=1$, the fragmentation technique is more expressive than $S_{k+3}$. Also, we can enlist more graphs where the larger pattern can be seen as the union of fragments of smaller patterns. Using this, we can see that it covers all the graphs that were mentioned in Chen et al. (2020). One can see that all the formulae use the function that can be computed by $1-$WL. The number of vertices and the number of edges can easily be computed after doing one iteration of $1-$WL. Also, the degree sequence can be computed after running $1-$WL for one iteration. All other functions(formulae) are just the linear combination of the functions computed earlier, the number of vertices or the number of edges. ∎
In the structures shown in the table 2 below, we have highlighted the key vertex by "light green" color and other vertices by "black color".
Table 2: Fragmentation for patterns for size of at most 4.
Vertices Structure Formula
G1 [Uncaptioned image] $\binom{n}{2}-|E|$
G2 [Uncaptioned image] $|E|$
G3 [Uncaptioned image] $\large\frac{\sum_{v}IndCount_{(u,v)}(G_{1},G-G[N_{G}[v]])}{3}$
G4 [Uncaptioned image] $\sum_{v}IndCount(G_{2},G-G[N_{G}[v]])$
G5 [Uncaptioned image] $\frac{\sum_{v}IndCount_{(u,v)}(G_{2},G[N_{G}(v)])}{3}$
G6 [Uncaptioned image] $\sum_{v}\binom{degree(v)}{2}-|E(N_{G}(v))|$
G7 [Uncaptioned image] $\frac{\sum_{v}IndCount_{(u,v)}(G_{3},G-G[N_{G}(v)]}{4}$
G8 [Uncaptioned image] $\frac{\sum_{v}IndCount_{(u,v)}(G_{5},G[N_{G}[v]])}{4}$
G9 [Uncaptioned image] $\frac{\sum_{v}IndCount_{(u,v)}(G_{4},G-G[N_{G}[v]])}{2}$
G10 [Uncaptioned image] $\frac{\sum_{v}(IndCount(G_{6},N_{G}(V))-IndCount(G_{5},N_{G}(v)))}{2}$
G11 [Uncaptioned image] $\sum_{v}(IndCount(G_{6},G-G[N_{G}[v]])$
G12 [Uncaptioned image] $\sum_{v}Count(attribute*-triangle,G[N_{G}(v)\cup N_{G}(N_{G}(v))])$
G13 [Uncaptioned image] $\binom{|E(G)|}{2}-Count(k_{1,2},G)$
G14 [Uncaptioned image] $\frac{\sum_{v}IndCount(attributed-k_{1,2},G[N_{G}(v)\cup N_{G}(N_{G}(v))])}{4}$
G15 [Uncaptioned image] $\sum_{v}IndCount(attribute-k_{1,2},G[N_{G}(v)\cup N_{G}(N_{G}(v))])$
G16 [Uncaptioned image] $\frac{\sum_{v}(n-deg(v)-1)\cdot|E(G[N_{g}(v)]|}{3}$
G17 [Uncaptioned image] $\frac{IndCount(attributed-G_{4},G[N_{G}(v)\cup N_{G}(N_{G}(v))]}{2}$
8 Theoretical Comparison of Graph Neural Networks
In this section we give a comparison of the time and expressiveness between the GNN hierarchies proposed in Papp and Wattenhofer (2022a) and our methods. From Theorem 3, it is clear that $k-$WL is less expressive than Local $k-$WL. Also, we have shown that the space and time required by Local $k-$WL are less compared to $k-$WL.
$S_{k}$ is a model in which a vertex is given an attribute based on the number of connected induced subgraphs of size at most $k,$ the key vertex. Even though it searches locally, the number of non-isomorphic graphs may be too many for a small value of $k$. Suppose the radius of the induced subgraph is $r$; then it has to search up to the $r$-hop neighborhood. Using brute force would require $O(n_{1}^{k})$-time to count the number of induced subgraphs of size $k,$ for each individual induced subgraph. To improve the time complexity, it either stores the previous computation, which requires a lot of space or further recomputes from scratch. Thus, it would require $O(t_{k}\times n_{1}^{k})$, where $t_{k}$ is the number of distinct graphs upto isomorphism of at most $k$-vertices. Using Theorem 6, one can easily see that running $k-$WL locally is more expressive than $S_{k+2}.$
The $N_{k}$ model has a preprocessing step in which it takes the $k$-hop neighborhood around vertex $v,$ and gives attributes based on the isomorphism type. In a dense graph, it is not feasible to solve the isomorphism problem in general, as the size of the induced subgraph may be some function of $n$. Also, using the best known algorithm for graph isomorphism by Babai (2016), the time required is $O(n_{1}^{O(\log n_{1})}).$ However, running Local $k-$WL would require $O(n_{1}^{k})$. Also, there are rare examples of graphs that are $3-$WL equivalent and non-isomorphic. So, if we run $3-$WL locally, then most of times expressive power matches with $N_{k}.$
The $M_{k}$ model deletes a vertex $v$ and then runs $1-WL.$ Papp and Wattenhofer (2022a) proposed that instead of deleting the vertices, $k$ vertices are marked in the local neighborhood and showed that it is more expressive than deletion. It identifies $k$ set of vertices in the local $r$-hop neighborhood of the graph. It would require $O(n_{1}^{(k+2)}\log(n_{1}))$ time as it has $O(n_{1}^{k})$ many possibilities of choosing $k$ many vertices. It requires $O(n^{2}\log n)$ time to run $1-$WL. The same time is required for Local $(k+1)-$WL. That’s why we compare $M_{k-1}$ with Local $k-$WL in Table 3. Also, it is known that with any $l$ markings and running $k-$WL is less expressive than running $(k+l)-$WL on the graph Fürer (2017). So, if we plug in the value, we can see that running Local $k-$WL is more expressive than doing $l$ marking and running $1-$WL. One can get an intuition by comparing with the $(k+1)$ bijective pebble game. If we mark the vertices, then the marking pebbles get fixed and give less power to spoiler. However, just running $k-$WL the spoiler is free to move all the pebbles. We present a simple proof that Local $k-$WL is at least as expressive as $M_{k-1}$.
Theorem 12.
Local $k-$WL is atleast as expressive as $M_{k-1}$.
Proof.
Let $G_{v}^{r}$ and $G_{u}^{r}$ be the induced subgraphs around the $r$-hop neighborhood for vertices $v$ and $u$, respectively. Let $M_{k}$ distinguish $G_{v}^{r}$ and $G_{u}^{r}$. We claim that Local $k-$WL can also distinguish the graph $G_{v}^{r}$ and $G_{u}^{r}$. To prove our claim, we use the pebble bijective game. $G_{v}^{r}$ is distinguished because there exists a tuple $(v_{1},v_{2},....v_{k-1})$ such that marking these vertices and running $1-$WL on the graph gives stabilised coloring of the vertices that does not match with that of $G_{u}^{r}.$ Now, consider two games. One game corresponds to $1-$WL and another to Local $k-$WL. For the first $(k-1)$ moves, the Spoiler chooses to play and places pebbles at $(v_{1},v_{2},....v_{k-1}).$ After that, in both games, there are two pebbles and the position of both the games are same. Let $S_{1}$ and $D_{1}$ be the spoiler and Duplicator in the $(k+1)$ bijective pebble game, and $S_{2}$ and $D_{2}$ be the spoiler and Duplicator in $2$ bijective pebble game. $S_{1}$ will follow the strategy of $S_{2}$ and $D_{2}$ follows the strategy of $D_{1}.$ We prove by induction on the number of rounds. Our induction hypothesis is that the position of games in both the games is the same and if $S_{2}$ wins, then $S_{1}$ also wins.
Base case : Duplicator $D_{1}$ proposes a bijection. $D_{2}$ will propose the same bijection. Now, $S_{2}$ places a pebble on some vertex $v.$ Similarly, $S_{1}$ will also place a pebble at $v.$ Note that the position of the game is the same, and if $S_{2}$ wins, then $S_{1}$ also wins.
Now, using the induction hypothesis, assume that the position of both the games is the same and $S_{2}$ has not won till round $i.$
Now, consider the game at round $(i+1).$ If $S_{2}$ decides to play / remove a pebble, then $S_{1}$ will do the same. If $S_{2}$ decides to play a pebble, then $S_{1}$ also decides to play a pebble. So, $D_{1}$ proposes a bijective function. $D_{2}$ proposes the same bijective function. Now, $S_{2}$ places pebble at some vertex $u,$ then $S_{1}$ also places pebble at $u$. Thus, the position of both the game is the same and if $S_{2}$ wins, then $S_{1}$ will also win. ∎
Table 3: Here, $n:$ number of nodes in the graph, $n_{1}:$ max number of nodes in a $r$-hop neighbourhood of a vertex, $n_{2}:$ maximum number of nodes in any two consecutive layers for a particular vertex, over all the vertices of the graph, and $t_{k}:$ number of distinct graphs upto isomorphism of at most $k$-vertices. The first row compares the expressiveness of the models , and the second row compares the time complexity, with respect t0 our models.
Local k -WL (Local+Layer) k -WL k- WL S_k M_k-1
Expressiveness - - Less
Less than
Local $(k-2)-$WL
Less
Time $O(n\times n^{k+1}_{1})$ $O((n_{2})^{k+1}\times rn)$ $O(n^{k+1})$ $O(t_{k}n^{k}_{1}n)$ $O(n^{k+1}_{1}n)$
9 Proposed Architecture
In this section, we propose two GNN architectures for subgraph as well as induced subgraph counting. The design of these networks follows from the discussions about subgraph and induced subgraph counting in the previous sections. Given a graph $G$ as an input, we extract the rooted subgraphs $G_{v}^{r}$ for each vertex $v.$ Each of these subgraphs is rooted at $v.$ The value of $r$ depends on the structure to be counted. For example, we extract $1-$hop induced subgraphs for counting triangles. These subgraphs are then given as inputs to the proposed models.
9.1 InsideOut GNN
We conducted the experiments using two different architectures, InsideOutGNN and InSigGNN.
Figure 3: Schematic of the InsideOutGNN Model
The InsideOut GNN model (Figure 3) consists of an internal network and an external network. The internal network learns the local substructure counts for each rooted subgraph, $G_{v}^{r}.$ A subgraph are given as input pass through an Internal GNN model. This model consists of GINConv Xu et al. (2019) layers. The Internal GNN outputs the embeddings of the nodes present in a given subgraph. These node embeddings are then aggregated using a GlobalAddPool layer to get the embedding of the root node $v$ of the subgraph. This root node embedding is the feature vector for $v,$ as given in line $7$ of Algorithm 1. The GINConv layers simulate the WL iterations on the nodes of the subgraph. This root node embedding, $emb_{v},$ is then passed through a linear transformation layer to learn the local counts of the required substructure in the subgraph.
Now, we take a union of all the subgraphs for a given graph. Using the node embeddings, $emb_{v},$ learnt from the Internal GNN model, the External GNN model learns the global count of the required substructure. The External GNN model also comprises of GINConv layers.
The intuition behind this design is to make the models learn the local counts well. Once these local counts are learned, the global count is just a linear function of the local counts, which can be learnt easily by the External GNN model.
9.2 InSigGNN
For the InSig model, the Internal GNN component is similar to that of the InsideOut GNN model. The architecture is shown in Figure 4. However, in this model, we do not transfer the node embeddings learnt from the Internal GNN part of the model. Instead, we add all the local counts predicted by the internal part of the model and pass the sum as an input to the external model, which is just a linear transformation layer. The weights learned in the Internal GNN model depend only on the local counts of the substructures in the rooted subgraphs. The external linear transformation is learned based on the global count of a substructure in the entire input graph.
Figure 4: Schematic of the InSigGNN model
9.3 Implementation Details
We experimented with both InsideOutGNN and InsigGNN models to predict the count of structures such as triangles, 3stars, 2stars, and Chordal cycles. For each of these structures, the true local counts are computed using various methods. For example, if our task is to count the number of triangles in a particular subgraph, for the root node of the subgraph, we label its local count to be the number of edges present in a 1-hop neighbouring subgraph without the root node. Using these local counts, from Equation Lemma 5, we assign the global counts to be $\frac{1}{3}\sum\limits_{u\in V(G_{v}^{1})}\prescript{x\mkern-0.5mu}{}{C}_{2}$, where $x$ is the local count corresponding to the root vertex $v.$ Here, $3$ is the orbit of a node in a triangle. Similarly, we assign the local counts and the global counts for different substructures.
When counting larger substructures such as $K_{4}$ and $C_{4},$, we use the fragmentation technique with the help of the InSignGNN model. For calculating $K_{4}$, we construct a set of subgraphs. For every node in the graph, a 1-hop neighbouring subgraph is generated, from which we remove the root node. In this subgraph, we count the number of triangles and process accordingly. Therefore, we use the model already trained for counting triangles in order to count $K_{4}.$ During the inference, we apply a rounding function, as the counts are integer numbers.
9.4 Hyperparameters
We use two GIN Convolutional layers for the Internal GNN part in both proposed models. For the models we have considered, as there can be paths of lengths more than 1 in the subgraph, $2$ GINConv layers are sufficient to capture the information well. In the InsideOut model, we use $2$ GINConv layers in the external part of the model as well.
We use a learning rate of $41e-$ 4 and a batch size of $1$ We also experimented with different hidden dimensions for the node embeddings and obtained the best results when we used $4$ as a hidden dimension size. The ablation studies have been mentioned in Section 10.2. The experiments were conducted using an NVIDIA A100 40GB GPU.
10 Experimental Observations
10.1 Results
Without Fragmentation Fragmentation
Models Triangle 3-star 2-star Chordal $K_{4}$ $C_{4}$ Tailed triangle
ID-GNN 6.00E-04 NA NA 4.54E-02 2.60E-03 2.20E-03 1.05E-01
NGNN 3.00E-04 NA NA 3.92E-02 4.50E-03 1.30E-03 1.04E-01
GNNAK+ 4.00E-04 1.50E-02 NA 1.12E-02 4.90E-03 4.10E-03 4.30E-03
PPGN 3.00E-04 NA NA 1.50E-03 1.65E-01 9.00E-04 2.60E-03
I2-GNN 3.00E-04 NA NA 1.00E-03 3.00E-04 1.60E-03 1.10E-03
InsideOut 0.00E+00 0.00E+00 0.00E+00 3.82E-03
InSig 0.00E+00 0.00E+00 0.00E+00 0.00E+00 0.00E+00 0.00E+00 0.00E+00
Table 4: MAE for the subgraph count of different patterns. Some results, such as $2-$star, $3-$star, are not conducted by the given baselines; therefore, it is mentioned NA. As mentioned in Section Section 10.1, InSig Model has a lesser number of parameters while giving the same performance as InsideOut Model; therefore, we used only InSig Model for fragmentation based counting.
We conducted our experiments***The dataset and the code are present in the anonymous repository[Link]. for counting the following patterns: 1. triangles, 2. $3$stars, 3. $2$stars, 4. chordal cycles, 5. $K4,$ 6. $C4$ and 7. tailed triangles. The dataset used for the experiments is a random graphs dataset prepared in Chen et al. (2020). We report our experiments’ Mean Absolute Error (MAE) in Table 4. We compare our results with those reported in ID-GNN You et al. (2021), NGNN Zhang and Li (2021), GNNAK+ Zhao et al. (2022), PPGN Maron et al. (2019) and I2-GNN Huang et al. (2023). We can observe that our approach of predicting the local counts first and then predicting the global counts, achieves zero error for almost all of the tests. Although we experimented with two models, we realise that InSig model needs much fewer parameters to learn yet gives the same results. For the structures where we used fragmentation, we apply the models for counting triangles and 2stars in order to count $K_{4}$ and $C_{4}$, respectively. Also, we use the InSig model to predict the degree of a node and the number of edges in the induced subgraph without the root node. We call them the Local Node and Local Edge models, respectively. Using the Local Node and Local Edge models, we count the number of Tailed Triangles.
For all the patterns, we observed that the model gets to zero error after only 2 to 3 epochs.
Models
Number of
Parameters
Inference Time
Memory Usage
ID-GNN 102k 5.73 2.35
NGNN 127k 6.03 2.34
GNNAK+ 251k 16.07 2.35
PPGN 96k 35.33 2.3
I2-GNN 143k 20.62 3.59
InSig 190 20 1.2
Table 5: The table shows the comparison of the number of parameters required by the baselines and InSig Model. The values shown here correspond to the triangle counting task with the hidden dimension set as $4$. It can be observed that our approach requires a considerably smaller number of parameters and beats all the baselines. Our inference time comprises model inference as well as the graph preprocessing time, where it creates the set of subgraphs corresponding to each node in the graph.
10.2 Ablation Study
We varied the hidden dimension of the GINConv layer by taking the hidden dimensions from the following set $\{2,4,8,16,32,64\}$. The loss corresponding to different hidden dimensions is plotted in Figure 5. The experiment was conducted using the InSig Model for counting triangles. We observe that our model was able to compute exact counts with just 4 hidden dimensions.
Figure 5: Ablation Study - Validation Loss VS Hidden Dimension
We also varied the number of GINConv Layers. The observations are present in Table 6. The task for this ablation study was counting triangles. As per the previous study, we set the hidden dimension as $4$. We observe that we two GINConv Layers we get the best result. It should also be noted that with other sets of layers, the loss is better than the baselines.
Number
of
Layers
Number of
Internal
Model
Parameters
Number of
External
Model
Parameters
Total
Parameters
Validation
Loss
Test
Loss
1 53 97 150 1.41E-04 4.24E-04
2 93 97 190 0.00E+00 0.00E+00
3 133 97 230 2.35E-05 9.42E-05
Table 6: This table shows the ablation study with the number of GINConv Layers in the proposed architecture.
The fragmentation technique helps us to use the previously learned models, such as the model to count triangles to count the number of $K_{4}$ in a graph. Although the loss corresponding to that experiment is zero, as mentioned in Table 4, we did another study where we used a different dataset to train previous models of fragmentation. For example, we trained the InSig model to count triangle and 2stars on $5000$ Erdos-Renyi random graphs Chen et al. (2020). Using the respective trained models, we predicted $K_{4}$ and $C_{4}$ counts in the random graphs dataset prepared in Chen et al. (2020). The reverse study was also conducted. In both cases, the loss was zero. This shows that our proposed architecture is not dependent on the statistics of the dataset. The fragmentation shows that a trained model to count the number of smaller patterns on any dataset can be used to count the larger patterns.
11 Conclusion
In this work, we progressed toward a more precise characterization of the localized versions of WL algorithms. We showed how the Local-$k-$WL lies between the global $k$ and $k+1-$WL in terms of expressiveness. We also developed strategies to make the Local-$k-$WL algorithm more efficient by introducing techniques such as layered WL, recursive WL, and pattern fragmentation. The hope is that such generalizations of the WL algorithm will lead to a finer subdivision of the WL hierarchy as well as more efficient and expressive graph neural networks.

### v5

@@
-Improving Expressivity of Graph Neural Networks using Localization
+DiffCoALG: Differentiable Learning of Combinatorial Algorithms
 Abstract
-In this paper, we propose localized versions of Weisfeiler-Leman (WL) algorithms in an effort to both increase the expressivity, as well as decrease the computational overhead.
-We focus on the specific problem of subgraph counting and give localized versions of $k-$WL for any $k$.
-We analyze the power of Local $k-$WL and prove that it is more expressive than $k-$WL and at most as expressive as $(k+1)-$WL.
-We give a characterization of patterns whose count as a subgraph and induced subgraph are invariant if two graphs are Local $k-$WL equivalent.
-We also introduce two variants of $k-$WL: Layer $k-$WL and recursive $k-$WL.
-These methods are more time and space efficient than applying $k-$WL on the whole graph.
-We also propose a fragmentation technique that guarantees the exact count of all induced subgraphs of size at most 4 using just $1-$WL.
-The same idea can be extended further for larger patterns using $k>1$.
-We also compare the expressive power of Local $k-$WL with other GNN hierarchies and show that given a bound on the time complexity, our methods are more expressive than the ones mentioned in Papp and Wattenhofer (2022a).
-11footnotetext: Indian Institute of Technology, Gandhinagar, India**footnotetext: Equal Contribution
+Subgraph counting is a fundamental task for analyzing structural patterns in graph-structured data, with important applications in domains such as computational biology and social network analysis, where recurring motifs reveal functional and organizational properties.
+In this paper, we propose localized versions of the Weisfeiler-Leman (WL) algorithms to improve both expressivity and computational efficiency for this task.
+We introduce Local $k$-WL, which we prove to be more expressive than $k$-WL and at most as expressive as $(k+1)$-WL, and provide a characterization of patterns whose subgraph and induced subgraph counts are invariant under Local $k$-WL equivalence.
+To enhance scalability, we present two variants—Layer $k$-WL and Recursive $k$-WL—that achieve greater time and space efficiency compared to applying $k$-WL on the entire graph.
+Additionally, we propose a novel fragmentation technique that decomposes complex subgraphs into simpler subpatterns, enabling the exact count of all induced subgraphs of size at most $4$ using only $1$-WL, with extensions possible for larger patterns when $k>1$.
+Building on these ideas, we develop a three-stage differentiable learning framework that combines subpattern counts to compute counts of more complex motifs, bridging combinatorial algorithm design with machine learning approaches.
+We also compare the expressive power of Local $k$-WL with existing GNN hierarchies and demonstrate that, under bounded time complexity, our methods are more expressive than prior approaches.
 1 Introduction
-Graphs have been used for representing relational and structural data that appear in a variety of domains, ranging from social network analysis and combinatorial optimization to particle physics and protein folding Dill et al.
-(2008).
-In order to learn representations of these data for various downstream learning tasks such as graph classification, graph neural networks (GNNs) have emerged as very effective models.
-Given the various types of GNN-based models developed in recent years Kipf and Welling (2017); Veličković et al.
-(2018); Hamilton et al.
-(2018); Xu et al.
-(2019), researchers have attempted to characterize the expressive power of these models.
-Morris et al.
-(2019) showed the equivalence between message-passing GNNs and 1-Weisfeiler-Leman (WL) algorithm Weisfeiler and Leman (1968), which is a well known combinatorial technique for checking graph isomorphism and similarity.
-They also showed the equivalence between $k$-GNNs and $k-$WL.
-Thus, in this paper, by $k-$WL, we will be referring to the equivalent $k$-GNN model.
-In general, the expressiveness of $k-$WL, for any $k,$ is measured by its ability to identify non-isomorphic graphs and subgraphs.
-In this paper, we are using Folklore WL and state the results accordingly.
-It has been shown that $(k+1)-$WL is more expressive than $k-$WL.
-The time and space complexity increases exponentially with $k$.
-Thus, it is infeasible to run $k-$WL on large graphs.
-Also, the $k-$WL hierarchy is crude as $3-$WL identifies almost all non-isomorphic graphs.
-Arvind et al.
-(2017) characterized the graphs that can be identified by $1-$WL.
-So, we are interested in coming up with a GNN hierarchy that can be easily extended without much computational overhead.
-More specifically, we want to define a GNN hierarchy whose expressiveness lies between $k-$WL and $(k+1)-$WL.
-A count of specific patterns is very useful in determining the similarity between two graphs.
-However, detecting and counting the number of subgraphs is generally NP-complete as it is a generalization of the clique problem.
-There have been various works on efficient algorithms for some fixed patterns and restricted host graph classes Bressan (2018); Shervashidze et al.
-(2009); Bouritsas et al.
-(2020); Zhao et al.
-(2022); Shervashidze et al.
-(2011); Komarath et al.
-(2023); Ying et al.
-(2019); Liu et al.
-(2019).
-In Arvind et al.
-(2020) characterizes patterns whose count is invariant for $1-$WL and $2-$WL equivalent graphs.
-Also, there exists a GNN hierarchy, $S_{k}$ Papp and Wattenhofer (2022a), where each node has an attribute that counts the number of induced subgraphs of size at most $k$, that the node is participating in.
-It would be interesting to see whether a scalable GNN hierarchy exists, that is comparable to the $k-$WL and $S_{k}$ hierarchy.
-There also exists a GNN hierarchy, $M_{k},$ Papp and Wattenhofer (2022a); Huang et al.
-(2023)in which $k$ vertices are marked or deleted and a GNN model is run on the modified graph.
-Various subgraph based GNN models have been proposed that tackle these questions Zhao et al.
-(2022); Zhang and Li (2021); Morris et al.
-(2018); Alvarez-Gonzalez et al.
-(2022); Maron et al.
-(2019); You et al.
-(2021); Frasca et al.
-(2022); Morris et al.
-(2021); Bevilacqua et al.
-(2021); Papp and Wattenhofer (2022b); Barceló et al.
-(2021); Huang et al.
-(2023).
-These GNNs have been effective in capturing more fine-grained patterns and relationships within the graphs, and are scalable for large graphs.
-Also, it has been shown that subgraph GNNs are more expressive than the traditional ones.
-Frasca et al.
-(2022) gave an upper bound on the expressive power of subgraph $1-$WL.
-This leads to the question of coming up with upper and lower bounds for the expressiveness of subgraph $k-$WL for arbitrary $k$.
-Consider the task of counting the occurrence of $H$ as subgraph or induced subgraph in the host graph $G$.
-Given the effectiveness of subgraph $k-$WL methods in increasing the expressiveness of GNNs, we want to extend this method to the subgraph itself and check whether we can fragment the subgraph and learn the count of fragments of the subgraphs to get the actual count of $H$ in the graph.
-We are interested in evaluating its expressiveness in terms of subgraph and induced subgraph counting.
-Also, if there exists a GNN hierarchical model, we want to compare its expressiveness to pre-existing GNN hierarchies, as done in Papp and Wattenhofer (2022a).
+Graphs naturally model relational and structural data that arise across diverse domains, from social network analysis and combinatorial optimization to computational biology, particle physics, and protein folding (dill2008protein).
+Graph Neural Networks (GNNs) have emerged as powerful methods for learning graph representations for downstream tasks such as classification and link prediction.
+The expressive power of GNNs has been extensively studied, with message-passing GNNs shown to be equivalent to the $1$-dimensional Weisfeiler–Leman (WL) test (morris2019wl), a classical combinatorial tool for graph isomorphism.
+More generally, $k$-GNNs have been shown to correspond to the $k$-WL test (weisfeiler1968reduction), and thus, throughout this paper, we use $k$-WL interchangeably to denote the corresponding $k$-GNN hierarchy.
+In practice, the expressiveness of $k$-WL is determined by its ability to distinguish non-isomorphic graphs and capture subgraph patterns.
+However, while $(k+1)$-WL is provably more expressive than $k$-WL, both the time and space complexity of these algorithms grow exponentially in $k$, making higher-order WL infeasible in large-scale settings.
+A central motivation driving this work lies in the task of subgraph counting, which is central to understanding complex graph structures.
+Counting specific subgraphs provides insights into graph similarity and functionality: in computational biology, this includes motifs critical for protein binding and molecular discovery; in social networks, structures such as triangles and stars reveal community dynamics and information propagation.
+Unfortunately, detecting and counting subgraphs — a generalization of the clique problem — is NP-complete in general.
+Prior works have therefore focused on specific families of patterns or restricted graph classes (shervashidze2009efficient; JMLR:v12:shervashidze11a; arvind2017graph; https://doi.org/10.48550/arxiv.1805.02089; https://doi.org/10.48550/arxiv.2006.09252; DBLP:conf/stacs/KomarathKMS23).
+At the same time, research has sought to enrich GNN hierarchies, including $S_{k}$-GNNs (papp2022theory), which augment nodes with counts of induced subgraphs of size up to $k$, and $M_{k}$-GNNs (papp2022theory; huang2023boosting), which modify graphs via marked or deleted vertices.
+Traditional GNNs, being limited to the power of $1$-WL, cannot count even simple motifs such as triangles or cycles beyond length three (xu2019powerfulgraphneuralnetworks; morris2021weisfeilerlemanneuralhigherorder).
+Subgraph-based GNN models (zhao2022from; zhang2021nested; frasca) have emerged as extensions that apply GNNs on extracted local subgraphs or multiple partitions of the input graph, thereby improving expressiveness.
+While these approaches capture more fine-grained structures, they suffer from scalability concerns since GNNs must be run repeatedly over many subgraphs, creating prohibitive memory and time overhead.
+To overcome these limitations, we propose localized variants of Weisfeiler–Leman algorithms that balance expressiveness and computational efficiency.
+Instead of running $k$-WL across the entire graph, we restrict its application to neighborhoods, resulting in the Local $k$-WL.
+Complementary to this, we introduce a novel fragmentation technique, which decomposes complex motifs into simpler subpatterns whose counts can be reliably obtained.
+Remarkably, we prove that all induced subgraphs of size at most 4 can be counted exactly using only $1$-WL, with natural extensions to larger motifs for higher $k$.
+In addition, we design a differentiable three-stage learning framework that operationalizes this fragmentation principle: (1) identify required subpatterns, (2) count them using Local $k$-WL, and (3) aggregate counts into the global motif count.
+This approach provides both theoretical guarantees and compatibility with gradient-based learning, bridging combinatorial counting algorithms with machine learning paradigms.
+Finally, we analyze the expressive power of our methods relative to existing GNN hierarchies, demonstrating that under bounded complexity, our models achieve strictly higher expressivity.
 1.1 Our Contributions
-In this paper, we attempt to answer these questions. The main contributions of our work are listed below:
+We summarize our key contributions:
   1. 1.
-    Characterizing expressiveness of Local $k-$WL: Given a graph $G=(V,E),$ we extract a $r$-hop subgraph rooted at each vertex $v\in V,$ say $G_{v}^{r},$ and run $k-$WL on $G_{v}^{r}.$ While GNNs based on subgraphs have been proposed in recent papers , this is the first work that gives both upper and lower bounds for the expressiveness of Local $k-$WL.
-We are also the first to characterize patterns that can be counted exactly by Local $k-$WL.
+    Expressiveness of Local $k$-WL: We formally introduce Local $k$-WL, which applies $k$-WL to $r$-hop neighborhoods $G_{v}^{r}$ rooted at $v\in V.$ We provide, for the first time, both upper and lower bounds for its expressiveness and characterize precisely the subgraph patterns that can be distinguished and counted.
   2. 2.
-    Layer $k-$WL: To improve the space and time complexity of Local $k-$WL, we propose the Layer $k-$WL method.
-For this method, instead of running $k-$WL on $G_{v}^{r}$, we run it on two consecutive layers of vertices.
-Here, the $i$th layer of vertices refers to the vertices that appear at an $i$-hop distance from $v$(or the $i$th layer breadth-first search(BFS)).
+    Layer $k$-WL: A scalable variant where $k$-WL is restricted to pairs of consecutive BFS layers, reducing time and space overhead compared to standard Local $k$-WL.
   3. 3.
-    Recursive WL : Recursive WL is an alternative to $k-$WL.
-In this method, we first run $1-$WL to get a partitioning of vertices.
-Then we run $(k-1)-$WL on the vertices of each partition separately.
-It can be shown that this method is more expressive than $(k-1)-$WL and less expressive than $k-$WL.
-Also, since we are running $(k-1)-$WL on a smaller set of vertices, it has better space and time complexity than running $k-$WL.
+    Recursive WL: A hierarchical alternative to $k$-WL.
+We first apply $1$-WL to partition nodes, and then run (k1)(k-1)-WL on smaller vertex sets.
+This achieves higher expressiveness than (k1)(k-1)-WL while maintaining better scalability than $k$-WL.
   4. 4.
-    Fragmentation : For the counting task, based on the pattern $H$ to be counted, the subgraph $G_{v}^{r}$ is further decomposed into simpler patterns for which the exact counts of subpatterns are already known.
-Thus, we need to learn the easier tasks in the subgraphs of $G_{v}^{r}$.
-So, a smaller $k$ is sufficient to count the number of patterns.
-Using this method, we show that all the patterns appearing as induced subgraphs of size four can be counted using just $1-$WL.
-This technique can be useful for counting larger and more complicated patterns.
-Thus, instead of training a GNN for the larger subgraph, we can train GNN models for the smaller patterns for counting and then combine their counts to get the count of the larger subgraph.
-Using the fragmentation technique, we use the model learned for predicting $K_{3}$ or a triangle to predict the number of $K_{4}$ in the graph.
-Similarly, if we have a model that can predict $K_{n}$ in a graph, then we can use it to predict $K_{n+1}$.
-In other words, we can reduce the counting of $K_{n+1}$ to a triangle counting problem with a minimal increase in the number of parameters.
+    Fragmentation for Subgraph Counting: We propose a decomposition technique that reduces the counting of complex motifs to that of simpler subpatterns.
+For example, we reduce $K_{4}$ counting to repeated triangle counting.
+We prove that all induced subgraphs of size at most 4 can be counted with $1$-WL and extend this methodology to larger patterns.
   5. 5.
-    Comparison with other GNN models : Papp and Wattenhofer (2022a) shows an analysis of four GNN models.
-We do a similar analysis for our models and compare them with the models mentioned in Papp and Wattenhofer (2022a).
-We show that our models are more expressive than the ones presented in that paper.
+    Comparative Expressiveness: We provide a rigorous comparison of Local $k$-WL, Layer $k$-WL, and Recursive WL against existing GNN hierarchies such as $S_{k}$ and $M_{k}$.
+Our results show that, under comparable complexity bounds, our models are strictly more expressive.
+In summary, our work advances subgraph counting in GNNs by combining the theoretical rigor of WL with scalable localized techniques, while introducing a new fragmentation paradigm that makes counting feasible, differentiable, and efficient for real-world graph applications.
 Outline of the paper :
-In Section 2, we introduce some of the terms used throughout the paper.
+Section 2 introduces some of the terms used throughout the paper.
 In Section 4, we introduce the localized variants of the $k$-WL algorithm and analyze their space and time complexities.
-In Section 5, we give theorems that characterize the expressiveness of the localized $k-$WL variants proposed in our work.
+Section 5 gives theorems that characterize the expressiveness of the localized $k$-WL variants proposed in our work.
 In Section 6, we characterize the expressiveness of our methods in terms of subgraph and induced subgraph counting.
 We also discuss how to count the occurrences of $H$ in $G,$ using localized algorithms.
-We discuss the fragmentation approach in Section 7, followed by a theoretical comparison of GNN models in Section 8.
-The model architecture, along with the parameters used for the experiment, is explained in Section 9 We report the results of our experiments in Section 10 and conclude the paper with Section 11.
+We discuss the fragmentation approach in Section 7.4, followed by a theoretical comparison of GNN models in Section 8.
+The model architecture in Section 7.
+Discusson on implementation details, hyperparameters and results of our experiments is in Section 9 and conclude the paper with Section 10.
 2 Preliminaries
-We consider a simple graph $G(V,E)$.
-For basic definitions of graph theory, we refer the reader to West et al.
-(2001).
-The neighbourhood of a vertex $v$ is the set of all vertices adjacent to it in $G$ (denoted as $N_{G}(v)$).
-The closed neighbourhood of $v$ is the set of all neighbours, including the vertex $v$ (denoted as $N_{G}[v]$).
-A graph whose all the vertices are of same degree are called regular graph.
-A graph $H$ is called a subgraph of $G$ if $V(H)\subseteq V(G)$ and $E(H)\subseteq E(G)$.
-The subgraph induced on $S\subseteq V(G)$ is a graph whose vertex set $S$ contains all the edges in $G$ whose endpoints are in $S$ and is denoted by $G[S]$.
-The induced subgraph on a $r$-hop neighbourhood around vertex $v$ is denoted by $G_{v}^{r}$.
-Attributed subgraphs are coloured subgraphs, also referred to as Motifs.
-The maximum distance from a vertex to all other vertices is called the eccentricity of the vertex.
-The minimum of the eccentricity over all the vertices is called the radius of the graph.
-The center of the graph is set of vertices such that eccentricity is minimum.
-For pattern counting, we pick one of the center vertex and call it key vertex.
-Homomorphism from graph $H$ to $G$ is a function from $V(H)$ to $V(G)$ such that if $\{u,v\}\in E(H)$ then $\{f(u),f(v)\}\in E(G)$.
-Given a pattern $H$, the set of all of its homomorphic images is called spasm of $H$.
-Two graphs $G$ and $H$ are isomorphic if there exists a bijective function $f:V(G)\xrightarrow{}V(H)$ such that $\{u,v\}\in E(G)$ if and only if $\{f(u),f(v)\}\in E(H)$.
-The orbit of a vertex $v$ in $G$ is the set of vertices to which $v$ can be mapped, and that mapping can be extended to automorphism (denoted by $Orbit_{G}(v)$).
-We mention some of the structural parameters of graphs. Problems on graph with bounded structural parameters can be solved efficiently on bounded graph parameters.
-Graph Parameters:
-We first define tree decomposition of the graph as:
-Given a graph $G$, we decompose it into tree structures, say $T,$ where the set of vertices of $T$ is a subset of set of vertices of $G.$ This decompostion has to satisfy the following constraints:
-  1. 1.
-    Every vertex of $G$ must lie in some bag associated with a vertex of $T$.
-  2. 2.
-    For each edge $\{v_{i},v_{j}\}$, there exists a bag containing having $v_{i}$ and $v_{j}$.
-  3. 3.
-    If a vertex $v_{i}\in V(G)$ belongs to two bags $B_{i}$ and $B_{j}$ associated with two vertices $u_{i}$ and $u_{j}$ of $T,$ then $v_{i}$ must be present in all the bags associated with the vertices belonging to the path connecting $u_{i}$ and $u_{j}$ in $T.$
-The width of the tree decomposition is the maximum size of the bag minus one.
-The treewidth of the graph $G,$ $tw(G)$ is the minimum over all such decompositions.
-It is NP-hard to compute tree-width of graphs.
-However, there exists an efficient algorithm that checks for a fixed $k$, whether $tw(G)$ is at most $k$ Korhonen (2022); Korhonen and Lokshtanov (2022).
-Graph of bounded treewidth implies sparse graphs.
-However, for some sparse graph, the treewidth is unbounded.
-For example, grid graph on $n$ vertices has treewidth $\sqrt{n}$.
-The maximum of the treewidth over all its homomorphic images is called the hereditary treewidth of pattern $H,$ denoted by $htw(H)$.
-Planar graphs are graphs that are $k_{5}$ and $k_{3,3}$ minor free.
-One can also say the graph that can be redrawn such that no edges cross each other are planar graphs.
-Also, the number of edges can be at most linear in the number of vertices.
-The Euler genus of a graph is defined in a similar manner.
-The genus of a graph is the minimum number such that the graph can be drawn on circle without intersecting edges.
-Now, we look at graph classes that are dense but have nice structure such as complete graphs.
-Clique width has been defined for dense graphs.
-However, there does not exist efficient algorithm to check whether clique width of the given graphs is $k$ for $k\geq 4$.
-Rankwidth has been defined by Robertson and Seymour to handle dense graph classes.
-Given a graph $G$, to construct rankwidth decomposition, we define subcubic tree $T$.
-Given a bijection from the set of leaves of a tree to the set of vertices of $G$, we can construct rankwidth decomposition with respect to that bijection.
-The parent vertices in the tree contains union of vertices belonging to its two children.
-Note that deletion of single edge disconnects the tree and vertices of graph get partitioned into two subparts say $X$ and $V(G)\setminus X$.
-We define submatrix of adjacency matrix $A(X,V(G)\setminus X)$ where $a_{i,j}=1$ if and only if $i\in X$ and $j\in V(G)\setminus X$ are adjacent.
-Also, it has been shown that bounded clique width means bounded rank-width and vice versa Oum (2017).
+We consider a simple undirected graph $G=(V,E)$ with vertex set $V$ and edge set $E$.
+For basic graph-theoretic definitions, we refer the reader to (west2001introduction).
+The neighbourhood of a vertex $v\in V$ is the set of all adjacent vertices, denoted by $N_{G}(v)$ (or $\mathcal{N}_{G}(v)$).
+The closed neighbourhood, denoted $N_{G}[v]$, includes $v$ itself.
+The degree of a node $v$ is denoted by $d_{v}$.
+The maximum distance from $v$ to any other vertex is its eccentricity, and the minimum eccentricity over all vertices gives the radius of the graph.
+A graph in which every vertex has the same degree is called a regular graph.
+A graph $H$ is a subgraph of $G$ if $V(H)\subseteq V(G)$ and $E(H)\subseteq E(G)$.
+The subgraph induced on $S\subseteq V(G)$, denoted $G[S]$, contains vertex set $S$ with all edges in $G$ whose endpoints lie in $S$.
+The induced subgraph on an $r$-hop neighbourhood of a vertex $v$ is denoted by $G_{v}^{r}$.
+The hop parameter $r$ depends on the pattern being counted.
+For example, $r=1$ suffices for counting triangles, while $r=2$ is required for counting $C_{4}$.
+Attributed subgraphs (or motifs) are subgraphs whose vertices or edges carry additional labels or colours.
+A graph homomorphism from $H$ to $G$ is a function $f:V(H)\to V(G)$ such that $\{u,v\}\in E(H)$ implies $\{f(u),f(v)\}\in E(G)$.
+The set of all homomorphic images of a pattern $H$ is called the spasm of $H$.
+Two graphs $G$ and $H$ are isomorphic if there exists a bijection $f:V(G)\to V(H)$ preserving adjacency.
+The orbit of a vertex $v$ in $G$, denoted $Orbit_{G}(v)$, is the set of vertices to which $v$ can be mapped under automorphisms of $G$.
+2.1 Graph Parameters
+Many hard graph problems become tractable on restricted graph classes characterized by structural parameters. We briefly review key notions.
+Treewidth.
+A tree decomposition of $G$ expresses it as a tree of vertex “bags” satisfying: (i) every vertex of $G$ appears in some bag, (ii) every edge is contained in some bag, and (iii) for each vertex $v$, bags containing $v$ form a connected subtree.
+The width of a decomposition is the maximum bag size minus one, and the treewidth $tw(G)$ is the minimum width across all decompositions.
+Computing $tw(G)$ is NP-hard, but fixed-parameter algorithms exist (korhonen2022single; korhonen2022improved).
+The hereditary treewidth $htw(H)$ of a pattern $H$ is the maximum treewidth among its homomorphic images.
+Planar graphs and genus.
+A graph is planar if it can be drawn without crossing edges, or equivalently, if it excludes $K_{5}$ and $K_{3,3}$ as minors.
+More generally, the Euler genus denotes the minimal surface on which a graph can be embedded without edge crossings.
+Clique-width and rank-width.
+For dense graphs, clique-width provides a measure of complexity, but recognizing graphs of clique-width at most $k\geq 4$ is NP-hard.
+Rank-width, introduced by Robertson and Seymour, offers an alternative based on rank decompositions of the adjacency matrix.
+It is known that bounded clique-width implies bounded rank-width, and vice versa (oum2017rank).
+2.2 Weisfeiler–Leman and Local $k$-WL
+The Weisfeiler–Leman (WL) test (weisfeiler1968reduction) is a colour refinement algorithm for testing graph isomorphism.
+Its $k$-dimensional extension, $k$-WL, refines colours of $k$-tuples of vertices, with expressiveness increasing in $k$.
+Applying $k$-WL to a full graph with $n$ vertices requires $O(n^{k+1}\log n)$ time.
+In contrast, the Local $k$-WL algorithm applies $k$-WL to restricted neighbourhoods $G_{v}^{r}$, reducing complexity.
+For graphs of maximum degree $d$, local $k$-WL can be run in $O(n\cdot d^{r(k+1)}\log d)$ time.
+2.3 Graph Neural Networks
+Graph Neural Networks (GNNs) (DBLP:journals/corr/KipfW16) generalize neural architectures to graphs using recursive MESSAGE and AGGREGATE functions.
+At the $\ell$-th iteration, node embeddings are updated as:
+$X_{v}^{(\ell)}=\texttt{UPDATE}^{(\ell)}\left(X_{v}^{(\ell-1)},\;\texttt{AGGREGATE}^{(\ell)}\Big\{\texttt{MESSAGE}^{(\ell)}\big(X_{v}^{(\ell-1)},\,X_{u}^{(\ell-1)}\big)\;\big|\;u\in N_{G}(v)\Big\}\right).$ (1)
+It is known that standard message-passing GNNs are bounded in expressiveness by $1$-WL, motivating higher-order and subgraph-based extensions.
 3 Weisfeiler Leman Algorithm
-Weisfeiler-Leman (WL) is a well known combinatorial algorithm that has many theoretical and practical applications.
-Color refinement($1-$WL) was first introduced in 1965 in Morgan (1965).
+Weisfeiler-Leman (WL) is a well-known combinatorial algorithm that has many theoretical and practical applications.
+Color refinement(11-WL) was first introduced in 1965 in (morgan1965generation).
 The algorithm goes as follows:
   • Initially, color all the vertices as color 1.
-  • In the next iteration $i,$ we color the vertices by looking at the number of colors of vertices adjacent to each vertex $v,$ in the $(i-1)$th iteration, as
+  • In the next iteration $i,$ we color the vertices by looking at the number of colors of vertices adjacent to each vertex $v,$ in the (i1)(i-1)th iteration, as
     $C_{i}(v)=(C_{i-1}(v),\{\{C_{i}(w)\}\}_{w\in N_{G}(v)})$
@@
 We can observe that the number of iterations is at most $n$ as a vertex set ,$V(G),$ can be partitioned at most $n$ many times.
-  • The color class of any vertex $v\in V(G)$ can appear at most $1+\log n$ times and the running time is $\mathcal{O}(n^{2}\log n)$ Immerman and Sengupta (2019).
+  • The color class of any vertex $v\in V(G)$ can appear at most $1+\log n$ times and the running time is $\mathcal{O}(n^{2}\log n)$ (immerman2019kdimensional).
 In case we need to run only $h$ iterations and stop before getting the stable color, then the running time is $O(nh)$.
@@
 The iteration in which no further partition is being done are called stable coloring.
-Weisfeiler Leman algorithm which is known as $2-$WL algorithm.
+Weisfeiler Leman algorithm which is known as 22-WL algorithm.
 Similar approach was extended later for coloring $k$-tuples and then do refinement of coloring in later iterations.
 Definition 1.
-Let $\vec{x}=(x_{1},...,x_{k})\in V^{k}$,$y\in V$ , and $1\leq j\leq k$.
+Let $\vec{x}=(x_{1},\dots,x_{k})\in V^{k}$$y\in V$ , and $1\leq j\leq k$.
 Then, let $x[j,y]\in V^{k}$ denote the $k$-tuple obtained from $x$ by replacing $x_{j}$ by $y$.
-The $k$-tuples $\vec{x}[j,y]$ and $\vec{x}$ are said to be $j$-neighbors for any $y\in V$ We also say $\vec{x}[j,y]$ is the j-neighbor of $\vec{x}$ corresponding to $y$.
+The $k$-tuples $\vec{x}[j,y]$ and $\vec{x}$ are said to be $j$-neighbors for any $y\in V$.
+We also say $\vec{x}[j,y]$ is the j-neighbor of $\vec{x}$ corresponding to $y$.
   • Color all the $k$-tuple vertices according to their isomorphic type.
@@
 We can observe that the number of iterations is at most $n^{k}$ as $V^{k}$ can be partitioned at most $n^{k}$ many times.
-  • The color class of any vertex $\vec{v}\in V^{k}$ can appear at most $\mathcal{O}(k\log n)$ times and running time is $\mathcal{O}(k^{2}n^{k+1}\log n)$ Immerman and Sengupta (2019).
+  • The color class of any vertex $\vec{v}\in V^{k}$ can appear at most $\mathcal{O}(k\log n)$ times and running time is $\mathcal{O}(k^{2}n^{k+1}\log n)$ (immerman2019kdimensional).
 Two graphs $G$ and $H$ are said to be $k$-WL equivalent ($G\simeq_{k}H$), if their color histogram of the stable colors matches.
 We say that $G$ is $k$-WL identifiable if there doesn’t exist any non-isomorphic graphs that are $k$-WL equivalent to $G$.
-Color refinement ($1-$WL) can recognise almost all graphs Babai et al.
-(1980), while $2-$WL can recognise almost all regular graphs Bollobás (1982).
+Color refinement (11-WL) can recognise almost all graphs (babai1980random), while $2$-WL can recognise almost all regular graphs (bollobas1982distinguishing).
 The power of $WL$ increases with an increase in the value of $k$.
@@
 Also, the power of $k$-WL to distinguish two non-isomorphic graphs is equivalent to spoiler’s winning condition in $(k+1)$-bijective pebble game.
-Recently, Dell et al.
-(2018) has shown that the expressive power of $k-$WL is captured by homomorphism count.
+Recently, (Dell2018LovszMW) has shown that the expressive power of $k$-WL is captured by homomorphism count.
 It has been shown that $G_{1}\simeq_{k}G_{2}$ if and only if $Hom(T,G_{1})=Hom(T,G_{2}),$ for all graphs $T$ of treewidth at most $k$.
-The graphs that are identified by $1-$WL are Amenable graphs.
-There is a complete characterization of the amenable graphs in Arvind et al.
-(2017); Kiefer et al.
-(2015).
-In the original algorithm, we initially color all the vertices with color $1$.
-However, if we are given a colored graph as input, we start with the given colors as the initial colors.
-Also, we can color the edges, and run $1-$WL Kiefer et al.
-(2015).
-Even if ${k-WL}$ may not distinguish two non-isomorphic graphs, two ${k-WL}$ equivalent graphs have many invariant properties.
-It is well known that two ${1-WL}$ equivalent graphs have the same maximum eigenvalue.
-Two graphs that are ${2-WL}$ equivalent are co-spectral and have the same diameter.
+The graphs that are identified by 11-WL are Amenable graphs.
+There is a complete characterization of the amenable graphs in (arvind2017graph; kiefer2015graphs).
+In the original algorithm, we initially colour all the vertices with colour $1$.
+However, if we are given a coloured graph as input, we start with the given colours as the initial colours.
+Also, we can color the edges and run $1$-WL (kiefer2015graphs).
+Even if kWL{k-WL} may not distinguish two non-isomorphic graphs, two kWL{k-WL} equivalent graphs have many invariant properties.
+It is well known that two 1WL{1-WL} equivalent graphs have the same maximum eigenvalue.
+Two graphs that are 2WL{2-WL} equivalent are co-spectral and have the same diameter.
 Recently, V.
 Arvind et al.
-have shown the invariance in terms of subgraph existence and counts Arvind et al.
-(2020).
-They show the complete characterization of subgraphs whose count and existence are invariant for ${1-WL}$ equivalent graph pairs.
-They also listed down matching , cycles and path count invariance for ${2-WL}$.
-Also, there is a relation between homomorphism count and subgraph count Curticapean et al.
-(2017).
-The count of subgraphs is a function of the number of homomorphism from set of all homomorphic image of patterns.
-Hereditary treewidth of graph is defined as maximum of treewidth over all homomorphic images.
-So, if two graphs are ${k-WL}$ equivalent, then the count of all subgraphs whose $htw(G)$ is almost $k$ are same .
-However, running ${k-WL}$ takes $O(k^{2}\cdot n^{k+1}logn)$ time and $O(n^{k})$ space Immerman and Sengupta (2019).
-So, it is not practically feasible to run ${k-WL},$ for large $k,$ for graphs .
-The expressive power of $k-$WL is equivalent to first order logic on $(k+1)$ variables with a counting quantifier.
-Let $G=(V,E),$ where $V$ is a set of vertices and $E$ is a set of edges.
+have shown the invariance in terms of subgraph existence and counts (arvind2020weisfeiler).
+They show the complete characterization of subgraphs whose count and existence are invariant for 1WL{1-WL} equivalent graph pairs.
+They also listed down matching, cycles, and path count invariance for 2WL{2-WL}.
+Also, there is a relation between homomorphism count and subgraph count (curticapean2017homomorphisms).
+The count of subgraphs is a function of the number of homomorphisms from a set of all homomorphic images of patterns.
+Hereditary treewidth of a graph is defined as the maximum treewidth over all homomorphic images.
+So, if two graphs are kWL{k-WL} equivalent, then the count of all subgraphs whose $htw(G)$ is almost $k$ is the same.
+However, running kWL{k-WL} takes $O(k^{2}\cdot n^{k+1}logn)$ time and $O(n^{k})$ space (immerman2019kdimensional).
+So, it is not practically feasible to run kWL{k-WL} for large $k$, for graphs.
+The expressive power of $k$-WL is equivalent to first-order logic on $(k+1)$ variables with a counting quantifier.
+Let $G=(V,E)$ where $V$ is a set of vertices and $E$ is a set of edges.
 In logic, we define $V$ as the universe and $E$ as a binary relation.
-In Cai et al.
-(1992), they have proved that the power to distinguish two non-isomorphic graphs using $k-$WL is equivalent to $C^{k+1},$ where $C^{k+1}$ represents first order logic on $(k+1)$ variables with counting quantifiers (stated in Theorem 1).
+In (cai1992optimal), they have proved that the power to distinguish two non-isomorphic graphs using $k$-WL is equivalent to $C^{k+1}$, where $C^{k+1}$ represents first-order logic on $(k+1)$ variables with counting quantifiers (stated in 1).
 To prove this, they define a bijective $k$-pebble game, whose power is equivalent to $C^{k}$.
 Bijective k-Pebble Game
-The bijective k-Pebble game ($BP_{k}(G,H)$) has been discussed in Kiefer (2020); Cai et al.
-(1992); Grohe and Neuen (2021).
+The bijective k-Pebble game ($BP_{k}(G,H)$) has been discussed in (kiefer2020power; cai1992optimal; rankwidth_WL).
 Let graphs $G$ and $H$ have the same number of vertices and $k\in\mathbb{N}$.
@@
 Definition 2.
-The position of the game in bijective pebble game is the tuples of the vertices where the pebbles are placed.
+The position of the game in a bijective pebble game is the tuples of the vertices where the pebbles are placed.
 The bijective $k$-pebble game is defined as follows:
@@
 Theorem 1.
-Cai et al. (1992) Let $G$ and $H$ be two graphs. Then $G\simeq_{k}H$ if and only if the Duplicator wins the pebble game $BP_{k+1}(G,H).$
+(cai1992optimal) Let $G$ and $H$ be two graphs. Then $G\simeq_{k}H$ if and only if the Duplicator wins the pebble game $BP_{k+1}(G,H).$
 A stronger result, namely, the equivalence between the number of rounds in the bijective $(k+1)$-pebble game and the iteration number of $k$-WL was stated in the following theorem.
 Theorem 2.
-Kiefer (2020) Let $G$ and $H$ be graphs of same size.
-The vertices may or may not be colored.
-Let $\vec{u}:=(u_{1},...,u_{k})\in(V(G))^{k}$ and $\vec{v}:=(v_{1},...,v_{k})\in(V(H))^{k}$ be any two arbitrary elements.
+(kiefer2020power) Let $G$ and $H$ be graphs of the same size.
+The vertices may or may not be coloured.
+Let $\vec{u}:=(u_{1},\dots,u_{k})\in(V(G))^{k}$ and $\vec{v}:=(v_{1},\dots,v_{k})\in(V(H))^{k}$ be any two arbitrary elements.
 Then, for all $i\in\mathbb{N}$, the following are equivalent :
@@
 4 Local k-WL based Algorithms for GNNs
-In this section, we present the local $k-$WL based algorithms for GNNs. We also give runtime and space requirements for such GNNs.
+This section presents the local $k$-WL based algorithms for GNNs. We also give runtime and space requirements for such GNNs.
 4.1 Local k-WL
@@
 Then, we run a GNN on the graph $G$ with the attributes on each vertex $v$.
-This is described in Algorithm 1.
+This is described in 1.
 Algorithm 1 Local k-WL
@@
 3:  Find the subgraph induced on the $r$-hop neighborhood rooted at vertex $v$ ($G_{v}^{r}$).
-4:     Color the vertices whose distance from $v$ is $i,$ by color $i$
+4:  Color the vertices whose distance from $v$ is $i,$ by color $i$.
 5:  Run $k$-WL on the colored graph until the colors stabilize.
 6:end for
-7:Each vertex has as an attribute the stable coloring of vertex $v$ obtained from $G_{v}^{r}$.
+7:Each vertex has as an attribute the stable colouring of vertex $v$ obtained from $G_{v}^{r}$.
 8:Run GNN on the graph $G$ with each vertex having attributes as computed above.
 Runtime and Space requirement Analysis : The time required to run $k$-WL on $n$ vertices is $O(n^{k+1}\log(n))$.
-Here, we run $k-$WL on a $r$-hop neighborhood instead of the entire graph.
-So, $n$ is replaced by $n_{1},$ where $n_{1}$ is the size of the neighborhood.
-If a graph has bounded degree $d$, and we run $k-$WL for a $2$-hop neighborhood, then $n_{1}$ is $O(d^{2})$.
+Here, we run $k$-WL on a $r$-hop neighbourhood instead of the entire graph.
+So, $n$ is replaced by $n_{1},$ where $n_{1}$ is the size of the neighbourhood.
+If a graph has bounded degree $d$, and we run $k$-WL for a 22-hop neighbourhood, then $n_{1}$ is $O(d^{2})$.
 Also, we have to run Local $k$-WL for each vertex.
@@
 4.2 Layer k-WL
-In order to make Local $k-$WL more time and space efficient, while maintaining the same expressive power, we propose a modification to Local $k-$WL.
+In order to make Local $k$-WL more time and space-efficient while maintaining the same expressive power, we propose a modification to Local $k$-WL.
 Instead of running $k$-WL on the entire $r$-hop neighbourhood, we run $k$-WL on consecutive layers of $G_{v}^{r}$ (i.e., run $k$-WL on the set of vertices with colour $i$ and colour $(i+1)$).
@@
 While running $k$-WL, initially, it partitions the $k$-tuples based on the isomorphism type.
-However, in this setting, we incorporate the stabilized colouring obtained in the previous round.
+However, we incorporate the stabilized colouring obtained in the previous round in this setting.
 For $l<k,$ we define the color of $l$ tuples as $col(u_{1},u_{2},...,u_{l}):=col(u_{1},u_{2},...,u_{l},\underbrace{u_{1},u_{1}..,,u_{1}}_{(k-l)\text{times}})$.
@@
 So, even if $(u_{1},v_{1},\ldots,u_{k})$ and $(u_{1}^{{}^{\prime}},v_{1}^{{}^{\prime}},\ldots,u_{k}^{{}^{\prime}})$ may be isomorphic, if $col(u_{1},u_{2},\ldots u_{l})\neq col(u_{1}^{{}^{\prime}},u_{2}^{{}^{\prime}},\ldots u_{l}^{{}^{\prime}})$ then $col(u_{1},v_{1},\ldots,u_{k})\neq col(u_{1}^{{}^{\prime}},v_{1}^{{}^{\prime}},\ldots,u_{k}^{{}^{\prime}})$.
-The algorithm is described in Algorithm 2.
-A GNN model incorporating Local+Layer $k-$WL is equivalent to running Layer $k-$WL in line 5 in Algorithm 1.
+The algorithm is described in 2.
+A GNN model incorporating Local+Layer $k$-WL is equivalent to running Layer $k$-WL in line 5 in 1.
 Algorithm 2 Layer k-WL($v$)
 1:Given $G_{v}^{r},k$.
-2:Run $k-$WL on the induced subgraph of levels $1$ and $2$.
+2:Run $k$-WL on the induced subgraph of levels 11 and $2$.
 3:for each layer $i$ of BFS($v$), $i\geq 2$ do
-4:     Initial colour of $k-tuple$ incorporates the stabilized colour obtained from the previous iteration.
+4:  Initial colour of ktuplek-tuple incorporates the stabilized colour obtained from the previous iteration.
 5:  Run $k$-WL on the subgraph induced on the vertices in layer $i$ and $(i+1)$
@@
 Here, we present another variant of WL.
-The central idea is to decompose the graphs initially by running $1-$WL.
+The central idea is to decompose the graphs initially by running 11-WL.
 Then, further, decompose the graphs by running $2$-WL and so on.
-One can note that the final vertex partition that $1-$WL outputs after color refinement are regular if we restrict to a single color class.
-In other words, let $G[X]$ be the induced graph on the vertices of same color.
+One can note that the final vertex partition that 11-WL outputs after colour refinement is regular if we restrict it to a single colour class.
+In other words, let $G[X]$ be the induced graph on the vertices of the same colour.
 Then, $G[X]$ is regular.
-Also, $G[X,Y]$ where $X$ and $Y$ are sets of vertices of two different color classes.
+Also, $G[X,Y]$ where $X$ and $Y$ are sets of vertices of two different colour classes.
 $G[X,Y]$ is also a bi-regular graph.
 We run $2$-WL on the regular graph.
-Using Bollobás (1982), we can guarantee that it would distinguish almost all regular graphs.
-Similarly, running $2-$WL on $G[X,Y]$ is bi-regular and thus can be distinguished by $2-$WL.
-We again run $1-$WL on $G,$ using the colors obtained after running $2-$WL.
-This further refines the colors of the vertices in $G$.
+Using (bollobas1982distinguishing), we can guarantee that it would distinguish almost all regular graphs.
+Similarly, running $2$-WL on $G[X,Y]$ is bi-regular and thus can be distinguished by 22-WL.
+We again run $1$-WL on $G,$ using the colours obtained after running 22-WL.
+This further refines the colours of the vertices in $G$.
 One can easily check that it is more expressive than $1$-WL and less expressive than $2$-WL.
-We give the graph Figure 1 that can not be distinguished by Recursive $(1,2)-$WL and $1-$WL but can be distinguished by $2-$WL.
+We give the graph 1 that can not be distinguished by Recursive $(1,2)$-WL and 11-WL but can be distinguished by 22-WL.
 This gives an intermediate hierarchy in the $k$-WL hierarchy.
 Also, the space and time required for running $2$-WL on the entire graph is more than that of Recursive $(1,2)$-WL.
-The running time and space required depend on the partition size obtained after running $1-$WL.
-Note that the color of vertex $v$ is $col(v,v)$ after running $2-$WL.
+The running time and space required depend on the partition size obtained after running 11-WL.
+Note that the colour of vertex $v$ is $col(v,v)$ after running 22-WL.
 Algorithm 3 Recursive(1,2) WL
 1:Given $G$
-2:Run $1-$WL and get the partition of vertices into colour classes.
-3:Let $S=\{C_{1},C_{2},\ldots C_{l}\}$ be the color classes obtained after running $1-$WL.
+2:Run 11-WL and get the partition of vertices into colour classes.
+3:Let $S=\{C_{1},C_{2},\ldots C_{l}\}$ be the color classes obtained after running 11-WL.
 4:for each color class $C_{i}$ in $S$ do
-5:     Run $2-$WL on the induced subgraph in $C_{i}$ and get color partition.
+5:  Run 22-WL on the induced subgraph in $C_{i}$ and get color partition.
 6:  Let $C_{i}$ get partitioned into $C_{i,1},C_{i,2},\ldots,C_{i,l}$
 7:end for
-8:Run $1-$WL on the colored graph $G$ whose colors are given by steps 5 and 6.
+8:Run 11-WL on the colored graph $G$ whose colors are given by steps 5 and 6.
 9:for each new color class $C^{\prime}_{i}$ and $C^{\prime}_{j}$ do
-10:     Run $2-$WL on the induced subgraph on the vertices in color partitions $C^{\prime}_{i}$ and $C^{\prime}_{j}$ and get new color partition.
+10:  Run 22-WL on the induced subgraph on the vertices in color partitions $C^{\prime}_{i}$ and $C^{\prime}_{j}$ and get new color partition.
 11:end for
@@
 This reduces the time and space required for running $k_{2}$-WL on the entire graph.
-One can easily see that it is less expressive than $k_{2}-$WL, however, more expressive than $k_{1}-$WL.
-More specifically, initially run $1-$WL and further run $(k-1)-$WL on the colored classes. One can check that it is more expressive than $(k-1)-$WL and less expressive than $k-$WL.
-Figure 1: Graph identifiable by 2-WL but not by Recursive $1-$WL
+One can easily see that it is less expressive than $k_{2}$-WL; however, it is more expressive than $k_{1}$-WL.
+More specifically, initially run $1$-WL and further run (k1)(k-1)-WL on the coloured classes. One can check that it is more expressive than (k1)(k-1)-WL and less expressive than $k$-WL.
+Figure 1: Graphs identifiable by 2-WL but not by Recursive 11-WL
 5 Theoretical Guarantee of Expressive Power
-In this section, we theoretically prove the expressive power of GNN models that we proposed in Section 4 in terms of graph and subgraph isomorphism.
+In this section, we theoretically prove the expressive power of GNN models that we proposed in 4 in terms of graph and subgraph isomorphism.
 In the discussion below, we say that a GNN model $A$ is at most as expressive as a GNN model $B$ if any pair of non-isomorphic graphs $G$ and $H$ that can be distinguished by $A$ can also be distinguished by $B$.
@@
 5.1 Local k-WL
-It has been shown in recent works that running Local $1-$WL has more expressive power as compared to running $1-$WL.
-The upper bound on the expressive power of Local $1-$WL has been shown in Frasca et al.
-(2022).
+It has been shown in recent works that running Local $1$-WL has more expressive power as compared to running 11-WL.
+The upper bound on the expressive power of Local $1$-WL has been shown in (frasca).
 However, the expressive power of Local $k$-WL, for arbitrary $k,$ has not been studied.
@@
 We also show that it is less expressive than $(k+1)$-WL.
-The proof techniques that we used are different from Frasca et al.
-(2022).
+The proof techniques that we used are different from (frasca).
 Theorem 3.
@@
 Let $G_{1}$ and $G_{2}$ be two graphs distinguishable by $k$-WL.
-So, the Spoiler has a winning strategy in the game ($G_{1}$,$G_{2}$).
+So, the Spoiler has a winning strategy in the game ($G_{1}$$G_{2}$).
 Suppose $G_{1}$ and $G_{2}$ do not get distinguished after running $k$-WL locally.
@@
 We can show using induction on the number of rounds that if $S_{1}$ wins the game, then $S_{2}$ also wins the game.
-Our induction hypothesis is that $S_{1}$ has not won till the $j^{th}$ round and the positions of both the games are same.
+Our induction hypothesis is that $S_{1}$ has not won till the $j^{th}$ round, and the positions of both games are the same.
 Let the current position of both the games after the $j^{th}$ round be $((v_{1},v_{2},\ldots,v_{l}),(f(v_{1}),f(v_{2}),\ldots,f(v_{l}))$.
@@
 $S_{2}$ also chooses the same vertices.
-So, the position of both the game is the same.
+So, the position of both the games is the same.
 Therefore, if $S_{1}$ wins the game, then $S_{2}$ also wins the game.
@@
 We can show that it is more expressive by looking at the simple example that running $1$-WL on a local substructure can count the number of triangles, whereas running $1$-WL on an entire graph does not recognize graphs having different triangle counts.
-Also, one can observe that running $k-$WL locally is running $k-$WL on the colored graph where vertices at distinct distances get distinct colors.
+Also, one can observe that running $k$-WL locally is running $k$-WL on the coloured graph where vertices at distinct distances get distinct colours.
 Its power is the same as individualizing one vertex and running $k$-WL.
@@
 Let $G_{1}$ and $G_{2}$ be two graphs that can be distinguished by running $k$-WL locally.
-Recall that, the key vertices refer to $u$ and $v$ in $G_{1}$ and $G_{2}$ such that they are the root vertices corresponding to $G_{1}$ and $G_{2}$, respectively.
+Recall that the key vertices refer to $u$ and $v$ in $G_{1}$ and $G_{2}$ such that they are the root vertices corresponding to $G_{1}$ and $G_{2}$, respectively.
 This means the Spoiler has a winning strategy in the $(k+1)$ bijective pebble game, where the key vertices are matched to each other.
@@
 Proof.
-Let $G$ and $H$ be the subgraphs induced on the $r$-hop neighborhood.
+Let $G$ and $H$ be the subgraphs induced on the $r$-hop neighbourhood.
 Let $(S,D)$ be the Spoiler-Duplicator pair for the game $(G,H)$.
@@
 We prove this using induction on the number of rounds of the game.
-Our induction hypothesis is that the positions of both the games are same, and if $S_{i}$ wins after $t$ rounds, then $S$ also wins after $t$ rounds.
+Our induction hypothesis is that the positions of both the games are the same, and if $S_{i}$ wins after $t$ rounds, then $S$ also wins after $t$ rounds.
 Base case: $D$ proposes a bijective function $f:V(G)\xrightarrow{}V(H)$.
-Note that the bijection must be color-preserving; otherwise, $S$ wins in the first round.
+Note that the bijection must be colour-preserving; otherwise, $S$ wins in the first round.
 Thus, we can assume that $f$ is color-preserving.
@@
 Now, $S_{i}$ plays a pair of pebbles in $(G_{i},H_{i})$, and $S$ also plays the same pair of pebbles in the game $(G,H)$.
-It is easy to see that both games’ positions are the same.
-Also, if $S_{i}$ wins, then the number of vertices of a particular color is different.
+It is easy to see that the positions of the two games are the same.
+Also, if $S_{i}$ wins, then the number of vertices of a particular colour is different.
 Hence, $S$ also has a winning strategy.
@@
 If $S_{i}$ chooses to remove a pebble, so will $S$.
-Again, the position of both the games is same.
-Now, if $S_{i}$ decides to play a pair of pebbles, then $S$ also decides to play a pair of pebbles.
-So, $D$ proposes a bijective function and $D_{i}$ proposes a restricted bijective function.
+Again, the position of both the games is the same.
+If $S_{i}$ decides to play a pair of pebbles, then $S$ also decides to play a pair of pebbles.
+So, $D$ proposes a bijective function, and $D_{i}$ proposes a restricted bijective function.
 Now, if $S_{i}$ plays a pair of pebbles at $(v,f_{i}(v))$, then $S$ also decides to play a pair of pebbles at $(v,f(v))$.
-Thus, the position of the game is same in both of the games.
+Thus, the position of the game is the same in both of the games.
 This ensures that if $S_{i}$ has won, then $S$ also wins.
@@
 Proof.
-Let $G$ and $H$ be the subgraphs induced on a $r$-hop neighborhood.
+Let $G$ and $H$ be the subgraphs induced on a $r$-hop neighbourhood.
 Let $(S,D)$ be the Spoiler-Duplicator pair for the game $(G,H)$.
@@
 We claim that if $S$ has a winning strategy in the game $(G,H),$ then there exists $S_{i}$ such that it has a winning strategy in the game $(G_{i},H_{i}).$ Here, the strategy of $D$ is copied by $D_{i}$ and the strategy of $S_{i}$ is copied by $S.$ We prove the lemma using induction on the number of rounds of the game.
-Our induction hypothesis is that the position of the game $(G,H)$ is same for $(G_{i},H_{i}),$ for all $i,$ if we restrict it to the subgraph induced by the vertices of color $i$ and $(i+1).$ Also, if $S$ wins after round $t,$ then there exists $S_{i}$ that wins after $t$ rounds.
+Our induction hypothesis is that the position of the game $(G,H)$ is the same for $(G_{i},H_{i}),$ for all $i,$ if we restrict it to the subgraph induced by the vertices of color $i$ and $(i+1).$ Also, if $S$ wins after round $t,$ then there exists $S_{i}$ that wins after $t$ rounds.
 Base case: $D$ proposes a bijective function $f:V(G)\longrightarrow V(H)$.
-Note that the bijection must be color-preserving; otherwise $S$ wins in the first round.
+Note that the bijection must be colour-preserving; otherwise, $S$ wins in the first round.
 Thus, we can assume that $f$ is color-preserving.
@@
 Now, $S$ will play a pair of pebbles in the game $(G,H).$ Suppose $S$ plays the pebbles at $(v,f(v))$ and $color(v)=i,$ then $S_{i}$ and $S_{i-1}$ play pebbles at $(v,f_{i}(v))$ in their first round.
-It is easy to see that the position of the games $(G,H)$ and $(G_{i},H_{i}),$ for all $i\in[r],$ is same if we restrict it to the subgraph induced by vertices of colors $i$ and $(i+1)$.
-Also, if $S$ wins, then the number of vertices of particular color are not same.
+It is easy to see that the position of the games $(G,H)$ and $(G_{i},H_{i}),$ for all $i\in[r],$ is the same if we restrict it to the subgraph induced by vertices of colours $i$ and $(i+1)$.
+Also, if $S$ wins, then the number of vertices of a particular colour is not the same.
 So, there exists some $i,$ such that $S_{i}$ also has a winning strategy.
-By induction hypothesis, assume that after the $t^{th}$ round, $S$ did not win and position of the game is same as defined.
+By induction hypothesis, assume that after the $t^{th}$ round, $S$ did not win, and the position of the game is the same as defined.
 Consider the $(t+1)^{th}$ round in both the games.
@@
 If $S$ chooses to remove a pebble from vertex $(v,f(v)),$ then, if $v$ is colored with color $i,$ then $S_{i}$ and $S_{i-1}$ will remove a pebble from vertex $(v,f_{i}(v))$.
-Again, the position of both the games is same.
+Again, the position of both the games is the same.
 Now, if $S$ decides to play a pair of pebbles, then each $S_{i}$ also decides to play a pair of pebbles.
-So, $D$ propose a bijective function and $D_{i}$ proposes a restricted bijective function.
+So, $D$ propose a bijective function, and $D_{i}$ proposes a restricted bijective function.
 Now, suppose $S$ plays a pair of pebbles at $(v_{1},f(v_{1}))$.
 If $color(v_{1})=i,$ then $S_{i}$ and $S_{i-1}$ also decides to play pebbles at $(v_{1},f_{i}(v_{1}))$.
-Thus, the position of the game is same as defined.
+Thus, the position of the game is the same as defined.
 Now, if $S$ wins, then there exists $u$ and $v$ such that either $(u,v)\in E(G)$ and $(f(u),f(v))\notin E(H)$ or $(u,v)\notin E(G)$ and $(f(u),f(v))\in E(H)$.
-Similarly, there exists $S_{i}$ for which these things happen as the position induced is same.
+Similarly, there exists $S_{i}$ for which these things happen as the position induced is the same.
 Therefore, $S_{i}$ wins for some $i.$
-Thus, from above two lemmas we can say that the expressive power of Layer $k-$WL is the same as local $k-$WL.
+Thus, from the above two lemmas, we can say that the expressive power of Layer $k$-WL is the same as local $k$-WL.
 6 Subgraph Counting Algorithms and Characterization of Patterns
 Here, we characterize the expressive power of the proposed methods in terms of subgraph as well as induced subgraph counting.
-In this section, we provide algorithms and characterization of subgraphs that can exactly count the number of patterns appearing as subgraph or induced subgraph.
+In this section, we provide algorithms and characterization of subgraphs that can exactly count the number of patterns appearing as subgraphs or induced subgraphs.
 As described above, we can see that the running time is dependent on the size of the local substructure and the value of $k$.
@@
 So, we have to take a $r$-hop neighbourhood for each vertex $v$ in the host graph $G$.
-In Section 6.1, we show how the value of $k$ can be decided based on the local substructure of the host graph.
+In 6.1, we show how the value of $k$ can be decided based on the local substructure of the host graph.
 It is independent of the structure of the pattern.
-Also, it gives an upper bound on the value of $k$ that can count pattern appearing as subgraph and induced subgraph.
-In Section 6.2, we first show that doing local count is sufficient for induced subgraph count and later we give the upper bound on $k$ based on the pattern size.
-Note that the value of $k$ for induced subgraph count is based only on the size of pattern not its structure.
-In Section 6.3, we again show that locally counting subgraph is sufficient.
-Also, we explore the value of $k$ based on the structure of pattern.
-For subgraph counting the structure of pattern can be explored to get a better upperbound for the value of $k$.
-Later, for the sake of completeness we give algorithm to count triangle, pattern of radius one and $r.$
+Also, it gives an upper bound on the value of $k$ that can count patterns appearing as subgraphs and induced subgraphs.
+In 6.2, we first show that doing local count is sufficient for induced subgraph count, and later, we give the upper bound on $k$ based on the pattern size.
+Note that the value of $k$ for induced subgraph count is based only on the size of the pattern, not its structure.
+In 6.3, we again show that locally counting subgraph is sufficient.
+Also, we explore the value of $k$ based on the structure of the pattern.
+For subgraph counting, the structure of the pattern can be explored to get a better upper bound for the value of $k$.
+Later, for the sake of completeness, we give an algorithm to count triangles, the pattern of radius one and $r.$
 6.1 Deciding k based on local substructure of host graph
-Here, we explore the local substructure of the host graph in which we are counting pattern appearing as graphs and subgraphs.
-For a given pattern of radius $r$, we explore $r$-hop neighbourhood around every vertex $v$ in the host graph $G$.
-If two graphs $G_{1}$ and $G_{2}$ are isomorphic then the number of subgraphs and induced subgraphs of both the graphs are the same.
-We use the same idea to count number of subgraphs.
-Cai et.
-al.
-Cai et al.
-(1992), shown that $\Omega(n)$ dimension is needed to guarantee graph isomorphism.
+Here, we explore the local substructure of the host graph in which we are counting patterns appearing as graphs and subgraphs.
+For a given pattern of radius $r$, we explore the $r$-hop neighbourhood around every vertex $v$ in the host graph $G$.
+If two graphs $G_{1}$ and $G_{2}$ are isomorphic, then the number of subgraphs and induced subgraphs of both the graphs are the same.
+We use the same idea to count the number of subgraphs.
+cai1992optimal, shown that $\Omega(n)$ dimension is needed to guarantee graph isomorphism.
 However, for restricted graph classes, we can still guarantee isomorphism for small dimensions.
-It has been shown that ${3-WL}$ is sufficient for planar graph Kiefer et al.
-(2019), ${k-WL}$ for graphs with treewidth at most $k$ Kiefer and Neuen (2019), $(3k+4)-WL$ for graphs with rankwidth at most $k$ Grohe and Neuen (2021), and $(4k+3)$ for graphs with Euler genus at most $k$ Grohe and Kiefer (2019).
+It has been shown that 3WL{3-WL} is sufficient for planar graph (kiefer2019weisfeiler), kWL{k-WL} for graphs with treewidth at most $k$ (decompose_graphs), (3k+4)WL(3k+4)-WL for graphs with rankwidth at most $k$ (rankwidth_WL), and $(4k+3)$ for graphs with Euler genus at most $k$ (grohe2019linear).
 We say these graph classes as good graph classes.
 Note that, for non-isomorphic graphs, the graphs is not $k$-WL equivalent.
-Thus, running corresponding $k-$WL can count pattern of radius $r$, appearing as subgraph and induced subgraph.
+Thus, running corresponding $k$-WL can count the pattern of radius $r$, appearing as a subgraph and induced subgraph.
 Theorem 4.
@@
   1. 1.
-    $(3-WL)$ if $G_{v}^{r}$ planar
+    (3WL)(3-WL) if $G_{v}^{r}$ planar
   2. 2.
-    $(k-WL)$ if $tw(G_{v}^{r})\leq k$***tw denotes treewidth
+    (kWL)(k-WL) if $tw(G_{v}^{r})\leq k$111tw denotes treewidth
   3. 3.
-    $((3k+4)-WL)$ if $rankwidth(G_{v}^{r})\leq k$
+    ((3k+4)WL)((3k+4)-WL) if $rankwidth(G_{v}^{r})\leq k$
   4. 4.
-    $((4k+3)-WL)$ if $Euler-genus(G_{v}^{r})\leq k$
+    ((4k+3)WL)((4k+3)-WL) if $Euler-genus(G_{v}^{r})\leq k$
 Proof.
-Consider the subgraph induced by the vertex $v$ and its $r$-hop neighborhood in $G,$ say $G_{v}^{r},$ and the subgraph induced by the vertex $u$ and its $r$-hop neighborhood in $H,$ say $H_{u}^{r}.$ Suppose both structures belong to good graph classes.
+Consider the subgraph induced by the vertex $v$ and its $r$-hop neighbourhood in $G,$ say $G_{v}^{r},$ and the subgraph induced by the vertex $u$ and its $r$-hop neighbourhood in $H,$ say $H_{u}^{r}.$ Suppose both structures belong to good graph classes.
 Now, we run corresponding $k$ based on the local substructure as mentioned in the theorem.
-If the color histogram of the stable color matches for both the graphs.
+If the colour histogram of the stable colour matches for both graphs.
 This implies that both the graphs are isomorphic.
 Thus, the number of subgraphs and induced subgraphs in both of the substructures are also the same.
-Also, we run respective $k-$WL on a colored graph, where vertices at a distance $i$ from $v$ is colored $i$.
-So, it is at least as expressive as running $k-$WL on an uncolored graph.
+Also, we run respective $k$-WL on a coloured graph, where vertices at a distance $i$ from $v$ are coloured $i$.
+So, it is at least as expressive as running $k$-WL on an uncoloured graph.
 We can also show that it is strictly more expressive in distinguishing non-isomorphic graphs.
@@
 Corollary 2.
-Running $1-$WL guarantees the exact number of subgraphs and induced subgraphs of all patterns of radius one, when the maximum degree of the host graph is bounded by 5.
+Running 11-WL guarantees the exact number of subgraphs and induced subgraphs of all patterns of radius one, when the maximum degree of the host graph is bounded by 5.
 Similarly, if the maximum degree of the host graph is bounded by $15$, then running $2$-WL is sufficient to count subgraphs and induced subgraphs of all patterns with a dominating vertex.
@@
 It is sufficient to compute $IndCount_{(u,v)}(H,G_{v}^{r}),$ for $i\in[n],$ where $G_{v}^{r}$ is induced subgraph of $r$-hop neighborhood vector of $v$.
-The following theorem gives a direct comparison with the $S_{k}$ model Papp and Wattenhofer (2022a), where each node has an attribute which is the count of induced subgraphs of size at most $k.$
+The following theorem gives a direct comparison with the $S_{k}$ model (papp2022theory), where each node has an attribute which is the count of induced subgraphs of size at most $k.$
 Theorem 6.
@@
 We will choose one vertex $v$ as a key vertex.
-Now, we want to count $P-v$ locally.
+Now, we want to count PvP-v locally.
 Assume that the distance from $v$ to any vertex is $^{\prime}r^{\prime}$.
@@
 The $k$-WL and $C^{k+1}$ are equivalent, so we have to add the unary relation putting the condition of the required colors.
-Again, using Lemma 3 we can say that running $k-$WL locally can count all patterns of size $(k+2)$ appearing as an induced subgraph.
+Again, using 3 we can say that running $k$-WL locally can count all patterns of size $(k+2)$ appearing as an induced subgraph.
 ∎
-We can see the corollary below in which we mention the set of patterns shown in Chen et al. (2020).
+We can see the corollary below in which we mention the set of patterns shown in (chen2020can).
 Corollary 3.
@@
 We can make informed choices for the values of $k$ and $r.$ Notice that the value of $k$ is chosen based on the local substructure.
-Also, the value of $r$ is the radius of $H.$ Suppose the local substructure is simple (planar, bounded treewidth, bounded rankwidth Theorem 4).
-In that case, $k-$WL, for small values of $k,$ is sufficient for counting induced subgraph $H.$ Otherwise, we have to run $(|H|-2)-$WL in the worst case.
+Also, the value of $r$ is the radius of $H.$ Suppose the local substructure is simple (planar, bounded treewidth, bounded rankwidth 4).
+In that case, $k$-WL, for small values of $k,$ is sufficient for counting induced subgraph $H.$ Otherwise, we have to run (|H|2)(|H|-2)-WL in the worst case.
 6.3 Deciding k based on the pattern for counting subgraphs
@@
 $Count(H,G)=\sum_{H^{\prime}}\alpha_{H^{\prime}}Hom(H^{\prime},G)$ (3)
-where $H^{\prime}$ ranges over all graphs in $H^{\prime}$. This equation has been used to count subgraphs by many authors (Please refer Alon et al. (1997); Curticapean et al. (2017)).
+where $H^{\prime}$ ranges over all graphs in $H^{\prime}$. This equation has been used to count subgraphs by many authors (Please refer (Alon1997; curticapean2017homomorphisms)).
 Theorem 7.
-Cai et al. (1992); Dell et al. (2018) For all $k\geq 1$ and for all graphs $G$ and $H$, the following are equivalent:
+(cai1992optimal; Dell2018LovszMW) For all $k\geq 1$ and for all graphs $G$ and $H$, the following are equivalent:
   1. 1.
@@
   3. 3.
-    Graph $G$ and $H$ are $C^{k+1}$ equivalent***Counting logic with (k+1) variables.
-Using Equation 2 and Theorem 7, we arrive at the following theorem:
+    Graph $G$ and $H$ are $C^{k+1}$ equivalent [^2] Counting logic with (k+1) variables.
+Using 3 and 7, we arrive at the following theorem:
 Theorem 8.
@@
 Thus, counting locally, we can easily see that every subgraph would be counted.
-Now to prove Equation 3, it is sufficient to show that for a given subgraph it is counted exactly $|Orbit_{H}(u)|$ many times.
+Now to prove 4, it is sufficient to show that for a given subgraph it is counted exactly $|Orbit_{H}(u)|$ many times.
 Note that two subgraphs are same if and only if their vertex sets and edge sets are same.
@@
 ∎
-Using Theorem 8 and Lemma 5, one can easily see that for counting pattern $H$ as a subgraph, it is sufficient to run Local $k-$WL on the local substructure and count the subgraph locally.
+Using 8 and 5, one can easily see that for counting pattern $H$ as a subgraph, it is sufficient to run Local $k$-WL on the local substructure and count the subgraph locally.
 Theorem 9.
 Local $k$-WL can exactly count any subgraph $H$ if $htw(H-v)\leq k$.
-The upper bound on the choice of $k$ for running $k-$WL can be improved from the default $|H|-2$ bound that we used for the induced subgraph count.
+The upper bound on the choice of $k$ for running $k$-WL can be improved from the default |H|2|H|-2 bound that we used for the induced subgraph count.
 The value of $k$ is now upper bound by $htw(H)$.
@@
 Corollary 4.
-Local $1-$WL can exactly count the number of patterns $P$ or $P-v$ appearing as a subgraph, when $P$ or $P-v,$ with a dominating vertex $v$, is $K_{1,s}$ and $2K_{2}.$
+Local $1$-WL can exactly count the number of patterns $P$ or PvP-v appearing as a subgraph, when $P$ or Pv,P-v, with a dominating vertex $v$, is $K_{1,s}$ and $2K_{2}.$
 Proof.
@@
 Also, for $2K_{2}$, the homomorphic image is either itself or the star.
-So, the treewidth of all its homomorphic images is $1$
+So, the treewidth of all its homomorphic images is $1.$
 Corollary 5.
@@
 Proof.
-We can see that $H-v$ is $K_{1,2}$ when choosing any vertex as a key vertex.
+We can see that HvH-v is $K_{1,2}$ when choosing any vertex as a key vertex.
 Also, the orbit size is $4$.
@@
 Proof.
-For subgraph counting, we can see that for all of the $5$ patterns, there exists a vertex $v$ such that $htw(P-v)=1.$ One can note that the attributed triangle can also be handled using Corollary 1.
-Since every pattern has a dominating vertex, running $1-$WL on the subgraph induced on the neighborhood is sufficient.
+For subgraph counting, we can see that for all of the $5$ patterns, there exists a vertex $v$ such that htw(Pv)=1.htw(P-v)=1.
+One can note that the attributed triangle can also be handled using Corollary 1.
+Since every pattern has a dominating vertex, running $1$-WL on the subgraph induced on the neighbourhood is sufficient.
 Now, we only have to argue for the patterns appearing as induced subgraphs.
@@
 Thus, Local $1$-WL can give the exact count of the number of triangles.
-For more details, please see 6.4.
+For more details, please see 6.4.1.
 The running time of $1$-WL depends on the number of iterations, $h$.
@@
 Lemma 6.
-It requires $O(n)$ time to guarantee the count of patterns which can be written using $2$-variable with a counting quantifier where the depth of the quantifier is constant.
-For more details, please see the equivalence between the number of iterations and quantifier depth in Theorem 2.
-A list of patterns that can be counted as subgraph and induced subgraphs using local $1-$WL and local $2-$WL are mentioned in Table 1.
-Also, the patterns including 3-star, triangle, chordal $4$-cycle (chordal $C_{4}$), and attributed triangle have been studied in Chen et al.
-(2020) and have been shown that it cannot be counted by $1-$WL.
+It requires $O(n)$ time to guarantee the count of patterns which can be written using 22-variable with a counting quantifier where the depth of the quantifier is constant.
+For more details, please see the equivalence between the number of iterations and quantifier depth in 2.
+A list of patterns that can be counted as subgraph and induced subgraphs using local $1$-WL and local $2$-WL are mentioned in 1.
+Also, the patterns including 3-star, triangle, chordal 44-cycle (chordal $C_{4}$), and attributed triangle have been studied in (chen2020can) and have been shown that it cannot be counted by 11-WL.
 Table 1: List of all the patterns that can be counted exactly (as a subgraph or induced subgraph), given $G$, using Local $k$-WL, for different $k$.
-Restriction on $G_{v}^{r}$ k Patterns, $H$
-Induced
-subgraph
-Subgraph Reference
-$G_{v}^{r}$ is amenable 1
-All
-Corollary 1
-Max degree $\leq$ 5 1
-Patterns with a dominating vertex
-Corollary 2
-Max degree $\leq$ 15 2
-Pattern with a dominating vertex
-Corollary 2
-No restriction 2
-3-star, triangle, tailed triangle, chordal cycle,
-attributed triangle
-Corollary 3
-No restriction 1
-Either $H$ or $H-v$ is $K_{1,s}$ or $2K_{2},$ where $v$ is the
-dominating vertex
-Corollary 4
+Restriction on $G_{v}^{r}$ k Patterns, $H$ Induced subgraph Subgraph Reference
+$G_{v}^{r}$ is amenable 1 All Corollary 1
+Max degree $\leq 5$ 1 Patterns with a dominating vertex Corollary 2
+Max degree $\leq 15$ 2 Pattern with a dominating vertex Corollary 2
+No restriction 2 3-star, triangle, tailed triangle, chordal cycle, attributed triangle Corollary 3
+No restriction 1 Either $H$ or HvH-v is $K_{1,s}$ or $2K_{2},$ where $v$ is the dominating vertex Corollary 4
 No restriction 1 $C_{4}$ Corollary 5
-No restriction 1
-3-star, tailed triangle, chordal cycle
-Corollary 6
-No restriction 1
-triangle, attributed triangle
-Corollary 6
+No restriction 1 3-star, tailed triangle, chordal cycle Corollary 6
+No restriction 1 triangle, attributed triangle Corollary 6
 6.4 Algorithms for subgraph counting
-Triangle counting in the host graph
-We describe an algorithm for counting the number of triangles in a given host graph $G$ in Algorithm 5.
-Note that counting the number of triangles is the same as counting the number of edges in the subgraph induced by $N_{G}(v).$ It is well known that two $1-$WL equivalent graphs have the same number of edges.
-This ensures that if we run $1-$WL on the induced subgraphs in the neighborhood of $v$, taking color as a feature, we can guarantee the count of the triangles.
-On the other hand, we can see that running $1$- WL on graph $G$ will not guarantee the count of triangles.
-Running $1-$WL on the entire graph takes $O(n+m)\log(n)$ and $O(n)$ space, where $m$ is the number of edges.
-Thus, running $1-$WL locally in the neighborhood is more space and time efficient.
-Note that the running time is also dependent on the number of iterations, $h.$ Running $1-$WL for $h-$ iteration requires $O(nh)$ time.
-The quantifier depth of counting logic with $(k+1)$ variables is equivalent to the number of iterations of $k-$WL ( See Theorem 2).
-For the case of triangle counting, we just need to count the number of edges, which can be done by running just one iteration of $1-$WL.
+6.4.1 Triangle counting in the host graph
+We describe an algorithm for counting the number of triangles in a given host graph $G$ in 5.
+Note that counting the number of triangles is the same as counting the number of edges in the subgraph induced by $N_{G}(v).$ It is well known that two 11-WL equivalent graphs have the same number of edges.
+This ensures that if we run $1$-WL on the induced subgraphs in the neighbourhood of $v$, taking colour as a feature, we can guarantee the count of the triangles.
+On the other hand, we can see that running 11- WL on graph $G$ will not guarantee the count of triangles.
+Running 11-WL on the entire graph takes $O(n+m)\log(n)$ and $O(n)$ space, where $m$ is the number of edges.
+Thus, running $1$-WL locally in the neighbourhood is more space and time-efficient.
+Note that the running time is also dependent on the number of iterations, $h.$ Running 11-WL for $h$- iteration requires $O(nh)$ time.
+The quantifier depth of counting logic with $(k+1)$ variables is equivalent to the number of iterations of $k$-WL ( See 2).
+For the case of triangle counting, we just need to count the number of edges, which can be done by running just one iteration of 11-WL.
 So, the time required is $O(deg(v))$ for each $v.$ This can be done in parallel for each vertex.
@@
 1:Let $G$ be the host graph.
-2:$num\_edges=0$
+2:num_edges=0num\_edges=0
 3:for each vertex $v$ in $V(G)$ do
@@
 5:  Find the number of edges in the induced subgraph on $N_{G}(v)$
-6:     Add it to $num\_edges$
+6:  Add it to num_edgesnum\_edges
 7:end for
-8:$c=num\_edges/3$
+[…]
 9:Output: The number of triangles in graph $G$ is $c$.
-Counting subgraph of radius one
+6.4.2 Counting subgraph of radius one
 We begin by explaining the procedure for counting the number of subgraphs having a dominating vertex (radius one).
@@
 If a subgraph exists, then the dominating vertex must be mapped to some vertex.
-We iteratively map the dominating vertex to each vertex in the host graph and count the number of patterns in the neighborhood of the dominating vertex.
-Here we present an algorithm for counting patterns of radius one Algorithm 6.
+We iteratively map the dominating vertex to each vertex in the host graph and count the number of patterns in the neighbourhood of the dominating vertex.
+Here, we present an algorithm for counting patterns of radius one 6.
 Note that running $k$-WL on the entire graph takes $O(k^{2}\cdot n^{k+1}\log n)$ time and $O(n^{k})$ space, whereas when we run locally, it requires less time and space.
-Suppose we run only on the neighborhood of each vertex.
+Suppose we run only on the neighbourhood of each vertex.
 Then, it requires $\sum_{v\in V(G)}(deg(v))^{k+1}\log(deg(v))$ and space $O(max_{i}(deg(v_{i}))^{k}+n).$ More specifically, suppose the given graph is $r$-regular.
@@
 3:  Find the induced subgraph on $N_{G}(v)$
-4:     if degree$(v)+1<|V(H)|$ then
+4:  if degree(v)+1<|V(H)|(v)+1<|V(H)| then
 5:   skip this iteration
@@
 10:return $\frac{\sum_{v\in V(G)}Count_{u,v}(H,G_{v}^{r})}{|Orbit_{H}(u)|}$
-Counting subgraphs of radius r
-Here, in Algorithm 7, we describe how to count the number of subgraphs of radius $r.$ We iterate over all the vertices and take the $r$-hop neighborhood around that vertex, say $v,$ and choose a suitable $k$ according to the structure of the pattern that can guarantee the count of subgraphs in the local substructure.
+6.4.3 Counting subgraphs of radius r
+Here, in 7, we describe how to count the number of subgraphs of radius $r.$ We iterate over all the vertices and take the $r$-hop neighbourhood around that vertex, say $v,$ and choose a suitable $k$ according to the structure of the pattern that can guarantee the count of subgraphs in the local substructure.
 Algorithm 7 Counting the number of subgraphs
@@
 5:  for each layer of the BFS tree rooted at $v$ do
-6:         if ##\## vertices at layer $i$ from $v$ $<$ ##\## vertices at layer $i$ from $u$ then
+6:   if #\# vertices at layer $i$ from $v$ << #\# vertices at layer $i$ from $u$ then
 7:     $c=0$
@@
 22:return $\frac{\sum_{v\in V(G)}Count_{u,v}(H,G_{v}^{r})}{|Orbit_{H}(u)|}$
-7 Fragmentation
-Here, we discuss the Fragmentation technique that is different from the localized methods we have seen so far.
-From Table 1, we have seen that Local $k-$WL (or ((Local + Layer) $k-$WL)) is sufficient for getting an exact count for the patterns given in the table.
-Given a pattern $P,$ that is more complicated than the patterns in Table 1, we fragment $P$ into simpler patterns such that their exact count is known.
-In the subgraph GNN proposed earlier, look into subgraph of the host graph.
-We have seen that this technique is scalable on large graphs.
-Also, we have seen that subgraph GNN is more expressive and efficient than traditional GNN.
-So, we tried to explore the expressibility when the pattern is also fragmented into smaller subpatterns.
-The fragmentation method involves fragmenting the pattern, $P$, into smaller subpatterns and counting these subpatterns to get the count of the $P$ in the host graph.
-As described in Section 6, the value of $k$ depends on the size of the pattern (induced subgraph) and its structure (subgraph count).
-Thus, even though the $htw(P)$ may be large, if we divide it into subpatterns, then $k$ is required to guarantee the count would be reduced.
-Thus, it provides more expressiveness for smaller $k$ in order to count the patterns that cannot be counted if we directly apply Local $k-$WL.
-Thus, given a pattern, we apply the same fragmentation on $G_{v}^{r}.$ Thus, the number of occurrences of $H$ in $G_{v}^{r}$ can be computed by combining the counts of the simpler patterns.
-Instead of training a GNN for counting $H,$ we can design GNNs for learning the easier tasks (i.e., for counting the simpler models) and combine the outputs of those models.
-It should be noted that the fragmentation into smaller subgraphs depends on the structure of the pattern $H.$
-We demonstrate this technique for counting induced tailed triangles in Figure 2.
-As seen in the figure, the tailed triangle pattern can be fragmented into two parts : the pendant vertex as the key vertex, and an attributed triangle.
-The colors assigned to the nodes of the attributed triangle depend on the distance of the nodes from the key node.
-Thus, the task of counting tailed triangles reduces to counting attributed triangles, as all the vertices at level 1 are connected to the root.
-Suppose the task is to count the number of chordal cycles appearing as induced subgraphs.
-If we pick the vertex of degree three as the key vertex, then it is enough to search the neighbourhood of $v$ in the host graph.
-Now, in $N_{G}(v),$ if we count the number of $2-stars,$ then it gives the count of chordal cycle appearing as subgraph.
-If we eliminate the appearance of $K_{4}$, then it would give the exact count of chordal cycles appearing as induced subgraphs.
-In that case, we count the number of triangles in $N_{G}(v),$ which gives the exact count of $K_{4}$.
-Using the fragmentation technique, we show that just $1-$WL is sufficient for getting exact counts of induced subgraphs of certain sizes.
-Figure 2: Fragmentation for counting tailed triangles.
-Idea of the fragmentation algorithm 8: Given a graph $G$ and a pattern $P$, we first fix a vertex $u\in V(P)$ as the key vertex.
-Now, assume that the radius of the pattern is $r$.
-Thus, for counting $P$ locally, it is sufficient to take the $r$-hop neighbourhood for each vertex $v$ of $G,$ say $G_{v}^{r},$ as has been shown in Lemma 3 and Lemma 5.
-Also, we have proved above that doing count locally is sufficient for both subgraph and induced subgraph.
-Now, we fragment pattern $P$ into smaller subpatterns, say $P_{1},P_{2},\ldots P_{l}$.
-Based on the structure of $P$, we consider the subgraphs of $G_{v}^{r}$ where the subpattern $P_{i}$ is required to be counted.
-For all subpattern $P_{i}$ in $P$, we make a list of subgraphs $G_{v}^{r}(1),G_{v}^{r}(2),\ldots G_{v}^{r}(t)$ of $G_{v}^{r}$ where $P_{i}$ needs to be counted.
-We aggregate these lists into a dictionary, $\mathcal{L},$ with $P_{i}$ as the keys.
-It should be noted that the decomposition of $P$ into $P_{i}$s is such that their counts can be calculated.
-That is, we have learnt models $M_{i},$ corresponding to each $P_{i}$, which counts the number of subpatterns $P_{i}$.
-The array $c$ stores the count of $P_{i}$’s in each subgraph of $G_{v}^{r}$.
-Now, for each vertex, we use appropriate functions $\alpha$ and $\beta$ to combine the counts in $c$ to get the count of $P$ in $G_{v}^{r}$.
-Finally, the function $\gamma$ finds the normalizing factor to get the actual count of the pattern in the $G.$
+7 Methodology
+This section describes our unified subgraph counting strategy, combining localized Weisfeiler-Leman (WL) algorithms, differentiable learning, and a novel fragmentation approach.
+The framework is broken into three primary components: (A) Pattern Learning, (B) Local Count Learning, and (C) Global Count Learning, as illustrated in Figure 2.
+Figure 2: Overview of the framework.
+(A) Pattern Learning: Each node’s $r$-hop neighbourhood, with directed edges marking the root, is input to a model that predicts relevant subpatterns.
+(B) Local Count Learning: Using predicted patterns, the model estimates subpattern counts per root node.
+(C) Global Count Learning: The total count is normalized to adjust for possible overcounting.
+7.1 Pattern Learning
+For a given subgraph pattern $P$ (e.g., triangle, $3$-star, or tailed triangle), the local subgraph $G_{v}^{r}$ centered at node $v$ must be preprocessed to match $P$’s structure.
+We augment $G_{v}^{r}$ by directing edges outward from $v$ to clearly indicate the root.
+Modified $G_{v}^{r}$ is passed to a GNN, which produces node and edge embeddings.
+These embeddings are used to classify edges as included/excluded based on their relevance to the target pattern, yielding a pruned subgraph suited as input to the next component.
+7.2 Local Count Learning
+Each root node’s updated local subgraph (from Pattern Learning) is used to estimate the "local count"—the number of occurrences of a subpattern in $G_{v}^{r}$.
+The local counting rule depends on the motif: for triangles, count edges among neighbours; for $k$-stars, compute $\binom{d_{v}}{k}$ where $d_{v}$ is node $v$’s degree.
+For arbitrary patterns, ground truth local counts are generated as supervision.
+The model is trained to predict these counts using a GNN applied to the modified $G_{v}^{r}$.
+7.3 Global Count Learning
+Summing local counts across all roots may overcount global subgraph occurrences due to overlaps.
+Hence, a final normalization step is applied.
+We use a learnable normalization function to infer the correction factor needed to yield the true global count, enabling an end-to-end differentiable pipeline for subgraph counting.
+7.4 Fragmentation
+For complex patterns beyond primitives like triangles or $k$-stars, we introduce the fragmentation technique.
+Instead of directly counting $P$, which may be computationally intensive, we recursively decompose $P$ into simpler subpatterns ("fragments") whose counts can be efficiently determined.
+The total count of $P$ is obtained by suitably combining the counts of these fragments, often requiring only small values of $k$ in the local $k$-WL hierarchy.
+For example, the tailed triangle can be fragmented into a pendant vertex (the key/root) and an attributed triangle, as shown in Figure 3.
+For chordal $C_{4}$ cycles, the pattern is broken into $2$-stars and triangles within the neighbourhood of a key vertex, adjusting counts to avoid overcounting $K_{4}$s.
+The approach applies for any pattern whose complexity can be reduced through such decomposition.
+Figure 3: Fragmentation example for a tailed triangle: identifying fragments (isolated node, edge) in the subgraph surrounding a root node.
+Fragmentation Algorithm:
+Given a host graph $G$ and a pattern $P$ to be counted, select a key vertex in $P$ and set the radius $r$.
+For each node $v\in V(G)$, extract the $r$-hop neighbourhood $G_{v}^{r}$.
+Decompose $P$ into subpatterns $\{P_{1},...,P_{l}\}$, each counted by a learned model $M_{i}$.
+Store local counts in arrays, aggregate via learnable combination functions $\alpha$, $\beta$ (for subpattern and root aggregation, respectively), and finally apply a normalization function $\gamma$ to compute the global count.
 Algorithm 8 Fragmentation Algorithm
-1:Let $G$ be the host graph and $P$ be the list of patterns,
-$\mathcal{L}$ be the dictionary of subgraph rules associated with the subpattern $P_{i},$
-3:$M=\{M_{1},\ldots,M_{l}\}$ the list of learnt models for counting $P_{i}$’s, where $l=|P|$
-4:$a\leftarrow$ Zero array of size $|V(G)|$
-5:for each vertex $v$ in $V(G)$ do
-6:     Extract $G_{v}^{r}$
-7:     $b\leftarrow$ Zero array of size $l$
-8:     for each pattern $P_{i}$ in $\mathcal{L}$ do
-9:         $c\leftarrow$ Zero array of size $s$, where $s=|\mathcal{L}(P_{i})|$
-10:         for each rule $k$ in $\mathcal{L}(P_{i})$ do
-11:              Extract $G_{v}^{r}(k)$
-12:              $c[k]$ = $M_{i}(G_{v}^{r}(k))$
+1:$G$; $\mathbb{P}$: List of subpatterns; $M^{\text{pattern}}_{i}$: learned model for generating pattern corresponding to $P_{i}\in\mathbb{P}$; $M^{\text{count}}_{i}$: learned model for counting $P_{i}\in\mathbb{P}$;
+[…]
+3:for each node $v\in V(G)$ do
+4:  $H_{v}=G_{v}^{r}\setminus\{v\}$
+5:  $b\leftarrow[]$
+6:  for each pattern $P_{i}\in\mathbb{P}$ do
+7:   $c\leftarrow[]$
+8:   for each node $u\in H_{v}$ do
+9:     $\mathcal{P}_{u}=M^{\text{pattern}}_{i}(H_{v},u)$
+10:     $c.append(M^{\text{count}}_{i}(\mathcal{P}_{u}))$
+11:   end for
+12:   $b.append(\alpha(c))\qquad\#$Learnable function
 13:  end for
-14:         $b[i]$ = $\alpha(c)$
+14:  $a.append(\beta(b))\qquad\#$Learnable function
 15:end for
-16:     $a[v]$ = $\beta(b)$
-17:end for
-18:$Count(P,G)$ = $\gamma(a)$
-19:return $Count(P,G)$
+[…]
+17:return CountCount
 Theorem 10.
-Using the fragmentation method, we can count induced tailed triangle, chordal $C_{4}$ and $3-star,$ by running Local $1-$WL.
+Using the fragmentation method, we can count induced tailed triangle, chordal $C_{4}$ and 3star,3-star, by running Local $1$-WL.
 Proof.
-For tailed triangle, we fix the pendant vertex as the key vertex (refer to Figure 2).
+For tailed triangle, we fix the pendant vertex as the key vertex (refer to 3).
 Now, we have to look for the count of triangles such that exactly one of the node of triangle is adjacent to the key vertex.
-We find induced subgraph of $2$-hop neighborhood for each key vertex.
+We find induced subgraph of 22-hop neighborhood for each key vertex.
 Now, we color the vertices at distance $i$ by color $i$.
 Then, the problem of counting the number of tailed triangles reduces to counting the number of colored triangles in the colored graph such that one node of triangle is colored $1$ and remaining two nodes are colored $2$.
-We can find the count of colored triangles using $1-$WL on the induced subgraph by Corollary 1.
+We can find the count of colored triangles using 11-WL on the induced subgraph by Corollary 1.
 This number of count of colored triangle is same as $IndCount_{(u,v}(tailed-triangle,G_{v}^{r})$.
-Now, using Lemma 3, we can say that fragmentation technique can count tailed triangles appearing as induced subgraphs using $1-$WL.
-Consider the pattern, chordal $C_{4}.$ We have proved that $1-$WL can count the number of subgraphs of chordal $C_{4}.$ So, to count the number of induced subgraphs of chordal $C_{4},$ we only have to eliminate the count of $K_{4}.$ When we fix one vertex of degree $3$ as key vertex, we can easily compute the count of $K_{1,2}$ in the neighborhood.
+Now, using 3, we can say that fragmentation technique can count tailed triangles appearing as induced subgraphs using 11-WL.
+Consider the pattern, chordal $C_{4}.$ We have proved that 11-WL can count the number of subgraphs of chordal $C_{4}.$ So, to count the number of induced subgraphs of chordal $C_{4},$ we only have to eliminate the count of $K_{4}.$ When we fix one vertex of degree $3$ as key vertex, we can easily compute the count of $K_{1,2}$ in the neighborhood.
 Now, we have to eliminate all three tuples appearing as triangles in the neighborhood of the key vertex.
@@
 This gives the exact count of chordal $C_{4}$ appearing as subgraph in the local structure.
-Using Lemma 3, we can find $IndCount(ChordalC_{4},G)$.
-Consider the pattern, $3-star$.
+Using 3, we can find $IndCount(ChordalC_{4},G)$.
+Consider the pattern, 3star3-star.
 Here, we choose any of the pendant vertex as key vertex.
 Now, we have to compute $K_{1,2},$ where one center vertex of the star is connected to the key vertex.
-We can easily count the number of colored $K_{1,2}$ in $2$-hop neighborhood of the key vertex.
+We can easily count the number of colored $K_{1,2}$ in 22-hop neighborhood of the key vertex.
 However, a triangle can be also be included in this count.
@@
 Again, using the approach discussed above, we can count the number of colored triangles and this will output the exact count of colored induced $K_{1,2}$.
-Again, using lemma 3, we can find $IndCount(3-star,G).$
-Theorem 11.
-Using the fragmentation technique we can count all patterns appearing as induced subgraphs of size $4$ by just running Local $1-$WL.
-Proof.
-In Table 2, we describe how the fragmentation technique can be leveraged to count all the induced subgraphs of size $4$.
-This shows that for $k=1$, the fragmentation technique is more expressive than $S_{k+3}$.
-Also, we can enlist more graphs where the larger pattern can be seen as the union of fragments of smaller patterns.
-Using this, we can see that it covers all the graphs that were mentioned in Chen et al.
-(2020).
-One can see that all the formulae use the function that can be computed by $1-$WL.
-The number of vertices and the number of edges can easily be computed after doing one iteration of $1-$WL.
-Also, the degree sequence can be computed after running $1-$WL for one iteration.
-All other functions(formulae) are just the linear combination of the functions computed earlier, the number of vertices or the number of edges.
-∎
-In the structures shown in the table 2 below, we have highlighted the key vertex by "light green" color and other vertices by "black color".
-Table 2: Fragmentation for patterns for size of at most 4.
+Again, using lemma 3, we can find IndCount(3star,G).IndCount(3-star,G).
+Table 2: Fragmentation for patterns for size of at most 4
 Vertices Structure Formula
@@
 G2 [Uncaptioned image] $|E|$
-G3 [Uncaptioned image] $\large\frac{\sum_{v}IndCount_{(u,v)}(G_{1},G-G[N_{G}[v]])}{3}$
+G3 [Uncaptioned image] $\frac{\sum_{v}IndCount_{(u,v)}(G_{1},G-G[N_{G}[v]])}{3}$
 G4 [Uncaptioned image] $\sum_{v}IndCount(G_{2},G-G[N_{G}[v]])$
@@
 G17 [Uncaptioned image] $\frac{IndCount(attributed-G_{4},G[N_{G}(v)\cup N_{G}(N_{G}(v))]}{2}$
+Theorem 11.
+Using the fragmentation technique we can count all patterns appearing as induced subgraphs of size $4$ by just running Local $1$-WL.
+Proof.
+In 2, we describe how the fragmentation technique can be leveraged to count all the induced subgraphs of size $4$.
+This shows that for $k=1$, the fragmentation technique is more expressive than $S_{k+3}$.
+Also, we can enlist more graphs where the larger pattern can be seen as the union of fragments of smaller patterns.
+Using this, we can see that it covers all the graphs that were mentioned in (chen2020can).
+One can see that all the formulae use the function that can be computed by 11-WL.
+The number of vertices and the number of edges can easily be computed after doing one iteration of 11-WL.
+Also, the degree sequence can be computed after running $1$-WL for one iteration.
+All other functions(formulae) are just the linear combination of the functions computed earlier, the number of vertices or the number of edges.
+∎
+Expressiveness Guarantee:
+Using fragmentation, we show that all patterns of up to $4$ nodes can be exactly counted using only Local $1$-WL.
+For specific motifs (tailed triangle, chordal $C_{4}$, $3$-star), the decomposition ensures algorithmic tractability and high expressiveness (see Theorems 10 and 11).
+The fragmentation technique thus enables scalable, modular, and highly expressive subgraph counting, leveraging learnable models for each subcomponent while ensuring practical feasibility and end-to-end differentiable training.
 8 Theoretical Comparison of Graph Neural Networks
-In this section we give a comparison of the time and expressiveness between the GNN hierarchies proposed in Papp and Wattenhofer (2022a) and our methods.
-From Theorem 3, it is clear that $k-$WL is less expressive than Local $k-$WL.
+n this section we give a comparison of the time and expressiveness between the GNN hierarchies proposed in (papp2022theory) and our methods.
+From 3, it is clear that $k$-WL is less expressive than Local $k$-WL.
 Also, we have shown that the space and time required by Local $k$-WL are less compared to $k$-WL.
@@
 Thus, it would require $O(t_{k}\times n_{1}^{k})$, where $t_{k}$ is the number of distinct graphs upto isomorphism of at most $k$-vertices.
-Using Theorem 6, one can easily see that running $k-$WL locally is more expressive than $S_{k+2}.$
+Using 6, one can easily see that running $k$-WL locally is more expressive than $S_{k+2}.$
 The $N_{k}$ model has a preprocessing step in which it takes the $k$-hop neighborhood around vertex $v,$ and gives attributes based on the isomorphism type.
 In a dense graph, it is not feasible to solve the isomorphism problem in general, as the size of the induced subgraph may be some function of $n$.
-Also, using the best known algorithm for graph isomorphism by Babai (2016), the time required is $O(n_{1}^{O(\log n_{1})}).$ However, running Local $k-$WL would require $O(n_{1}^{k})$.
-Also, there are rare examples of graphs that are $3-$WL equivalent and non-isomorphic.
+Also, using the best-known algorithm for graph isomorphism by babai2016graph, the time required is $O(n_{1}^{O(\log n_{1})}).$ However, running Local $k$-WL would require $O(n_{1}^{k})$.
+Also, there are rare examples of graphs that are 33-WL equivalent and non-isomorphic.
 So, if we run $3$-WL locally, then most of times expressive power matches with $N_{k}.$
-The $M_{k}$ model deletes a vertex $v$ and then runs $1-WL.$ Papp and Wattenhofer (2022a) proposed that instead of deleting the vertices, $k$ vertices are marked in the local neighborhood and showed that it is more expressive than deletion.
+The $M_{k}$ model deletes a vertex $v$ and then runs 1WL.1-WL.
+papp2022theory proposed that instead of deleting the vertices, $k$ vertices are marked in the local neighborhood and showed that it is more expressive than deletion.
 It identifies $k$ set of vertices in the local $r$-hop neighborhood of the graph.
 It would require $O(n_{1}^{(k+2)}\log(n_{1}))$ time as it has $O(n_{1}^{k})$ many possibilities of choosing $k$ many vertices.
-It requires $O(n^{2}\log n)$ time to run $1-$WL.
+It requires $O(n^{2}\log n)$ time to run 11-WL.
 The same time is required for Local $(k+1)$-WL.
-That’s why we compare $M_{k-1}$ with Local $k-$WL in Table 3.
-Also, it is known that with any $l$ markings and running $k-$WL is less expressive than running $(k+l)-$WL on the graph Fürer (2017).
-So, if we plug in the value, we can see that running Local $k-$WL is more expressive than doing $l$ marking and running $1-$WL.
+Also, it is known that with any $l$ markings and running $k$-WL is less expressive than running $(k+l)$-WL on the graph (furer2017combinatorial).
+So, if we plug in the value, we can see that running Local $k$-WL is more expressive than doing $l$ marking and running 11-WL.
 One can get an intuition by comparing with the $(k+1)$ bijective pebble game.
-If we mark the vertices, then the marking pebbles get fixed and give less power to spoiler.
+If we mark the vertices, then the marking pebbles get fixed and give less power the to spoiler.
 However, just running $k$-WL the spoiler is free to move all the pebbles.
@@
 $G_{v}^{r}$ is distinguished because there exists a tuple $(v_{1},v_{2},....v_{k-1})$ such that marking these vertices and running $1$-WL on the graph gives stabilised coloring of the vertices that does not match with that of $G_{u}^{r}.$ Now, consider two games.
-One game corresponds to $1-$WL and another to Local $k-$WL.
-For the first $(k-1)$ moves, the Spoiler chooses to play and places pebbles at $(v_{1},v_{2},....v_{k-1}).$ After that, in both games, there are two pebbles and the position of both the games are same.
+One game corresponds to 11-WL and another to Local $k$-WL.
+For the first (k1)(k-1) moves, the Spoiler chooses to play and places pebbles at $(v_{1},v_{2},....v_{k-1}).$ After that, in both games, there are two pebbles and the position of both games are the same.
 Let $S_{1}$ and $D_{1}$ be the spoiler and Duplicator in the $(k+1)$ bijective pebble game, and $S_{2}$ and $D_{2}$ be the spoiler and Duplicator in $2$ bijective pebble game.
@@
 ∎
-Table 3: Here, $n:$ number of nodes in the graph, $n_{1}:$ max number of nodes in a $r$-hop neighbourhood of a vertex, $n_{2}:$ maximum number of nodes in any two consecutive layers for a particular vertex, over all the vertices of the graph, and $t_{k}:$ number of distinct graphs upto isomorphism of at most $k$-vertices.
-The first row compares the expressiveness of the models , and the second row compares the time complexity, with respect t0 our models.
-Local k -WL (Local+Layer) k -WL k- WL S_k M_k-1
-Expressiveness - - Less
-Less than
-Local $(k-2)-$WL
-Less
-Time $O(n\times n^{k+1}_{1})$ $O((n_{2})^{k+1}\times rn)$ $O(n^{k+1})$ $O(t_{k}n^{k}_{1}n)$ $O(n^{k+1}_{1}n)$
-9 Proposed Architecture
-In this section, we propose two GNN architectures for subgraph as well as induced subgraph counting.
-The design of these networks follows from the discussions about subgraph and induced subgraph counting in the previous sections.
-Given a graph $G$ as an input, we extract the rooted subgraphs $G_{v}^{r}$ for each vertex $v.$ Each of these subgraphs is rooted at $v.$ The value of $r$ depends on the structure to be counted.
-For example, we extract $1-$hop induced subgraphs for counting triangles.
-These subgraphs are then given as inputs to the proposed models.
-9.1 InsideOut GNN
-We conducted the experiments using two different architectures, InsideOutGNN and InSigGNN.
-Figure 3: Schematic of the InsideOutGNN Model
-The InsideOut GNN model (Figure 3) consists of an internal network and an external network.
-The internal network learns the local substructure counts for each rooted subgraph, $G_{v}^{r}.$ A subgraph are given as input pass through an Internal GNN model.
-This model consists of GINConv Xu et al.
-(2019) layers.
-The Internal GNN outputs the embeddings of the nodes present in a given subgraph.
-These node embeddings are then aggregated using a GlobalAddPool layer to get the embedding of the root node $v$ of the subgraph.
-This root node embedding is the feature vector for $v,$ as given in line $7$ of Algorithm 1.
-The GINConv layers simulate the WL iterations on the nodes of the subgraph.
-This root node embedding, $emb_{v},$ is then passed through a linear transformation layer to learn the local counts of the required substructure in the subgraph.
-Now, we take a union of all the subgraphs for a given graph.
-Using the node embeddings, $emb_{v},$ learnt from the Internal GNN model, the External GNN model learns the global count of the required substructure.
-The External GNN model also comprises of GINConv layers.
-The intuition behind this design is to make the models learn the local counts well.
-Once these local counts are learned, the global count is just a linear function of the local counts, which can be learnt easily by the External GNN model.
-9.2 InSigGNN
-For the InSig model, the Internal GNN component is similar to that of the InsideOut GNN model.
-The architecture is shown in Figure 4.
-However, in this model, we do not transfer the node embeddings learnt from the Internal GNN part of the model.
-Instead, we add all the local counts predicted by the internal part of the model and pass the sum as an input to the external model, which is just a linear transformation layer.
-The weights learned in the Internal GNN model depend only on the local counts of the substructures in the rooted subgraphs.
-The external linear transformation is learned based on the global count of a substructure in the entire input graph.
-Figure 4: Schematic of the InSigGNN model
+9 Experiments
 9.1 Implementation Details
-We experimented with both InsideOutGNN and InsigGNN models to predict the count of structures such as triangles, 3stars, 2stars, and Chordal cycles.
-For each of these structures, the true local counts are computed using various methods.
-For example, if our task is to count the number of triangles in a particular subgraph, for the root node of the subgraph, we label its local count to be the number of edges present in a 1-hop neighbouring subgraph without the root node.
-Using these local counts, from Equation Lemma 5, we assign the global counts to be $\frac{1}{3}\sum\limits_{u\in V(G_{v}^{1})}\prescript{x\mkern-0.5mu}{}{C}_{2}$, where $x$ is the local count corresponding to the root vertex $v.$ Here, $3$ is the orbit of a node in a triangle.
-Similarly, we assign the local counts and the global counts for different substructures.
-When counting larger substructures such as $K_{4}$ and $C_{4},$, we use the fragmentation technique with the help of the InSignGNN model.
-For calculating $K_{4}$, we construct a set of subgraphs.
-For every node in the graph, a 1-hop neighbouring subgraph is generated, from which we remove the root node.
-In this subgraph, we count the number of triangles and process accordingly.
-Therefore, we use the model already trained for counting triangles in order to count $K_{4}.$ During the inference, we apply a rounding function, as the counts are integer numbers.
+We refer to our proposed model as InSig.
+In the experiments, we predict the counts of the following substructures occurring as subgraphs : triangles, $3$-Star, 22-Star, chordal $C_{4}$, $K_{4}$, $C_{4}$, and tailed triangles.
+Task
+Total pattern
+count
+Zero Count
+graphs
+Standard Deviation
+of Count
+Average number
+of Nodes
+Average number
+of Edges
+Number
+of graphs
+Triangle $25209$ $195$ $3.072$ $18.7976$ $62.678$ $5000$
+2-Star $429463$ 0 $18.015$
+3-Star $309525$ 0 $17.777$
+Chordal $19088$ $1786$ $4.742$
+K4 $643$ $4447$ $0.387$
+C4 $53002$ $16$ $6.938$
+Tailed Triangle $177968$ $195$ $25.943$
+Table 3: Dataset statistics. The total number of graphs in the dataset is $5000.$ We used $4000$ graphs for training, $500$ for validation and $500$ for testing.
+When counting larger substructures like $K_{4}$, $C_{4}$, and Tailed Triangles, we use the fragmentation technique with the help of a model learned to predict triangles, 3-Star, 2-Star, and Chordal $C_{4}$.
+For counting $K_{4}$, the task of Pattern Counting Component is to learn which edges to prune.
+Later Local Count Learning component counts the number of substructures, which in the case of $K_{4}$, is triangles.
+Once the number of triangles is predicted using models learned to count triangles, the normalization factor is learned to output the global count.
+In other structures like tailed triangles, we have two patterns to learn: the nodes in the $1$-hop neighborhood and the edges between them.
+During the inference phase, we use a rounding function, as the counts are integer numbers.
 9.2 Hyperparameters
-We use two GIN Convolutional layers for the Internal GNN part in both proposed models.
+We use two GIN Convolutional layers for the Pattern Learning Local Count Learning component.
+We consider Linear transformations as readout layers in the previously mentioned components.
+Since we need to classify whether an edge should be present or not in a subgraph for counting a subpattern, we use Binary Cross-Entropy (BCE) loss for the pattern learning component.
+For the local counting and global counting components, we use Mean Absolute Error (MAE).
 For the models we have considered, as there can be paths of lengths more than 1 in the subgraph, $2$ GINConv layers are sufficient to capture the information well.
-In the InsideOut model, we use $2$ GINConv layers in the external part of the model as well.
-We use a learning rate of $41e-$ 4 and a batch size of $1$ We also experimented with different hidden dimensions for the node embeddings and obtained the best results when we used $4$ as a hidden dimension size.
-The ablation studies have been mentioned in Section 10.2.
+We use a learning rate of 1e41e-4 and a batch size of $1.$ We also experimented with different hidden dimensions for the node embeddings and obtained the best results when we used $4$ as a hidden dimension size.
 The experiments were conducted using an NVIDIA A100 40GB GPU.
-10 Experimental Observations
-10.1 Results
-Without Fragmentation Fragmentation
-Models Triangle 3-star 2-star Chordal $K_{4}$ $C_{4}$ Tailed triangle
-ID-GNN 6.00E-04 NA NA 4.54E-02 2.60E-03 2.20E-03 1.05E-01
+The source code of the implementation is available at this Github Link.
+9.3 Experimental Results
+Models Without Fragmentation Fragmentation
+Triangle 3-Star 2-Star Chorcal C4 $K_{4}$ $C_{4}$ Tailed Triangle
+ID-GNN 6.00E-04 NA NA 4.52E-02 2.60E-03 2.20E-03 1.05E-01
 NGNN 3.00E-04 NA NA 3.92E-02 4.50E-03 1.30E-03 1.04E-01
@@
 PPGN 3.00E-04 NA NA 1.50E-03 1.65E-01 9.00E-04 2.60E-03
-I2-GNN 3.00E-04 NA NA 1.00E-03 3.00E-04 1.60E-03 1.10E-03
-InsideOut 0.00E+00 0.00E+00 0.00E+00 3.82E-03
-InSig 0.00E+00 0.00E+00 0.00E+00 0.00E+00 0.00E+00 0.00E+00 0.00E+00
-Table 4: MAE for the subgraph count of different patterns.
-Some results, such as $2-$star, $3-$star, are not conducted by the given baselines; therefore, it is mentioned NA.
-As mentioned in Section Section 10.1, InSig Model has a lesser number of parameters while giving the same performance as InsideOut Model; therefore, we used only InSig Model for fragmentation based counting.
-We conducted our experiments***The dataset and the code are present in the anonymous repository[Link].
-for counting the following patterns: 1.
-triangles, 2.
-$3$stars, 3.
-$2$stars, 4.
-chordal cycles, 5.
-$K4,$ 6.
-$C4$ and 7.
-tailed triangles.
-The dataset used for the experiments is a random graphs dataset prepared in Chen et al.
-(2020).
+I2-GNN 4.00E-04 NA NA 1.00E-03 3.00E-04 1.60E-03 1.10E-03
+InSig 0.00E-$0$ 0.00E-$0$ 0.00E-$0$ 0.00E-$0$ 0.00E-$0$ 0.00E-$0$ 0.00E-$0$
+Table 4: MAE for the subgraph count of different patterns. Some results, such as 22-star, 33-star, are not conducted by the given baselines; therefore, it is mentioned NA.
+The dataset used for the experiments is a random graphs dataset prepared in (chen2020can).
+In Table 3, we report the dataset statistics, specifically the counts of the various patterns used for our experiments, number of graphs where these patterns do not appear, and so on.
 We report our experiments’ Mean Absolute Error (MAE) in Table 4.
-We compare our results with those reported in ID-GNN You et al.
-(2021), NGNN Zhang and Li (2021), GNNAK+ Zhao et al.
-(2022), PPGN Maron et al.
-(2019) and I2-GNN Huang et al.
-(2023).
-We can observe that our approach of predicting the local counts first and then predicting the global counts, achieves zero error for almost all of the tests.
-Although we experimented with two models, we realise that InSig model needs much fewer parameters to learn yet gives the same results.
-For the structures where we used fragmentation, we apply the models for counting triangles and 2stars in order to count $K_{4}$ and $C_{4}$, respectively.
-Also, we use the InSig model to predict the degree of a node and the number of edges in the induced subgraph without the root node.
-We call them the Local Node and Local Edge models, respectively.
-Using the Local Node and Local Edge models, we count the number of Tailed Triangles.
+We compare our results with those reported in ID-GNN (https://doi.org/10.48550/arxiv.2101.10320), NGNN (zhang2021nested), GNNAK+ (zhao2022from), PPGN (https://doi.org/10.48550/arxiv.1905.11136) and I2-GNN (huang2023boosting).
+We can observe that our approach of predicting substructures, the local counts, and then predicting the global counts, achieves zero error for all of the tests.
 For all the patterns, we observed that the model gets to zero error after only 2 to 3 epochs.
+From Table 5, it can be observed that our approach requires a considerably smaller number of parameters and beats all the baselines.
+Our inference time comprises model inference as well as the graph preprocessing time, where it creates the set of subgraphs corresponding to each node in the graph.
 Models
@@
 Parameters
-Inference Time
-Memory Usage
+Inference
+Time (ms)
+Memory
+Usage (GB)
 ID-GNN 102K 5.73 2.35
@@
 I2-GNN 143K 20.62 3.59
-InSig 190 20 1.2
+InSig 268 20 1.2
 Table 5: The table shows the comparison of the number of parameters required by the baselines and InSig Model.
 The values shown here correspond to the triangle counting task with the hidden dimension set as $4$.
-It can be observed that our approach requires a considerably smaller number of parameters and beats all the baselines.
-Our inference time comprises model inference as well as the graph preprocessing time, where it creates the set of subgraphs corresponding to each node in the graph.
-10.2 Ablation Study
-We varied the hidden dimension of the GINConv layer by taking the hidden dimensions from the following set $\{2,4,8,16,32,64\}$.
-The loss corresponding to different hidden dimensions is plotted in Figure 5.
-The experiment was conducted using the InSig Model for counting triangles.
-We observe that our model was able to compute exact counts with just 4 hidden dimensions.
-Figure 5: Ablation Study - Validation Loss VS Hidden Dimension
-We also varied the number of GINConv Layers.
-The observations are present in Table 6.
-The task for this ablation study was counting triangles.
-As per the previous study, we set the hidden dimension as $4$.
-We observe that we two GINConv Layers we get the best result.
-It should also be noted that with other sets of layers, the loss is better than the baselines.
-Number
-of
-Layers
-Number of
-Internal
-Model
-Parameters
-Number of
-External
-Model
-Parameters
-Total
-Parameters
-Validation
-Loss
-Test
-Loss
-1 53 97 150 1.41E-04 4.24E-04
-2 93 97 190 0.00E+00 0.00E+00
-3 133 97 230 2.35E-05 9.42E-05
-Table 6: This table shows the ablation study with the number of GINConv Layers in the proposed architecture.
-The fragmentation technique helps us to use the previously learned models, such as the model to count triangles to count the number of $K_{4}$ in a graph.
-Although the loss corresponding to that experiment is zero, as mentioned in Table 4, we did another study where we used a different dataset to train previous models of fragmentation.
-For example, we trained the InSig model to count triangle and 2stars on $5000$ Erdos-Renyi random graphs Chen et al.
-(2020).
-Using the respective trained models, we predicted $K_{4}$ and $C_{4}$ counts in the random graphs dataset prepared in Chen et al.
-(2020).
-The reverse study was also conducted.
-In both cases, the loss was zero.
-This shows that our proposed architecture is not dependent on the statistics of the dataset.
-The fragmentation shows that a trained model to count the number of smaller patterns on any dataset can be used to count the larger patterns.
+(a) Figure shows an example of $G^{1}_{0}$ for a random graph with $10$ nodes
+(b) t-SNE plot of the updated node embeddings from the Pattern learning component
+Figure 4: Figure showing the input and updated node embedding from the pattern learning component
+In Figure 4, we show an example of a graph sent as input to the pattern learning component and the updated node embeddings outputted from the model.
+In Figure 4(a), we see that the root node is node 0 and there are directed edges from the root node to its neighboring nodes.
+Figure 4(b) shows the updated node embeddings of each node in the graph.
+We can observe that $8,2,4,6$ nodes can be separated using some separator from the rest of the nodes.
+This indicates that the model is able to learn the updated node embedding such that we can distinguish nodes which has edges between them.
 10 Conclusion
-In this work, we progressed toward a more precise characterization of the localized versions of WL algorithms.
-We showed how the Local-$k-$WL lies between the global $k$ and $k+1-$WL in terms of expressiveness.
-We also developed strategies to make the Local-$k-$WL algorithm more efficient by introducing techniques such as layered WL, recursive WL, and pattern fragmentation.
-The hope is that such generalizations of the WL algorithm will lead to a finer subdivision of the WL hierarchy as well as more efficient and expressive graph neural networks.
+Subgraph counting is a fundamental combinatorial problem that arises in the study of graphs and graph structured problems as well as data.
+Exactly counting the number of subgraphs in a graph is also a computationally hard problem.
+In this paper, we present a learnable algorithm that is able to compute exact counts of a number of commonly occurring patterns.
+The proposed fragmentation method has proven to be beneficial for the task of counting subgraphs.
+Additionally, since fragmentation results in smaller subgraphs, the parameter requirements of the proposed GNN-based architecture is orders of magnitude less than previous methods.
+As future work, we plan to analyse the fragmentation algorithm from a theoretical perspective.
+We also plan to investigate the effectiveness of this technique for increasing the expressiveness of GNNs for downstream tasks.
+Along with that, the current approach is limited to patterns that can be computed using $1$-hop or 22-hop subgraphs, hence we can study our algorithm for bigger patterns.